A Complete Bibliography of Publications in Journal of Parallel and Distributed Computing: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

03 June 2024 Version 2.18

Title word cross-reference

2 [CFNH24, DQH^{+}21, FLFZTS20, JP22, ZLCL24]. 3 [BHB^{+}21, MCC20, NE23]. 8 × 8 [JP22]. g [GSM22, MAL^{+}23]. Γ [WYH^{+}24]. h [MAL^{+}23, ZZLM22]. K [FBL^{+}21, LLM^{+}20, AMM21a, Amm21b, CSS21, CTA20, LLM24, LFC^{+}24, NA23, YZM23, ZLCL24]. κ_{3} [WHC21]. G [RBS21]. SSI − FL [UHAH^{+}24]. n [LLM24, LFC^{+}24, MZMM21, ZLCL24]. n_{4} [PB20].

-and- [MAL^{+}23]. -ary [LLM24, LFC^{+}24, ZLCL24]. -component [YZM23]. -connectivity [WHC21]. -core [CSS21]. -coverage [AMM21a, Amm21b, NA23]. -CSqu [NA23]. -cube [LFC^{+}24]. -cubes [LLM24, MZMM21, ZLCL24]. -disjoint [ZLCL24]. -extra [GSM22]. -means [FBL^{+}21, LLM^{+}20]. -mismatch [CTA20]. -packing [FLFZTS20]. -restricted [ZZLM22]. -robust [WYH^{+}24].

/Special [Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t].

1 [IKK20]. 1-Full [Ano20w, Ano24u]. 100 [BRK^{+}21]. 11 [QM21]. 148 [YWF23]. 183 [SSA24a]. 19 [BMC^{+}24].

2020 [Ano20z, Ano20-29, Ano20y, Ano20-27, Ano20-30, Ano20-31, Ano20-28]. 2021 [Ano21-32, Ano21-29, Ano21-34, Ano21-27, Ano21-35, Ano21-30, Ano21-28, Ano21-33, Ano21z, Ano21-36, Ano21-31, Ano21y]. 2022 [Ano22z, Ano22-30, Ano22-36, Ano22y, Ano22-35, Ano22-31, Ano22-29, Ano22-28, Ano22-33, Ano22-34, Ano22-27, Ano22-32]. 2023 [Ano23a, Ano23z, Ano23-27, Ano23-30, Ano23-28, Ano23-32, Ano23-33, Ano23-29, Ano23-35, Ano23-34, Ano23-31, Ano23-36]. 2024 [Ano24a, Ano24b, Ano24l, Ano24v, Ano24x, Ano24w, Ano24y, Ano24z, Ano24-28].

3 [QM21]. 3.0 [DM20].

5G [CN22, CL22b, Jar20, WHY^{+}21]. 5G/beyond [CL22b].

Abaci [SWF^{+}22]. Abaci-finder [SWF^{+}22]. ABC [CGG^{+}23, GKB^{+}20]. ABE [GLY^{+}21]. abstraction [VGTSG^{+}21]. abstraction-level [VGTSG^{+}21]. abstractions [DGWD21, JB20]. Accelerated [NRdA^{+}20, JKM^{+}22, MN24, MMA22, YNI^{+}22]. Accelerating [CSS^{+}23, DW23, LLG^{+}24, OZ22, STG^{+}20, CFNH24, NZ23]. Acceleration [SALP20, BGDMF^{+}24, KTMB22, NMA^{+}24, SBL20, YSZL23]. accelerator [CSY^{+}24, MKP22, MRB20, TAG23, ZCY^{+}21]. accelerators [CaTZ^{+}24, HGC^{+}23]. acceptance [BMC^{+}24]. access [ARL20, BKL^{+}20, CL22b, HZY^{+}21, JTV^{+}22, KYZ^{+}20, OK21, RGB20, TTD24, WCG^{+}24, ZTKL^{+}21]. accountability [CLT^{+}20]. Accountable [BKL^{+}20]. accuracy [SFML21]. Accurate [ZZS^{+}21a]. achieve [YTLF22, ZLD22]. Achieving [GLY^{+}21, JHZ20, SSA24a, SSA24b, MCD^{+}21]. ACO [CH23]. ACO-based [CH23]. across [SPP^{+}23]. actions [MLB21]. active [ZYW24]. acyclic [SSH23, TZZ^{+}20a, YSZL23]. Ad [ABB22, SLFC22]. adaptation [AGC^{+}21, KG20, SUD^{+}22]. Adaptive [DF22, GXYH21, KRK20, LLW^{+}20, MM24, NGS21, WXAL22, XLL^{+}20, BCM^{+}21, BSY24, IWS22, JYH22, JJJ21, JKK^{+}23, LLM^{+}20, RK24, RGESG^{+}21, SADM24, SFT^{+}21, WYH^{+}23, WCMS24, ZLD22, ZJW22]. adaptively [LKAB^{+}22]. Address [VGMG20, SLZP24]. Addressable [AS21]. addressing [JTV^{+}22]. ADI [AR20]. ADI-type [AR20]. Adjusting [PYYO22, LKAB^{+}22]. ADMM [SBL20]. adoption [GMS^{+}21]. Advances [MMM22, dARR21, MS20]. adversaries [JYH22]. aerial [SLSN24, ZSL^{+}23]. AES [LFJ^{+}20]. Affinity [MYS^{+}23]. Affinity-aware [MYS^{+}23]. Agent [GGTSFD23, BCM^{+}21, CMR20, TB22, WXZ^{+}23]. agents [MLB21, PK21b]. aggregation [GC23]. aging [MPR^{+}21]. agreement [XLL^{+}21]. agricultural [FFS^{+}22]. AI [CSY^{+}24, KGTK20, TlW23]. AI-specific [CSY^{+}24]. AIoT [LXC^{+}22]. air [MD20]. air-jet [MD20]. airborne [SW22]. airfoil [CGL^{+}22]. ALBERT [CWHC22]. algebra [MFAB23, VLCM^{+}20, ZAB20]. algorithm [AAZMS20, CHJ^{+}24, CH23, DCM^{+}23, DT21, DED^{+}20, DPSD21, FFGEL21, FLFZTS20, GRZT22, GHD20, GSMÖ23, GMA^{+}22, HKTG20, HMCG20, HWR^{+}24, IRA20, IMP^{+}23, KKH^{+}23, KD21, KSS23, LLM^{+}20, LWHF22, LLYZ23, LFRBGV^{+}21, MN24, MMA22, MST24, NMA^{+}24, PC21, PF22, PSBB21, QZW^{+}24, RdCR^{+}24, SS21, ŠK23, SWM23, TLD^{+}23, TYOC24, TYM^{+}22, VP20, VJR20, WCLD21, WC22, WCTW22, XBX^{+}22, YWF21, YWF23, YS21, YNI^{+}22, ZZZG21, ZZ22, ZZZ^{+}20, CH24, TSTY22, AGMG23, SW22]. Algorithmic [Stp20, GHKKL23]. Algorithms [JTV^{+}22, LLXX24, BL23, BLB^{+}20, CDY23, CTA20, FPGLSA24, FW23, GK24, JYH22, KGPT23, LFC^{+}24, MBM^{+}20, MS20, Mar20, NRGL22, PB20, PK21a, Sok21, YHKR20, ZAB20]. aligned [ZYL^{+}21]. all-pair [CTA20]. all-to-all [PW21]. allocation [CN22, DVS24, HWM^{+}23, KK21, LLR^{+}21, XTGJ21, ZZZ^{+}23]. allreduce [NMA^{+}24]. amalgamation [KGTK20]. AMBLE [PYYO22]. Amnis [XPW^{+}22]. analogy [GSA21]. analyses [GPC23]. analysing [BDFG21]. Analysis [GKTW21, LCK23, TMB^{+}21, VP22, ACCN20, Ati20, BMK^{+}22, BPBD23, BV21, CML^{+}24, CC23, CL22a, DWR^{+}23, HMS20, KST^{+}23, LTBY20, LHL21, LAPB20, QGP24, Sch24, SCZ24, SFZ23, VFB^{+}24, YZM23, ZYL^{+}21, ZXY21]. analytic [AKY20]. Analytical [EHH^{+}23, ACC^{+}23, LCK23]. analytics [ESA24, GKSS24]. analyze [WYH^{+}21]. analyzing [GMS^{+}21]. anchor [SM22]. anchor-assisted [SM22]. Angara [MST24]. annealing [VBB22]. Anomaly [AGMG23, GKB^{+}20, HR23, MGE20, RKAA20, WLZ20]. Anomaly-based [AGMG23, HR23, RKAA20]. anonymity [JZWX20]. anonymous [LCZL21, LZY23]. ant [LYC23, ZFL^{+}23]. Antecedent [Uhl20]. antennas [SMMG20]. Anthropomorphic [WLL^{+}23b]. anti [PWL^{+}22]. anti-interference [PWL^{+}22]. Antipaxos [MGW24]. any [Jea22]. AOR [SOS^{+}24]. Apache [CYWL21]. APC [CN22]. APC-based [CN22]. API [DFP20, HZY^{+}21, SM23]. API-images [DFP20]. Appends [CAG^{+}23]. Application [IWS22, SW22, GRZT22, GSMÖ23, GLL21, HZY^{+}21, JKM^{+}22, MSRB20, SWM23, WZC^{+}20, XZY^{+}23]. Application-aware [IWS22]. application-level [WZC^{+}20]. Applications [KSB^{+}20, AKY20, ACC^{+}21, BPBD23, BHP^{+}24, BSWO23, CMFV^{+}20, CL22a, DMSB20, DM20, DK24, EL20, GF20, GM21, GBC^{+}22, IRLN23, KPS^{+}22, KGP^{+}21, LYF^{+}24, LLD^{+}23, MTR22, MYS^{+}23, NPO^{+}23, NMA^{+}24, PS22, RLW^{+}24, RH20, STG^{+}20, SPBR20, SALP20, SMMG22, VGTSG^{+}21, VD21, WXAL22, WLL^{+}23b, XJR21, YM21]. applying [KSLN24]. Approach [AkBA^{+}20, AMM21a, AKL22, BS24, BRK^{+}21, CSY^{+}24, CGL^{+}22, CMFV^{+}20, DWWX22, DF21, DSZ^{+}21, DHF23, DZZ^{+}23, GJL^{+}24, HKTG20, HLL^{+}22, HLBZ20, JSJC22, KuR24, KHO22, KKG^{+}23, LFJ^{+}20, LYF^{+}24, LZZ^{+}20, NT20, NTT^{+}23, NHR22, PXY^{+}20, PS22, QM21, SSG24, SAATK21, Sch24, SNSK20, TZC^{+}24, URC20, VTT^{+}22, WCR^{+}20, WYH^{+}24, YSMB21]. approaches [QGP24, VFB^{+}24]. Approaching [CCSI21]. Approximate [KGPT21, CSS21, GC23, KGPT23, RSSP23, WZO^{+}21]. approximating [ŠK23]. approximation [SK21]. approximations [MAL^{+}23]. apps [MRC21]. April [Ano21-32, Ano22z, Ano23-30, Ano24a]. APT [HLZ22, LLXG21]. arbitrable [TY23]. Arbitrarily [BLB^{+}20]. Architecting [SOL22]. architectural [Gow21, SKS21]. Architecture [BDL22, CRS22, AHAB23, AZF^{+}24, AAZMS20, AGC^{+}21, BRK^{+}21, DMKFJ20, FQL^{+}23, GF20, HCY^{+}21, PD21, PAD22, SLSN24, SFZ23, SOS^{+}24, Yaz23b, YM21]. architectures [CCAACS21, CIH^{+}23, EMCE20, FRAK23, FSL^{+}21, HCC^{+}20, JB20, MTG^{+}24, NE23, TV22]. archive [ZTKL^{+}21]. Area [AJH^{+}20, Amm21b, LZZ^{+}20, PD21, PAD22, SADM24]. arguments [ACR23]. arithmetic [WZO^{+}21]. ARM [ZYL^{+}21]. arrangement [ZZLM22]. array [DQZZ21, SBSB20]. arrays [AT24]. Artificial [SATJ^{+}20, dARR21, BJ23, GKB^{+}20]. ARVMEC [XLL^{+}20]. ary [LLM24, LFC^{+}24, ZLCL24]. Ascend [CSY^{+}24]. ascent [MMA22, PXY^{+}20]. aspects [GHKKL23]. assessment [AM20, BAC22, LLM24]. assignment [KN24]. assisted [ARL20, LMG^{+}21, Li24, MLTT20, SM22, XLL^{+}21]. assumptions [BL23]. asymptotic [Ati20, Sch24]. Asynchronous [CAG^{+}23, VBB22, DG22, DGFR21, PXY^{+}20, SMMG22, ZLD22]. asynchronously [BLNP23]. atom [SPBR20]. atom-based [SPBR20]. Atomic [CAG^{+}23, GHNS22, MM24, RdCR^{+}24]. attack [HA21, HR23, ZSL^{+}23]. attacks [DT21, KKT^{+}22, MPAS24, TZDC21]. attribute [BKL^{+}20]. Auction [KK21, SWM23]. auction-based [SWM23]. audit [TY23, XZH^{+}22]. auditable [HSX^{+}21]. auditing [CML^{+}24, GPK21]. auditor [HR23]. augmented [KBW20, Pou20, WHC21, ZXY21]. August [Ano20z, Ano21-29, Ano22-30, Ano23-28, Ano24b]. authenticated [BKL^{+}20]. authentication [ABMPL22, ABB22, JZWX20, LCZL21, Li23, MBB22, OK21, VD21, XLL^{+}21]. authorized [YQZ^{+}20]. Auto [ESA24, DK24, DZZ^{+}23, GA21, PSBB21, VLCM^{+}20]. auto-scaling [DK24, GA21]. Auto-tuned [ESA24, PSBB21, VLCM^{+}20]. auto-tuning [DZZ^{+}23]. Autoencoders [DFP20]. automata [PAD22, SKK21, SFZ23, RSGA20]. automated [HFP^{+}22, XLCL20]. Automatic [LTBY20, BkB^{+}23, CWHC22, VLCM^{+}20]. automation [PAD22]. automaton [PD21]. Automotive [TMB^{+}21]. autonomic [JZS^{+}20]. autonomous [SLSN24, TAG23]. autotuning [FHN^{+}22]. availability [LSZL20, TK23]. available [KKW23]. avoidance [PC21, SLSN24, SOL22]. Avoiding [AA21]. Aware [MLR^{+}23, Ato23, BMK^{+}22, CPZ^{+}20, CLW^{+}23b, CLMH22, CTFW22, HKTG20, HLBZ20, HWR^{+}24, IMP^{+}23, IWS22, JHZ20, JKK^{+}23, KD22, KHO22, LCH^{+}21, LZS^{+}24, MSRB20, MYS^{+}23, MCC20, NHR22, OZ22, PC21, SATJ^{+}20, TZZ^{+}20b, TZC^{+}24, WWH^{+}21, WYH^{+}23, WHL^{+}23, ZDZ^{+}21, ZZ22, ZFL^{+}23].

B [CGG^{+}23]. B5G [CL22b]. backward [WCT^{+}22]. balanced [GS20, LZZJ23, SOL22]. balancer [FPdLS^{+}21]. Balancing [KSB^{+}20, AM22, BAC22, DED^{+}20, GGTSFD23, MBM^{+}20, PSBB21, TDCM21, VP20, ZZS^{+}21b]. ball [ŠK23]. band [SMMG20]. bandwidth [BHP^{+}24]. bandwidth-sharing [BHP^{+}24]. barrier [CSS21, SAATK21]. barrier-reinforcing [SAATK21]. based [ABMPL22, Ala24, AMP20, AMM21a, ACCN20, AKS^{+}20, AGMG23, AKL22, ABB22, BW22, BSY24, BDSQO22, BGDMF^{+}24, BKL^{+}20, BPBD23, BPT^{+}22, BDRJ24, CLT^{+}20, CLW^{+}23a, CWHC22, CLZ^{+}22, CN22, CGK20, CH23, CL22a, DFP20, DWW^{+}21, DWWX22, DGWD21, DSZ^{+}21, DQZZ21, DZ24, DK24, DZZ^{+}23, DKS21, EMSEMM20, FQL^{+}23, FPdLS^{+}21, FFS^{+}22, GLC^{+}22, GJL^{+}24, GKB^{+}20, GLF20, GVC^{+}22, GSMÖ23, GVI24, GPC23, GLY^{+}21, HWM^{+}23, HA21, HGC^{+}23, HNKÖ21, HFP^{+}22, HR23, HLK^{+}22, HFA20, HZY^{+}21, HSX^{+}21, JSJC22, JL23, JZS^{+}20, KK22, KVMR23, KuR24, KBS^{+}21, KTM^{+}21, KTMB22, LCZL21, LZGL22, LWHF22, LLGC22, LWL^{+}22, LLX^{+}23, LWW^{+}23, LYF^{+}24, LLXX24, LD21, LYZ^{+}22a, LLD^{+}23, LZZ^{+}20, LHL21, LZWZ22, LJH^{+}22, LYZ^{+}22b, LGZZ23, LZY23, MLTT20, MRB20, MST24, MLR^{+}23, NRdA^{+}20, NGS21, NHR22, NRGL22, PJV^{+}22, PD21, PAD22, PS22, PSU^{+}21, PLBG21, PG20, QBS21, QZW^{+}24]. based [QDD^{+}22, QMB21, RGB20, RWF^{+}21, RKAA20, Rub22, SHI22, SKB21, SS21, SCZ24, SM23, SPBR20, SWM23, SEM20, SKS21, SATJ^{+}20, Stp20, SFZ23, TK23, TDL^{+}22, TLD^{+}23, TYOC24, TTD24, TNM^{+}22, TYM^{+}22, UHAH^{+}24, VLCM^{+}20, VFB^{+}24, WMJ^{+}20, WHY^{+}21, WYH^{+}21, WLL^{+}23a, WXZ^{+}23, WCG^{+}24, WDY^{+}24, XLCL20, XWCJ22, XZH^{+}22, XZY^{+}23, XLL^{+}21, XRBT21, XDM^{+}22, YWF21, YWF23, YZM23, YS21, YLL21, YSMB21, YLZ^{+}20, ZWS^{+}20, ZYL^{+}21, ZNX^{+}21, ZLS23, ZCY^{+}21, ZFL^{+}23, ZZZ^{+}20, ZLLC22, ZD22]. Bat [RK24]. batch [BG21, PYYO22, ZdCLT22]. Batched [AR20, KNK^{+}23]. batches [ATD20]. Bayesian [PS22, YHKR20]. BCube [DLWF23]. bee [GKB^{+}20]. Behavior [LHL21, AB22b, WYH^{+}21]. behaviors [WLL^{+}23b]. belief [NVE^{+}21]. benchmark [MKP22]. benchmarks [GM21, MKP22]. beneficial [VBB22]. benefits [OS20]. best [ESA24]. between [Mar20]. beyond [CL22b]. BFT [ZLD22]. bi [CTGJ22, PB21]. bi-directional [CTGJ22]. biclustering [LFRBGV^{+}21]. bidirectional [YZC22]. Big [MMM22, AHAB23, AKY20, DF21, DZZ^{+}23, GQW^{+}21, JJJ21, LZZ^{+}22, LLW^{+}20, LBHW20, LYZ^{+}22b, QDD^{+}22, RLW^{+}24, SAY20, SEM20, WWL^{+}21, ZDZ^{+}21]. Bilateral [ZQL^{+}21]. billion [RSSP23]. billion-edge [RSSP23]. binary [LFRBGV^{+}21, YNI^{+}22, KVMR23]. bio [GMA^{+}22]. bio-inspired [GMA^{+}22]. biogeography [KK22]. biogeography-based [KK22]. bit [CSS^{+}23, PD21, YNI^{+}22]. bit-level [CSS^{+}23]. black [WCLD21]. black-box [WCLD21]. BLAS [IKK20]. blind [KSV^{+}20a]. Block [SBB20, AA20, GS20, LLG^{+}24]. block-diagonal [AA20]. Blockchain [ARL20, BPT^{+}22, CLT^{+}20, CGC21, HFP^{+}22, HSX^{+}21, KGTK20, NYZ^{+}20, Pou20, RGB20, WLZ20, YLZ^{+}20, CW20, CTGJ22, GLC^{+}22, GLF20, GLY^{+}21, HZL^{+}20, JJ22, KBS^{+}21, KTM^{+}21, KKT^{+}22, KKG^{+}23, LCZL21, LZD21, LWW^{+}23, LLG^{+}24, LTBY20, LZZ^{+}20, LHL21, LZY23, MLTT20, MGE20, NVE^{+}21, SMHK21, SZQ^{+}23, TTD24, URC20, VD21, WHY^{+}21, WDY^{+}24, XZH^{+}22, XLL^{+}21, XBX^{+}22, ZWS^{+}20, ZNX^{+}21, DM20]. Blockchain-assisted [ARL20]. Blockchain-based [BPT^{+}22, CLT^{+}20, HFP^{+}22, HSX^{+}21, RGB20, YLZ^{+}20, GLC^{+}22, WHY^{+}21, XZH^{+}22, XLL^{+}21]. blockchain-empowered [TTD24]. Blockchain-enabled [CGC21, KKT^{+}22, MLTT20]. blockchain-enhanced [HZL^{+}20]. blockchain-orchestrated [KKG^{+}23]. blocks [URS21]. Board [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano23r, Ano23s]. Board [Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k]. bonded [BNOS21]. Boosting [BADP22]. bound [GSMÖ23]. Bounding [JH21]. bounds [SHI22]. box [WCLD21]. BP [ZLLC22]. BPS [HZL^{+}20]. brain [CGW23]. branch [GSMÖ23, LW20]. branch-and-bound [GSMÖ23]. Breaking [CSS21]. broadcast [NRGL22, RdCR^{+}24]. BSF [Sok21]. BTNC [ZWS^{+}20]. bubble [KKH^{+}23]. bubble-sort [KKH^{+}23]. Budgeted [WYW^{+}20]. buffers [KLL^{+}21]. building [DMSB20]. burials [MM24]. burst [KLL^{+}21]. bus [HNN^{+}20]. bypassing [EL20]. Byzantine [CAG^{+}23, PB21].

C [CGG^{+}23, KGPT21, QM21]. C11 [QM21]. C11/C [QM21]. Cache [TLC20, CGK20, KSLN24, LZL22, NMPS20, PB20, SAY20, SWY^{+}21]. cache-oblivious [SAY20]. caching [HLX^{+}22, LZGL22, LJH^{+}22, ZWSL22]. calculation [WWL^{+}21]. calculations [CZR^{+}24]. camera [BCM23]. can [CDA20]. cancer [VJR20]. capability [ABB22]. capacitated [LLR^{+}21]. capsule [dRBB21]. CAPTCHA [AMP20]. Capuchin [AGMG23]. carbon [XB20, ZZ22]. carbon-aware [ZZ22]. Carlo [DPSD21]. cars [TAG23]. Case [TLC20, GM21, SMMG22]. categorical [LZZ^{+}22]. causal [SSH23, ZWCL21]. causality [KSV20b]. CC [KGPT23]. CCFTL [SLZP24]. ccountable [CGG^{+}23]. cell [AAZMS20, Li22a, PD21, DSZ^{+}21]. cellular [PD21, PAD22, SKK21, SFZ23]. center [BW22, BDRJ24, DLWF23]. centers [CW21a, FWZ^{+}20, FHG^{+}20, GVC^{+}22, GPH^{+}22, GGTSFD23, HSHT22, HLBZ20, KHO22, PC21, PPN^{+}20, SM23, ZZ22]. centres [MRPH20]. centric [CSS^{+}23, FDT^{+}24, GF20, SBBP20, ZBF^{+}24]. Certificateless [GPK21]. certified [AAD^{+}20]. CFD [DDC^{+}24, XJR21]. CFSec [AKS^{+}20]. chain [JJ22, LWL^{+}22]. challenge [MKP22]. Challenges [KGTK20, LZD21, FTA^{+}22, HQL^{+}22, IOG20, SSG21]. Challenging [ABMPL22]. chameleon [WDY^{+}24, KSB^{+}20]. change [Ada21]. channel [VTT^{+}22]. chaos [PJV^{+}22]. chaotic [LYF^{+}24]. character [MLTT20]. characteristic [ZCD^{+}21]. characteristics [KGP^{+}21, QDD^{+}22]. Characterization [ZZ24, GBEFBC24, RPM24]. characterize [KSV20b]. cheap [KEK^{+}20]. checking [WHY^{+}21]. checkpointing [SNSK20]. Chip [JHML21, BV21, KK22, RK24, SMMG20, SOL22, Yaz23b]. chips [AM20, ZDL^{+}24]. chunk [ALS23]. chunk-level [ALS23]. churn [HNKÖ21]. CIC [ZJW^{+}21]. CIC-PIM [ZJW^{+}21]. cities [Ala24, AKB^{+}20, CGC21, GF20, MLTT20, OK21, Pou20]. city [CJZ^{+}22, KPS^{+}22]. class [CRS22]. classical [MFAB23]. classification [CRS22, FHG^{+}20, LD21, LSC22, SWF^{+}22, URS21, URM23, VJR20, XLCL20, ZdCLT22]. classifying [GMS^{+}21]. client [CLW^{+}23a]. clock [KSV20b, PPN^{+}20]. clone [ZLS23]. closed [VP20]. Cloud [BGA^{+}21, DED^{+}20, FQL^{+}23, Kur21, MSRB20, RCVA22, WYZ^{+}24, XLL^{+}20, ZZS^{+}21b, ASHO20, AKY20, Ala24, AKS^{+}20, AAD^{+}20, BS24, BJ23, CLMH22, CTFW22, CML^{+}24, CGK20, DMKFJ20, DF21, DFL^{+}23, GKB^{+}20, GA21, GPK21, GSV21, GGTSFD23, HA21, HSHT22, HR23, HLBZ20, HLL^{+}21, HZY^{+}21, HZL^{+}20, HSX^{+}21, HCY^{+}21, IRA20, Jar20, JKK^{+}23, KuR24, KHO22, KSS23, KGTK20, LSZL20, LMG^{+}21, LCH^{+}21, LWHF22, LYF^{+}24, Li24, LCC20, LYZ^{+}22a, MBM^{+}20, NGS21, PS22, PPN^{+}20, QMB21, Rub22, SSG^{+}20, SNSK20, SM23, TK23, TZC^{+}24, TDCM21, VGMG20, WMJ^{+}20, WCT^{+}22, WYH^{+}24, XWCJ22, XB20, YSZL23, YY22, YQZ^{+}20, YLZ^{+}20, ZGTM24, ZDZ^{+}21, ZZ22, ZFL^{+}23, DVS24, KK21]. cloud-assisted [Li24]. cloud-based [Ala24, NGS21]. Cloud-Edge [FQL^{+}23]. Cloud-edge-end [WYZ^{+}24]. cloud-enabled [AKY20]. cloud-fog [DMKFJ20]. cloud-to-edge [YY22]. cloud/edge [BJ23]. CloudFNF [ASHO20]. Clouds [SBBP20, BKL^{+}20, CWHC22, HWR^{+}24, IMP^{+}23, LQX^{+}20, WWH^{+}21]. Cluster [KGPT21, ABB22, BMC^{+}24, DZZ^{+}23, GHT^{+}21, KGPT23, Sok21, SATJ^{+}20, YS21]. cluster-based [ABB22, DZZ^{+}23, YS21]. Cluster-combining [KGPT21]. clustering [CKS22, FBL^{+}21, GHD20, HMCG20, PJV^{+}22, WCTW22, ZZZ^{+}20]. clusters [ALS23, BGDT22, FFGEL21, LZS^{+}24, MYS^{+}23, NRGL22, SPBR20, WCMS24, XZY^{+}23, ZJW22]. CNN [CaTZ^{+}24, LKAB^{+}22]. co [YES22]. co-design [YES22]. Coalition [HLL^{+}21]. coalitions [GGTSFD23]. code [EHH^{+}23, ZLS23, ZT20]. coded [BW22, CPZ^{+}20]. codes [DO22, MRPH20, PB20]. coding [KD21, Zha23, LLC20]. coffer [GHT^{+}21]. coflow [CW21a, ZQL^{+}21]. coflows [Che23]. cognitive [SATJ^{+}20]. coherence [URM23]. Cohort [HLK^{+}22]. Cohort-based [HLK^{+}22]. collaboration [HLK^{+}22, PWL^{+}22, YSZL23]. Collaborative [FSL^{+}21, GMMP24, YZC22, YWF21, YWF23, CN22, DFL^{+}23, JKK^{+}23, LLD^{+}23, MGE20]. Collecting [AKB^{+}20]. collection [CLZ^{+}22]. collections [BYW^{+}22]. collision [CZR^{+}24, KNK^{+}23, SLSN24]. colocation [ZPN^{+}21]. colony [GKB^{+}20, LYC23, ZFL^{+}23]. colossal [VP20]. column [AA20, DQH^{+}21, DQZZ21]. column-overlapped [AA20]. combat [CW20]. combination [RGESG^{+}21]. combine [KD21]. combining [KGPT21, KGPT23, LJW^{+}22, OZ22, QM21, SFML21]. COMITMENT [AkBA^{+}20]. commodity [BRK^{+}21]. common [CTA20]. communicating [MLB21]. Communication [ODXX21, SHI22, AKS^{+}20, CCC23, DGA^{+}24, HZL^{+}20, KMS22, NMA^{+}24, PSU^{+}21, TB22, VD21]. communication-efficient [DGA^{+}24]. communications [CGC21, GF20]. Compact [BT20]. comparative [VFB^{+}24]. Compare [KW20]. Compare-and-Swap [KW20]. Comparing [SMT22]. comparison [HDJ21]. Compatibility [CFRGPM22]. compensation [YDX^{+}22]. completion [Che23, LoKS24, SZW^{+}22, WXZ^{+}23]. complex [ATD20, DF22, TV22, ZT20]. complexity [Ati20, RKAA20]. Component [LZWZ22, KTMB22, YZM23, ZXY21]. component-based [KTMB22]. components [BLNP23, ZAB20]. composite [AAD^{+}20, ZR22]. composition [BLNP23, IRA20]. compound [LZWZ22]. compressed [SLZP24]. compression [CCC23, DGA^{+}24, DFL^{+}23, DKS21, DW23, YDX^{+}22]. compromised [LLXG21]. Comput [YWF23]. computation [CDY23, CHJ^{+}24, FLZ^{+}20, GQW^{+}21, Li22b, LZZ^{+}22, LLYZ23, Sok21]. computational [CN22, NE23, Sch24, TLD^{+}23]. computationally [FQL^{+}23]. computations [SHI22]. compute [JL23, WHC21]. Computer [BDL22, DPEL24]. computers [WCTW22]. Computing [AkBA^{+}20, BDL22, DED^{+}20, MLR^{+}23, SSA24a, dARR21, AZF^{+}24, Alm20, AT24, BS24, CTKdS21, CFRGPM22, CLMH22, CDY23, CH23, DWWX22, DCM^{+}23, DAG24, DF21, DGMS20, DBAC^{+}22, DFL^{+}23, DK24, DWR^{+}23, FLZ^{+}20, FBL^{+}21, GMS^{+}21, HA21, HNN^{+}20, HR23, HKTG20, HLL^{+}21, HQL^{+}22, Jar20, KD22, KPS^{+}22, LLR^{+}21, LSZL20, LMG^{+}21, LZD21, Li22b, LZGL22, LLX^{+}23, LYF^{+}24, LLYZ23, LYZ^{+}22b, LYC23, LWD^{+}20, MSRB20, MB21, MS23, NRdA^{+}20, NE23, OGR^{+}24, PWL^{+}22, PGB^{+}22, QBS21, QMB21, RWF^{+}21, SSG^{+}20, SSG21, SWM23, Sok21, SMT22, SFZ23, TZZ^{+}20b, TDL^{+}22, TDCM21, TYM^{+}22, WMJ^{+}20, WCT^{+}22, WYH^{+}24, WCG^{+}24, XWL^{+}20, XB20, XPW^{+}22, XJR21, YQZ^{+}20, YLL21, ZJW^{+}21, ZZZ^{+}23, ZLLC22, KK21]. concept [CC23, KST^{+}23]. concepts [CTKdS21]. concise [JP22, MPAS24]. concurrency [MPR^{+}21, SC23]. Concurrent [TGFPRA20, MLB21, OGR^{+}24, PMV20, ZQL^{+}21]. condition [VBB22]. conditional [DLWF23, TSTY22]. configurable [ZWSL22]. configuration [DZZ^{+}23, GJL^{+}24]. configuration-based [DZZ^{+}23]. configure [MMESG^{+}21]. CONGEST [SK21]. congestion [JHML21, KGP^{+}21, SOL22]. Conjugate [TV22]. Connected [Amm21b, MMM22, DQH^{+}21, GPH^{+}22, GQX20, KBW20, NA23, ZAB20]. connection [WDL22, ZWS^{+}20]. connections [ZQL^{+}21]. connectivity [GPC23, GSM22, LZWZ22, WHC21, YZM23, ZXY21, ZZ24, ZZLM22]. conscious [PAD22]. Consensus [PB21, ABD^{+}23, ACHP22, DGFR21, Pou20, XBX^{+}22]. considerations [BRK^{+}21]. consistency [DMSB20, MCD^{+}21, MGW24, TA24]. consistent [SÖAOA20]. consolidation [AKL22, HLBZ20, KHO22, PC21]. consortium [LCZL21, WDY^{+}24]. constant [SDLM20]. constrained [CLMH22, HLL^{+}21]. constraint [Li24, ZD22]. constraints [ALP21, AT24, Ati20, GBC^{+}22, Mal21, NMPS20, TDCM21, WWH^{+}21, WHL^{+}23]. constructing [KKH^{+}23]. Construction [LFC^{+}24, XWJ^{+}23]. consumption [BCEH23, BSWO23, SBB21, WXZ^{+}23, YZC22, Yaz23a, ZLLC22]. container [DK24, YLL21]. container-based [DK24]. contaminant [SW22]. contention [SSD^{+}20]. contexts [CC23]. continuity [SLZP24]. continuous [BCEH23, NT20]. Continuum [RCVA22, YY22]. contract [YLL21]. control [GVI24, HZY^{+}21, IWS22, KD21, MRC21, OK21, RGB20, SLSN24, TTD24]. Controller [EMSEMM20, MMESG^{+}21]. convergence [FHN^{+}22]. convex [CFNH24, TSTY22]. conveying [MD20]. convolution [BDSQO22, SNT^{+}20, WCMS24]. Convolutional [WLZ20, AGMG23, BDRJ24, GA21, KTMB22, LXC^{+}22, XLCL20]. Cooperative [KJA^{+}22, CDTS24, CDY23, SADM24, SC23, XTGJ21, ZZS^{+}21b]. Coordinate [TDL^{+}22, PXY^{+}20]. Coordinate-based [TDL^{+}22]. Coordinated [NMPS20, GK21]. copy [HCC^{+}20, HLZ22]. Core [OZ22, BV21, CSS21, HMCG20, JB20, NMPS20, NRGL22, QBS21, SBSB20, SOS^{+}24, TlW23, TLC20, XWJ^{+}23, XZPL24]. Core-aware [OZ22]. cores [Ato23]. correctness [AR21, KKW23]. Corrigendum [SSA24a, YWF23]. corruption [AA21]. Cost [AAD^{+}20, HWR^{+}24, KYZ^{+}20, SSA24a, YPD^{+}20, BMK^{+}22, CKS22, DLL^{+}21, DZZ^{+}23, FWZ^{+}20, HSHT22, HLX^{+}22, JHZ20, LJH^{+}22, LWD^{+}20, SSA24b, TB22, WCR^{+}20]. Cost-aware [HWR^{+}24, BMK^{+}22]. cost-driven [LJH^{+}22]. Cost-effective [AAD^{+}20, HLX^{+}22, WCR^{+}20]. cost-efficient [DZZ^{+}23, LWD^{+}20]. costs [NPAM20]. Coulomb [HMS20]. counters [FHN^{+}22]. coupled [JZS^{+}20]. Coupling [ARL20]. courses [GMS^{+}21, QBS21]. Cover [Ano20x]. coverage [AMM21a, Amm21b, NA23]. covering [AT24]. covers [ZLCL24]. COVID [BMC^{+}24]. COVID-19 [BMC^{+}24]. CP [YDX^{+}22]. CP-SGD [YDX^{+}22]. CPU [BMK^{+}22, Gow21, HCC^{+}20, JKK^{+}23, LGZZ23, SFT^{+}21, WYH^{+}23]. CPU-GPU [LGZZ23, WYH^{+}23]. CPUs [LLM^{+}20, SFML21, TLC20]. cracking [SAY20]. crash [DGFR21, SWF^{+}22]. crash-prone [DGFR21]. credit [TY23]. criteria [BAC22]. critical [OZ22]. Cross [SMHK21, BW22, CZR^{+}24, LWL^{+}22, ZSL^{+}23]. cross-chain [LWL^{+}22]. cross-data [BW22]. Cross-domain [SMHK21]. cross-layer [ZSL^{+}23]. cross-section [CZR^{+}24]. crossed [GSM22]. crowds [DT21]. crowdsensing [TNM^{+}22, WYW^{+}20]. crowdsensing-based [TNM^{+}22]. crowdsourced [HNN^{+}20]. crowdsourcing [MRC21]. Crypto [KVMR23]. CS [GMS^{+}21]. CS-Materials [GMS^{+}21]. CScript [DMSB20]. CSMV [NCR23]. CSqu [NA23]. CtrlEvents [TAGEL23]. cube [LFC^{+}24]. cubes [GSM22, KBW20, LLM24, MZMM21, PCC20, WHC21, ZXY21, ZLCL24].

CUDA [BGDMF^{+}24, KN24, Stp20]. current [SSG21]. curricula [DF21]. curriculum [Ada21]. Curve [KVMR23, ZD22]. cusp [NA23]. Customer [CLMH22]. Customer-satisfaction-aware [CLMH22]. customization [HCY^{+}21]. cut [GS20]. Cyber [NVE^{+}21, HFP^{+}22, MPAS24]. cyber-attacks [MPAS24]. Cyber-physical [NVE^{+}21]. cycles [LZZJ23]. cyclic [YNI^{+}22]. cyclic-min [YNI^{+}22].

D [BHB^{+}21, CFNH24, DQH^{+}21, JP22, MCC20, NE23]. DAG [RWF^{+}21]. Data [CKS22, FWZ^{+}20, MMM22, ZBF^{+}24, AHAB23, AKY20, ALS23, AA21, AKB^{+}20, BW22, BSY24, BMK^{+}22, BkB^{+}23, BSWO23, BDRJ24, CW21a, CTGJ22, DF21, DHF23, DFL^{+}23, DLWF23, DZZ^{+}23, DMM^{+}21, FHG^{+}20, FLZ^{+}20, GF20, GLF20, GHKKL23, GVC^{+}22, GPH^{+}22, GC23, GQW^{+}21, GPK21, GGTSFD23, HSHT22, HLBZ20, JJJ21, JJ22, JSJC22, JHZ20, JZS^{+}20, JKM^{+}22, KYZ^{+}20, KHO22, KTM^{+}21, KKG^{+}23, LRV20, LZL22, LZZ^{+}22, LLGC22, LW20, LLW^{+}20, LBHW20, LYZ^{+}22b, LGZZ23, MTG^{+}24, NT20, OK21, PC21, PJV^{+}22, PPN^{+}20, QDD^{+}22, RLW^{+}24, RH20, RPM24, RAJ^{+}23, SHI22, SKB21, SAY20, SM23, SMHK21, SEM20, TY23, URM23, WLZ20, WHY^{+}21, XXZ^{+}24, YZC22, YLZ^{+}20, ZNX^{+}21, ZDZ^{+}21, ZZ22]. Data-centric [ZBF^{+}24]. data-oriented [DMM^{+}21]. data-parallel [BSWO23, GQW^{+}21, RPM24]. database [BYW^{+}22]. datacenter [LLGC22]. datacenters [AJH^{+}20, NTT^{+}23]. Dataflow [CaTZ^{+}24]. dataset [VP20]. datasets [LFRBGV^{+}21]. datatype [HCC^{+}20]. DaVinci [TlW23]. DBSCAN [WCTW22]. DCA [PXY^{+}20]. DCTs [JP22]. DDF [PLBG21]. DDMTS [TDCM21]. DDoS [DT21, HA21, HR23, KKT^{+}22]. deadline [CLMH22, HLL^{+}21, WWH^{+}21]. deadline-constrained [CLMH22, HLL^{+}21]. deadlock [MMESG^{+}21]. deadlock-free [MMESG^{+}21]. December [Ano20-29, Ano21-34, Ano22-36, Ano23a]. Decentralized [HDJ21, RGB20, SÖAOA20, DAG24, GF20, GLF20, HNKÖ21, KD21, SWM23, TSTY22]. decision [CNFMA20, MLGC^{+}21]. decoder [LLC20]. Decomposition [KBW20, AM22, CSS21, FSL^{+}21, SAY20, Zha23]. Decoupling [MTR22]. deduplication [SSG^{+}20]. Deep [CGDS20, DPSD21, LGZZ23, MLR^{+}23, YWF21, YWF23, BDSQO22, DFL^{+}23, GVI24, JJJ21, KKG^{+}23, LZS^{+}24, LJH^{+}22, MBS^{+}20, MTR22, MRB20, NVE^{+}21, NMA^{+}24, ODXX21, RLW^{+}24, RPM24, SSG24, SUD^{+}22, XLCL20, YZC22, YHKR20, ZLS23, DWWX22, Kur21]. deep-learning [MTR22]. Deep-RL [DWWX22]. defined [AB22a, HA21, LCC20, YPD^{+}20, AJH^{+}20]. degrees [AMPT23]. DeJong [JH21]. delay [HSHT22, WXZ^{+}23, ZSL^{+}23]. delay-guaranteed [HSHT22]. delayed [Mal21]. delta [BW22]. demand [CCSI21, DCM^{+}23, GHD20, PPN^{+}20, XRBT21, YSZL23]. demands [CBO^{+}23, KK21]. democracy [Pou20]. Dengue [SKS21]. dense [CIH^{+}23]. densities [MS23]. dependable [Yaz23a]. dependent [LYZ^{+}22a]. deployment [AAD^{+}20, GVC^{+}22]. depth [MKP22]. derived [HCC^{+}20]. descent [YDX^{+}22]. Design [BRK^{+}21, CML^{+}24, IKK20, AR21, BG21, CaTZ^{+}24, DMM^{+}21, KL22, MAR21, PAD22, YES22]. Designing [HMCG20]. destination [XRBT21]. detect [KKT^{+}22]. detecting [LCC20, MPAS24]. Detection [XTGJ21, Alm20, AGMG23, DFP20, FFS^{+}22, GKB^{+}20, HR23, HLS^{+}23, HLL^{+}22, JJJ21, KGPT21, KTM^{+}21, KKT^{+}22, MGE20, RKAA20, RAJ^{+}23, TA24, WLZ20, ZSL^{+}23, ZLS23, CH24]. detector [SADM24]. detectors [GKP21]. determination [BDRJ24]. deterministic [FW23]. detour [MCC20]. Developing [YSMB21]. development [HGC^{+}23]. device [CN22, TAGEL23]. devices [AGC^{+}21, HWM^{+}23, HLZ22, PYYO22, PYX^{+}22, SCZ^{+}23, SUD^{+}22, XZH^{+}22, YPD^{+}20]. DHTs [HNKÖ21]. diagnosability [LZWZ22]. diagnosis [SUD^{+}22, WLL^{+}23b]. diagonal [AA20, GXYH21]. diameter [CSS21]. diffeomorphic [BHB^{+}21]. difference [Mar20]. different [LXC^{+}22]. Differential [URC20, FHG^{+}20, WMJ^{+}20]. diffusion [BSY24, DZ24]. digital [ACCN20]. dimensional [Amm21b, TZC^{+}24, VP20]. DINSaaS [Kur21]. directed [TZZ^{+}20a, YSZL23]. direction [ZWCL21, ZZZ^{+}20]. direction-based [ZZZ^{+}20]. directional [AMM21a, CLLM23, CTGJ22]. directions [KGTK20]. disaggregated [GPH^{+}22]. disaggregation [SM23]. discovery [LCW^{+}21, LD21, VFB^{+}24, ZCD^{+}21]. discrete [BCEH23]. discretization [FPdLS^{+}21]. Discriminating [DT21]. disease [HFA20]. disjoint [Lai21, LLW^{+}20, LFC^{+}24, MZMM21, WHC21, ZLCL24]. disk [LLXX24]. dispatch [XRBT21]. dispersed [PK21b]. Dispersion [KMS22, GMMP24]. distance [GQX20, KSLN24]. Distrib [YWF23]. Distributed [BkB^{+}23, CGDS20, CSS21, CAG^{+}23, CC23, FFGEL21, GXJ^{+}24, GHD20, KGPT21, Li22b, PWL^{+}22, RSSP23, RCVA22, SSA24a, SZW^{+}22, SEM20, VP20, WWL^{+}21, YHKR20, YDX^{+}22, AJH^{+}20, BSY24, BL23, CZR^{+}24, CTKdS21, CCC23, CLZ20, CDA20, DF22, DMKFJ20, DCM^{+}23, DGA^{+}24, DMSB20, DBAC^{+}22, DPEL24, DHF23, DW23, FWZ^{+}20, FLFZTS20, FPdLS^{+}21, GHKKL23, GHNS22, GMS^{+}21, GC23, GQW^{+}21, GKP21, JYH22, KST^{+}23, KSV20b, KTM^{+}21, KKT^{+}22, Kur21, LZS^{+}24, LLD^{+}23, MD20, MCD^{+}21, MBN22, NPAM20, NE23, ODXX21, PPN^{+}20, PSU^{+}21, PGB^{+}22, RBS21, RPM24, SHI22, SK21, SLFC22, SC23, TV22, VTT^{+}22, WCTW22, WDY^{+}24, XXZ^{+}24, XBX^{+}22, ZZZG21, ZJW22, ZZ22, GDF^{+}23]. Distributed-memory [SZW^{+}22, DHF23, MBN22, SHI22, WCTW22]. distribution [BDRJ24, SDLM20, SFZ23]. distribution-based [BDRJ24]. distributions [XWL^{+}20]. divergence [LW20]. diverse [MRPH20]. divisible [MDS20]. Django [ZQL^{+}21]. DLHub [LCW^{+}21]. DNN [HGC^{+}23, LYF^{+}24, AAAS24, ZJW22]. DNN-based [LYF^{+}24]. DNS [LLXG21]. docking [VFB^{+}24]. docking-based [VFB^{+}24]. documents [BYW^{+}22]. domain [LXC^{+}22, LZZ^{+}20, MBN22, SMHK21, SUD^{+}22]. domains [ARL20]. domination [GSMÖ23]. double [LLYZ23, PXY^{+}20]. DPC [CFRGPM22]. DPDK [BPBD23]. DPDK-based [BPBD23]. DQN [LZGL22]. DQPFS [SEM20]. DQS [SSG24]. Dragonflies [MMESG^{+}21]. drift [BDRJ24]. driven [BDRJ24, CCC23, HFA20, KPS^{+}22, LJH^{+}22, MBM^{+}20, SSG24, ZJW22, ZGTM24]. Driver [AB22b]. DriverRep [AB22b]. driving [AB22b]. DRL [WXZ^{+}23]. drug [VFB^{+}24]. dual [MMA22, PXY^{+}20]. DuMato [FDT^{+}24]. Dumont [CBO^{+}23]. duplicate [OS20]. duplicate-free [OS20]. duplication [TZZ^{+}20a]. durable [MLB21]. during [EHH^{+}23]. DVFS [NMPS20]. Dynamic [ALS23, AM22, DG22, GSV21, PS22, TZZ^{+}20b, CGK20, GPK21, JYH22, KYZ^{+}20, LLM^{+}20, LZGL22, LWHF22, LLX^{+}23, MPR^{+}21, TA24, TYOC24, TDCM21, WYH^{+}21, WXAL22, YY22, KK21]. dynamic-multithreading [WXAL22]. dynamics [BNOS21, SCZ24].

Early [BADP22, GMS^{+}21, HFA20]. early-stage [HFA20]. EASS [KRK20]. easy [CGG^{+}23]. EasyPAP [LNW21]. Edge [FQL^{+}23, GDF^{+}23, LLR^{+}21, MLR^{+}23, RCVA22, WMJ^{+}20, XLL^{+}20, Ala24, AMPT23, BJ23, CSS21, CJZ^{+}22, CDY23, CN22, CH23, DWWX22, DAG24, DFL^{+}23, DK24, GVC^{+}22, GPC23, GQX20, HWM^{+}23, HGC^{+}23, HLK^{+}22, HFA20, HLS^{+}23, HZL^{+}20, HLZ22, HLX^{+}22, Jar20, KPS^{+}22, LSZL20, LCH^{+}21, LZGL22, LLX^{+}23, LYF^{+}24, Li24, LLYZ23, LJH^{+}22, LYZ^{+}22b, LYC23, MZMM21, MB21, PJV^{+}22, PWL^{+}22, RSSP23, SLSN24, SWM23, SÖAOA20, SUD^{+}22, TDL^{+}22, TYM^{+}22, WCR^{+}20, WYZ^{+}24, WCT^{+}22, WCG^{+}24, XPW^{+}22, YSZL23, YZM23, YY22, YLL21, YLZ^{+}20, ZXY21, ZZZ^{+}23, ZLLC22, ZZLM22, ZLCL24, Alm20]. Edge-based [WMJ^{+}20]. Edge-Cloud [XLL^{+}20, DFL^{+}23, LSZL20, LCH^{+}21, LYF^{+}24, WCT^{+}22, YSZL23, YLZ^{+}20]. edge-connectivity [YZM23, ZXY21]. Edge-of-Things [Alm20]. Edge-to-Cloud [RCVA22]. EdgeDecAp [SWM23]. EdgeKV [SÖAOA20]. edges [LZZJ23, LLM24]. Editorial [CL22b, MS20, RSV21, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano20m, Ano20n, Ano20o, Ano20p, Ano20q, Ano20r, Ano20s, Ano20t, Ano20u, Ano20v, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano21n, Ano21o, Ano21p, Ano21q, Ano21r, Ano21s, Ano21t, Ano21u, Ano21v, Ano21w, Ano21x, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano22u, Ano22v, Ano22w, Ano22x, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano23n, Ano23o, Ano23p]. Editorial [Ano23q, Ano23r, Ano23s, Ano23t, Ano23u, Ano23v, Ano23w, Ano23x, Ano23y, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k]. Edward [KVMR23]. effect [JH21]. Effective [LSZL20, NPY^{+}23, AAD^{+}20, HLX^{+}22, KYZ^{+}20, TB22, WCR^{+}20, XLCL20]. Effectively [AT24]. effectiveness [RK24]. Efficiency [SSA24a, dRBB21, AR21, FQL^{+}23, KSLN24, LLC20, SSA24b, Sch24]. Efficient [AZF^{+}24, AMP20, BDSQO22, BMK^{+}22, BG21, CYWL21, DGMS20, FW23, GQW^{+}21, JZWX20, KRK20, LFJ^{+}20, LSC22, MLGC^{+}21, QZW^{+}24, RWF^{+}21, SSG^{+}20, SNT^{+}20, URS21, YM21, ZT20, AB22a, CHJ^{+}24, DWWX22, DGA^{+}24, DT21, DED^{+}20, DZZ^{+}23, FLZ^{+}20, FDT^{+}24, GLY^{+}21, HGC^{+}23, HLBZ20, HZL^{+}20, HMCG20, IRA20, Jar20, JL23, KL22, KYZ^{+}20, LMG^{+}21, Li23, LYF^{+}24, LYC23, LWD^{+}20, MRPH20, PAD22, PB20, RK24, RGESG^{+}21, SMMG20, SNSK20, SKS21, SBSB20, TA24, TDL^{+}22, TTD24, TAGEL23, WZO^{+}21, XBX^{+}22, XZPL24, YQZ^{+}20, ZZS^{+}21a, ZTKL^{+}21, ZZZ^{+}20]. eHealth [HSX^{+}21]. EHRs [HSX^{+}21]. Elastic [KG20, BW22, DWR^{+}23]. elasticity [FPGLSA24]. Election [ADD^{+}20, BT20]. electricity [JJJ21]. electron [SKK21]. electron-hole [SKK21]. electronic [LQX^{+}20]. Elementary [KW20]. elephant [LCC20]. elevating [DAG24]. elliptic [ZD22]. Embedded [EMSEMM20, AAAS24, SLSN24]. embedding [LLX^{+}23, LZZJ23, PW21]. embeddings [AB22b]. emergency [KK21]. empirical [HDJ21, KLL^{+}21]. empowered [CJZ^{+}22, TTD24, WLZ20]. Empowering [MRC21]. Emulous [SS21]. En-ABC [GKB^{+}20]. enabled [AKY20, Ala24, BMC^{+}24, CGC21, KKT^{+}22, KKG^{+}23, MLTT20, NVE^{+}21, SCZ^{+}23, WLZ20, HZY^{+}21]. Enabling [BGA^{+}21, Jar20, NZ23, PLBG21, SLSN24]. enclosures [LLXX24]. Encoded [KSV20b]. encrypted [BKL^{+}20]. encryption [AKY20]. end [OGR^{+}24, WYZ^{+}24]. Energy [BGA^{+}21, HWM^{+}23, JKK^{+}23, KHO22, LZGL22, LYF^{+}24, MZZW21, SKS21, WZO^{+}21, WHL^{+}23, XZY^{+}23, XZPL24, BCEH23, BSWO23, DED^{+}20, DPEL24, GBC^{+}22, GLY^{+}21, HSHT22, HKTG20, HLBZ20, IRA20, Jar20, KL22, KJA^{+}22, KSS23, KGTK20, Li24, LCC20, LLW^{+}20, LYC23, MBM^{+}20, MDS20, MBRR24, NMPS20, PB20, RK24, SMMG20, SPP^{+}23, SFML21, SBB21, SATJ^{+}20, SBSB20, WXZ^{+}23, XB20, YZC22, Yaz23a, ZZ22, ZLLC22, KRK20, RKAA20]. Energy-aware [JKK^{+}23, WHL^{+}23, HKTG20]. Energy-efficient [LYF^{+}24, XZPL24, HLBZ20, KL22, LYC23, RK24, SBSB20]. Energy-latency [LZGL22]. Energy-saving [XZY^{+}23]. enflame [CaTZ^{+}24]. enforced [HA21]. enforcement [OMCW23]. engine [RBS21]. engineering [CH24]. engines [MMESG^{+}21]. enhanced [AGMG23, GMLW24, HNN^{+}20, HZL^{+}20, AAAS24, TK23, TZC^{+}24, WYA^{+}21]. enhancement [LLM^{+}20]. Enhancing [RH20, LWW^{+}23]. ensemble [GKB^{+}20, ZWCL21]. ensembles [MLGC^{+}21]. Ensuring [NA23]. enumeration [FDT^{+}24]. Environment [XLL^{+}20, AKS^{+}20, BDRJ24, DVS24, GKB^{+}20, HA21, HFA20, HZY^{+}21, IRA20, IRLN23, LSZL20, LCH^{+}21, Li22b, RKAA20, TZZ^{+}20b, TB22, ZZZ^{+}23, ZDZ^{+}21, ZFL^{+}23]. environments [BS24, BPT^{+}22, CDY23, DFP20, DED^{+}20, EBV22, GA21, HR23, HSX^{+}21, JKK^{+}23, KSS23, Li23, MSRB20, MBB22, NPAM20, SSG^{+}20, XXZ^{+}24, XB20]. epidemic [BMC^{+}24, SCZ24]. epidemic-like [SCZ24]. epoch [PYYO22]. equation [SOS^{+}24]. equilibrium [MZZW21]. equipments [DLL^{+}21]. erasure [BW22, CPZ^{+}20, MRPH20]. erasure-coded [BW22, CPZ^{+}20]. Error [FF22, RPM24]. Error-sensitive [FF22]. errors [KGP^{+}21]. ESDU [BW22]. ESMA [DAG24]. Estimating [NPAM20, MS23]. estimation [BDRJ24, CaTZ^{+}24, CRS22, EHH^{+}23, Sok21]. Evaluating [ACC^{+}23, DDC^{+}24, RK24, XMJG22]. Evaluation [CFRGPM22, EMCE20, CFNH24, CGL^{+}22, CNFMA20, LLXX24, MKP22, MST24]. events [SFML21]. Eventually [GK24]. every [ACHP22]. Everything [MBM^{+}20]. evolutionary [GRZT22, LYF^{+}24, LD21]. Evolving [Ada21, CHJ^{+}24]. Ewald [HMS20]. exascale [BkB^{+}23, IOG20]. Excavating [WLL^{+}23a]. exchange [GHNS22, PW21]. exclusion [DGFR21]. exclusive [MLTT20]. execution [ACC^{+}23, CWHC22, FSL^{+}21, Mal21, OZ22, YM21]. existing [BAC22]. expected [CGW23]. Expelliarmus [SBBP20]. experience [BKY21, VGTSG^{+}21]. explicit [SOS^{+}24]. exploitation [CRS22]. Exploiting [FPGLSA24, Gow21]. exploration [SPP^{+}23]. Explore [ZDZ^{+}21]. Exploring [HLS^{+}23, MBRR24, QGP24, YLDY23]. Exponential [MZMM21, LZZJ23]. Extending [SAY20, SMMG22]. extension [MLB21]. extensions [AS20, VLCM^{+}20]. extra [GSM22, ZXY21]. extracting [LCK23]. extraction [XLCL20]. Extreme [Zha23].

fabric [DWW^{+}21, PAD22]. fabrics [SMMG20]. face [ZTKL^{+}21]. facility [PSU^{+}21]. factor [Li23, MBB22]. factorizations [CIH^{+}23]. failure [BDRJ24, GKP21, JHZ20, SADM24, ZDZ^{+}21]. failure-aware [JHZ20]. failures [DGFR21, GSV21, KKW23, RMJM21]. fair [GLC^{+}22]. FALCON [HCC^{+}20]. FALCON-X [HCC^{+}20]. family [KGPT23]. far [WLL^{+}23a]. Fargraph [WLL^{+}23a]. farmland [LWHF22]. Fast [AA21, BHB^{+}21, JSA21, LoKS24, LLXX24, PB21, WC22, CDA20, JP22, LSC22, MMA22, PF22, ZTKL^{+}21]. Fat [RGESG^{+}21, BCM23, WDL22]. Fat-Tree [RGESG^{+}21, WDL22]. Fault [DLWF23, PG20, WFS^{+}22, YZM23, AM20, GQX20, GSV21, LFC^{+}24, MBRR24, SNSK20, SUD^{+}22, TGFPRA20, ZLCL24]. Fault-tolerability [WFS^{+}22]. Fault-tolerant [DLWF23, PG20, AM20, LFC^{+}24]. faulty [LZZJ23, LLM24, WFS^{+}22, YZM23]. FCA [KST^{+}23]. FCP [Kur21]. FD [SADM24]. Feature [VJR20, BGDT22, BGDMF^{+}24, LLX^{+}23, SEM20, XLCL20]. Features [MB21, ASHO20, ARL20, Gow21]. February [Ano21-27, Ano22y, Ano23-32, Ano24l]. Federated [FFS^{+}22, GLC^{+}22, XXZ^{+}24, ARL20, CLW^{+}23a, CLW^{+}23b, DK24, GVI24, HDJ21, HLK^{+}22, KSS23, LLD^{+}23, PYYO22, UHAH^{+}24, dSSE23]. FELIDS [FFS^{+}22]. fence [CP24]. fence-free [CP24]. fertility [LWHF22]. fever [SKS21]. FFT [IOG20]. FGFL [GLC^{+}22]. field [MAL^{+}23]. file [AA21, CDA20, ZTKL^{+}21, ZTKL^{+}21]. files [ZTKL^{+}21]. filter [DKS21, URM23]. filters [CFNH24, ZZZG21]. finance [MS23]. finder [SWF^{+}22]. finding [ZAB20]. Fine [URM23, LCZL21, OK21, SBSB20, YQZ^{+}20]. Fine-grain [URM23]. fine-grained [LCZL21, OK21, SBSB20, YQZ^{+}20]. fingerprint [STG^{+}20]. Fire [VP22]. firefly [DED^{+}20, MBM^{+}20]. five [DDC^{+}24]. flame [SS21]. flash [AMP20, DT21, SLZP24]. flash-based [AMP20]. Flexible [DQH^{+}21, ZCY^{+}21]. flooding [CJZ^{+}22]. Flow [MRB20, BRK^{+}21, DQZZ21, FSL^{+}21, HLL^{+}22, PF22, Zha23]. flow-vector [HLL^{+}22]. flows [GVI24, LCC20, NE23]. fluid [FSL^{+}21]. FMapper [XDM^{+}22]. FMM [IOG20]. Fog [AkBA^{+}20, MSRB20, SSG21, AKS^{+}20, DMKFJ20, HNN^{+}20, HKTG20, IRLN23, Li22b, LWD^{+}20, SKS21]. folded [GQX20, GSM22, Lai21]. folding [MM24, PB20]. footprint [XB20]. forecast [HWM^{+}23]. forecasting [EBV22, YZC22]. forensics [ACCN20]. forest [ZCY^{+}21]. Forged [Kur21]. form [AA20]. formal [CC23, KST^{+}23]. formation [HLL^{+}21]. formats [ZYL^{+}21]. Forseti [ALS23]. forward [JP22, WCT^{+}22]. forwarding [FRAK23]. FP16 [CSY^{+}24]. FP16-only [CSY^{+}24]. FPC [PB21]. FPC-BI [PB21]. FPGA [HGC^{+}23, JKK^{+}23, KVMR23, MKP22, NPY^{+}23, PD21, PAD22, STG^{+}20, ZCY^{+}21]. FPGA-based [HGC^{+}23]. FPGAs [KTMB22, MKP22, MLGC^{+}21]. fractional [ALP21, SOS^{+}24]. framework [AJH^{+}20, BKL^{+}20, CLW^{+}23b, DAG24, GBEFBC24, HFA20, HZY^{+}21, KST^{+}23, LNW21, LLD^{+}23, QMB21, RK24, TK23, TTD24, XLCL20, XJR21, YZC22, YLZ^{+}20, ZSL^{+}23]. frameworks [DZZ^{+}23, RWF^{+}21]. free [AR20, AR21, CP24, HLZ22, MMESG^{+}21, MAR21, OS20]. freedom [ACHP22]. frequency [BSWO23, LKAB^{+}22, MTR22, MLTT20]. frequency-based [MLTT20]. frequent [VP20]. freshmen [DGWD21]. friendly [LSWY20]. Front [Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t]. FRQ [RWF^{+}21]. FST [BGDT22]. Fugaku [CDTS24]. Full [Ano20x, Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t]. Fully [HNKÖ21, AKY20, JKK^{+}23, VLCM^{+}20]. fully-adaptive [JKK^{+}23]. functional [ASHO20, PLBG21]. functions [Ati20, IKK20, SZW^{+}22, WDY^{+}24, ZR22]. functions-as-a-service [ZR22]. fundamental [GSA21]. funnel [NT20]. fusion [DGA^{+}24, KNK^{+}23, DGA^{+}24]. Future [GDF^{+}23, MMM22, BKY21, KGTK20]. futuristic [URC20]. fuzzy [KHO22].

GA [BS24]. GA-GWO [BS24]. game [CDY23, DWWX22, DBAC^{+}22, PJV^{+}22, XWCJ22]. game-based [DWWX22]. games [MMM21]. gaming [AMP20]. GAN [LYZ^{+}22a]. GAN-based [LYZ^{+}22a]. garbage [CLZ^{+}22]. gathering [LLW^{+}20]. Gaussian [MAL^{+}23, SSH23]. Gbps [BRK^{+}21]. GCN [WYH^{+}23]. GEMM [BDSQO22]. GEMM-based [BDSQO22]. General [JKM^{+}22, MTG^{+}24, BT20, DHF23, SSD^{+}20, SMT22]. General-purpose [JKM^{+}22, MTG^{+}24]. generalized [SZW^{+}22]. Generalizing [CW21b]. generation [CIH^{+}23, EHH^{+}23, GRZT22, HLL^{+}22, LYZ^{+}22a, Rub22]. generative [XXZ^{+}24]. generator [PD21, Stp20]. generic [GBEFBC24, GQW^{+}21]. genetic [AMM21a, HWR^{+}24, PC21]. genetic-based [AMM21a]. genomics [JKM^{+}22]. geo [Kur21, LRV20, MRPH20]. geo-distributed [Kur21]. geo-diverse [MRPH20]. geo-replicated [LRV20]. global [AMPT23, GK21, KMS22]. Goal [KPS^{+}22]. Goal-driven [KPS^{+}22]. gossip [HDJ21]. gossiping [PG20]. governor [GLC^{+}22]. GPGPU [MTR22]. GPGPUs [AR20, EL20]. GPU [ACC^{+}23, BNOS21, BMK^{+}22, BG21, BHB^{+}21, CFNH24, CMR20, FFGEL21, FDT^{+}24, GXYH21, GJL^{+}24, GM21, Gow21, HCC^{+}20, JL23, JKM^{+}22, LW20, LGZZ23, LFRBGV^{+}21, MN24, NRdA^{+}20, NZ23, SBL20, SSD^{+}20, SM23, SFT^{+}21, SNT^{+}20, VP22, VFB^{+}24, WYH^{+}23, WCMS24, XJR21, YNI^{+}22, ZDD^{+}24, ZKL21]. GPU-accelerated [MN24, YNI^{+}22]. GPU-based [JL23]. GPU-optimized [VFB^{+}24]. GPU-parallelization [BG21]. GPUs [BLB^{+}20, CLLM23, CSS^{+}23, DSZ^{+}21, DEFQO21, EHH^{+}23, FHN^{+}22, KNK^{+}23, LCK23, LSC22, NCR23, RCX^{+}21, TMB^{+}21, TA24, WZC^{+}20, YM21]. Gradient [TV22, DGA^{+}24, MMA22, YDX^{+}22, TSTY22]. grain [URM23]. grained [LCZL21, OK21, SBSB20, YQZ^{+}20]. Graph [CLW^{+}23a, HNKÖ21, MST24, WLZ20, YM21, ZZS^{+}21a, CSS^{+}23, DW23, JSA21, JSJC22, LoKS24, LTBY20, SKB21, SZQ^{+}23, WLL^{+}23a, WCMS24, XMJG22, XWJ^{+}23, YSZL23, ZJW^{+}21, ZKL21]. Graph-based [CLW^{+}23a, HNKÖ21, JSJC22]. graph-tensor [ZKL21]. Graph-Waving [YM21]. GraphCS [CLW^{+}23a]. graphics [IKK20, LoKS24]. graphs [AMPT23, CHJ^{+}24, FLFZTS20, RBS21, RSSP23, TZZ^{+}20a, THPM22, WWL^{+}21, YZM23]. graybox [CDTS24]. greedy [QZW^{+}24]. greedy-memetic [QZW^{+}24]. grey [KuR24]. grey-wolf [KuR24]. Grid [ARL20, KRK20, Uhl20, GS20, GLY^{+}21, KSV^{+}20a, Li22a, MM24, NE23, QMB21, TLD^{+}23]. grids [JJJ21, KJA^{+}22, Li22a, XTGJ21, XJR21]. group [GPK21, KSV^{+}20a, SOS^{+}24]. grouping [CLW^{+}23b]. growth [CW20]. guaranteed [HSHT22]. guarantees [KYGG20]. GuardHealth [WLZ20]. Guest [CL22b]. Guide [CGDS20]. gVMP [SM23]. GWO [BS24].

Hadoop [CWHC22, ZTKL^{+}21]. half [ATD20]. half-complex [ATD20]. Halide [JP22]. Halin [FLFZTS20]. Hamiltonian [LZZJ23]. Hammer [SKB21]. Hamming [YZM23]. handle [JJJ21]. handling [TGFPRA20]. handover [LCZL21]. happiness [DAG24]. hardness [BL23]. hardware [ABMPL22, AZF^{+}24, BRK^{+}21, FHN^{+}22, LLG^{+}24, YES22]. hardware/software [YES22]. harvesting [LLW^{+}20]. hash [AR21, Ati20, LSC22, MAR21, WDY^{+}24, XDM^{+}22]. hashing [JKM^{+}22]. HBPB [KSLN24]. HDFS [ZTKL^{+}21]. HDOT [CMFV^{+}20]. health [LQX^{+}20, OK21]. healthcare [JJ22, KKG^{+}23, NVE^{+}21, WLZ20]. HeRAFC [DVS24]. heterogeneity [CLW^{+}23a]. heterogeneous [ALS23, BS24, BSWO23, CFRGPM22, Che23, CNFMA20, DGMS20, EL20, FFGEL21, FSL^{+}21, KD22, LTBY20, LYC23, MDS20, MTG^{+}24, OZ22, PYYO22, PSBB21, QBS21, QDD^{+}22, SMT22, TDL^{+}22, TAGEL23, WWH^{+}21, WCMS24, XXZ^{+}24, XZPL24]. Heuristic [DVS24, AB22a, HKTG20, IRA20, IMP^{+}23, Rub22, SM23]. heuristic-based [Rub22]. heuristics [KEK^{+}20, SNSK20]. hexa [AAZMS20]. hexa-cell [AAZMS20]. hexagonal [Amm21b]. hidden [WLL^{+}23b]. Hierarchical [TYOC24, CCAACS21, GBC^{+}22, GPC23, JB20, LLF^{+}20, LSC22, MDS20, PW21, RdCR^{+}24, WCMS24]. hierarchy [ACHP22]. High [BDL22, JB20, KVMR23, LLC20, RCX^{+}21, ZKL21, dARR21, AT24, CLW^{+}23b, DDC^{+}24, DQZZ21, DWR^{+}23, JL23, JHZ20, LMG^{+}21, LJW^{+}22, OGR^{+}24, PGB^{+}22, SMMG22, VP20, XWL^{+}20, YTLF22]. High-efficiency [LLC20]. high-end [OGR^{+}24]. high-order [DDC^{+}24, JL23]. High-Performance [BDL22, RCX^{+}21, CLW^{+}23b, DQZZ21, DWR^{+}23, PGB^{+}22, XWL^{+}20]. highly [KKW23, NCR23, XBX^{+}22, ZZZ^{+}23]. histograms [GC23]. history [GSMÖ23]. history-based [GSMÖ23]. HITA [KVMR23]. Hitchhiker [CGDS20]. hoc [ABB22, SLFC22]. hole [SKK21]. holistic [ZSL^{+}23]. homogeneous [TZZ^{+}20a]. homomorphic [AKY20]. homonyms [ADD^{+}20]. honeypots [Ala24]. HoneyTwin [Ala24]. hood [CBO^{+}23]. host [TAGEL23]. host-device [TAGEL23]. hosts [LLXG21]. HOTD [ZSL^{+}23]. Hotspot [WYH^{+}24]. HPC [BHP^{+}24, BMC^{+}24, CBO^{+}23, CCSI21, GHT^{+}21, HQL^{+}22, KLL^{+}21, KGP^{+}21, MKP22, MBRR24, MMM21, NPO^{+}23, ZPN^{+}21, ZJW22]. HPC-enabled [BMC^{+}24]. HPL [CSY^{+}24]. HPL-AI [CSY^{+}24]. HSI [DKS21]. hull [CFNH24]. HY [WCTW22]. HY-DBSCAN [WCTW22]. Hybrid [Gow21, KN24, KSB^{+}20, Li22a, Rub22, BS24, CCC23, CMFV^{+}20, GMA^{+}22, HKTG20, HWR^{+}24, IRA20, LYF^{+}24, MZZW21, MFAB23, QM21, RLW^{+}24, RKAA20, SCZ^{+}23, WWH^{+}21, WCTW22, ZZZG21, ZGTM24, CH24, PXY^{+}20]. Hybrid-DCA [PXY^{+}20]. Hybridization [DED^{+}20, KK22]. hydraulic [SFZ23]. hydrophilic [WC22]. hydrophobic [WC22]. hydrophobic-hydrophilic [WC22]. HyLAC [KN24]. hyper [AAZMS20]. hypercube [LZWZ22, PW21, WFS^{+}22]. hypercube-based [LZWZ22]. hypercubes [GQX20, GMLW24, Lai21, LZZJ23]. hyperledger [DWW^{+}21]. hyperspectral [FFGEL21]. HyperTP [NTT^{+}23]. hypervisor [NTT^{+}23].

I/O [BHP^{+}24, CBO^{+}23, KLL^{+}21, XWL^{+}20, ZCD^{+}21]. IaaS [IMP^{+}23, SBBP20]. IAP [WYH^{+}23]. IAP-SpTV [WYH^{+}23]. ICN [DPSD21]. identification [AB22b, DZ24, STG^{+}20]. Identifying [HQL^{+}22, LLXG21]. identity [UHAH^{+}24]. IDSs [XTGJ21]. III [Li22a]. IIoT [BPT^{+}22, LJH^{+}22, WXZ^{+}23]. Image [WYA^{+}21, BMK^{+}22, BHB^{+}21, FFGEL21, KTM^{+}21, RAJ^{+}23, SBBP20, SBB20]. imagery [LD21]. images [DFP20]. imbalanced [JJJ21]. Impact [GVI24, BMC^{+}24, LXC^{+}22, MHdC^{+}23]. Implementation [GMA^{+}22, CML^{+}24, DMM^{+}21, HMS20, IKK20, JP22, LFJ^{+}20, MAR21, SKK21, SBB20, TA24, VP22]. implemented [VLCM^{+}20]. Implementing [GHNS22]. implicit [NE23]. improve [CLZ^{+}22, DGMS20, GMS^{+}21, KSLN24, RKAA20]. Improved [CMR20, DED^{+}20, SK21, FWZ^{+}20, VJR20, XJR21]. Improvement [YWF21, YWF23, FQL^{+}23, KK22]. Improving [BJ23, LKAB^{+}22, MFAB23, SFML21, SPBR20, SC23, SBB21, ZdCLT22, LSZL20]. In-depth [MKP22]. in-memory [AZF^{+}24, HQL^{+}22, YES22]. incentive [GLC^{+}22]. incompressible [MBN22]. increased [THPM22]. independency [CW21b]. independent [BCEH23, KKH^{+}23, Li24]. index [PK21b, XDM^{+}22]. indexing [OGR^{+}24, TDL^{+}22]. individualized [Li22b]. industrial [HLK^{+}22, KTM^{+}21, SMHK21, SUD^{+}22, ZNX^{+}21]. infection [SKS21]. inference [HGC^{+}23, JSJC22, YSZL23, ZWCL21]. InfiniBand [MMESG^{+}21]. Information [GHKKL23, DGFR21, DZ24, GK21, HWM^{+}23, LZZ^{+}22]. information-coordinated [GK21]. Information-theoretic [GHKKL23]. Infrastructures [PB21]. inherent [FPGLSA24]. input [WYH^{+}23]. input-aware [WYH^{+}23]. insertion [PAD22]. Insight [Kur21]. inspired [DPSD21, GMA^{+}22, HWR^{+}24, SATJ^{+}20]. instantiation [CSY^{+}24]. instruction [LLF^{+}20]. Instructions [KW20]. integer [XZY^{+}23]. Integrate [VP22]. Integrate-and-Fire [VP22]. integrated [MSRB20, TYM^{+}22]. Integrating [DF21, KNK^{+}23, KJA^{+}22]. Integrity [AA21, WHY^{+}21, YLZ^{+}20]. Intel [AS20, CFRGPM22, MKP22, TGFPRA20]. Intelligence [GDF^{+}23, dARR21, BJ23, CJZ^{+}22, Kur21, RCVA22, SATJ^{+}20, ZLLC22]. Intelligent [CN22, LYC23, ZPN^{+}21, LYZ^{+}22b, PJV^{+}22, Rub22, SCZ^{+}23, TDL^{+}22, XBX^{+}22]. Intelligently [LCC20]. intensive [FQL^{+}23]. Inter [PSU^{+}21, FRAK23, LZZ^{+}20, NPY^{+}23, SMMG20]. inter-domain [LZZ^{+}20]. inter-FPGA [NPY^{+}23]. Inter-kernel [PSU^{+}21]. inter-thread [FRAK23]. interacting [BLNP23]. interaction [LLF^{+}20]. interactions [VGTSG^{+}21]. Interactive [HR23, MCD^{+}21, MGW24]. interconnect [KGP^{+}21, MST24]. interconnecting [AJH^{+}20]. interconnection [SMMG20]. interest [TYOC24]. Interference [LZS^{+}24, PWL^{+}22]. Interference-aware [LZS^{+}24]. Interlaced [HNKÖ21]. Intermediate [LZL22]. internally [WHC21]. Internet [GDF^{+}23, AGMG23, CJZ^{+}22, FTA^{+}22, FFS^{+}22, GLF20, JZWX20, MBM^{+}20, RKAA20, XLL^{+}21, XBX^{+}22, YS21]. interoperability [LWW^{+}23]. Interval [RSGA20]. intra [Zha23]. intra-session [Zha23]. introducing [DGWD21]. introduction [QBS21]. Intrusion [Alm20, AGMG23, FFS^{+}22, HLS^{+}23, KKT^{+}22, RKAA20]. inverse [JP22]. inversion [CIH^{+}23]. Investigating [BMC^{+}24]. investments [DBAC^{+}22]. IoE [SCZ^{+}23]. Ion [CZR^{+}24]. Ion-molecule [CZR^{+}24]. IoT [AGC^{+}21, BDFG21, CN22, GKSS24, HLS^{+}23, HLX^{+}22, KKT^{+}22, KKG^{+}23, Li23, MLTT20, MBB22, MGE20, PJV^{+}22, PWL^{+}22, PYX^{+}22, SMHK21, SKS21, SUD^{+}22, TDL^{+}22, TTD24, VD21, WHY^{+}21, XLL^{+}20, YZC22, Yaz23a, ZWS^{+}20, dSSE23]. IoT-assisted [MLTT20]. IoT-enabled [KKG^{+}23]. IoT-Fog [SKS21]. IoTA [VD21]. IoTSim [AJH^{+}20]. IoTSim-SDWAN [AJH^{+}20]. IPFS [JJ22, KTM^{+}21]. irregular [BMK^{+}22, FPGLSA24, TLC20]. ISA [SBB21]. Issue [Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, BDL22, GDF^{+}23, MS20, MMM22, dARR21]. issues [Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, JTV^{+}22]. itemset [VP20]. iterations [DG22]. iterative [BLB^{+}20, KNK^{+}23, MFAB23, Sok21, SOS^{+}24, XMJG22].

J [YWF23]. jamming [LSWY20]. January [Ano21-35, Ano22-35, Ano23-33, Ano24v]. jet [MD20]. JMDC [DFL^{+}23]. job [BCM^{+}21, LZS^{+}24, PC21, TLD^{+}23, YLDY23, ZdCLT22]. join [Gow21]. Joint [CW21a, SWY^{+}21, DFL^{+}23, SM23, WXZ^{+}23]. Joker [KG20]. Journal [SSA24a]. Julia [HMCG20]. July [Ano20-27, Ano21-28, Ano22-29, Ano23-35, Ano24w]. June [Ano20y, Ano21-30, Ano22-31, Ano23-29, Ano24x].

KASLR [VGMG20]. KASLR-MT [VGMG20]. Keeping [PGB^{+}22]. Kernel [VGMG20, CLLM23, PSU^{+}21, SWF^{+}22, WCLD21, ZDD^{+}24]. kernels [ACC^{+}23, GJL^{+}24, NZ23, SZW^{+}22]. key [JZWX20, Rub22, XLL^{+}21]. keyword [YQZ^{+}20]. knapsack [WYH^{+}24]. KNN [Gow21]. KNN-join [Gow21]. Knowledge [HFA20, AKB^{+}20, LoKS24, LD21, NZ23]. Knowledge-driven [HFA20]. knows [ESA24]. Kokkos [DMM^{+}21, HMS20]. Kunpeng [CSY^{+}24].

labeling [BDRJ24, FF22, Mal21]. Lagrangian [Uhl20]. LakeHouse [AHAB23]. Lakes [AHAB23]. lane [dRBB21]. language [DMSB20, Stp20]. language-based [Stp20]. Languages [MS20]. Large [CDTS24, BYW^{+}22, BLB^{+}20, CC23, HQL^{+}22, KGP^{+}21, LLXX24, MAL^{+}23, PK21a, SBL20, STG^{+}20, YZM23]. Large-scale [CDTS24, MAL^{+}23, PK21a, SBL20, STG^{+}20, YZM23]. LASs [VLCM^{+}20]. latency [GKTW21, LZGL22, ZR22]. latest [NE23]. lattice [GK24]. lattice-linear [GK24]. layer [CaTZ^{+}24, DGA^{+}24, XTGJ21, ZSL^{+}23]. layer-wise [CaTZ^{+}24]. Layered [SW22]. Layout [VGMG20]. lazy [PMV20]. leader [BT20]. Leaderless [ABD^{+}23, RdCR^{+}24]. leaf [CW21a]. leaf-spine [CW21a]. Leaky [VP22]. Learning [CGK20, GLC^{+}22, MLR^{+}23, SSA24a, YWF21, YWF23, ZGTM24, Ala24, ACC^{+}23, BSY24, CLW^{+}23a, CLW^{+}23b, CWHC22, DK24, DW23, DEFQO21, DPSD21, FTA^{+}22, FFS^{+}22, GJL^{+}24, GBEFBC24, GVI24, GLL21, HDJ21, HLK^{+}22, HFA20, HLX^{+}22, JH21, KK22, KKG^{+}23, LNW21, LCW^{+}21, LZS^{+}24, LLYZ23, LJH^{+}22, LGZZ23, MTR22, MRB20, MMM21, NGS21, NMA^{+}24, PS22, PYYO22, QMB21, RLW^{+}24, RPM24, RAJ^{+}23, SSG24, SSA24b, SPP^{+}23, SSH23, SWF^{+}22, UHAH^{+}24, VGTSG^{+}21, WCR^{+}20, XXZ^{+}24, YZC22, YHKR20, YSMB21, ZKL21, ZLS23, ZDL^{+}24, ZDZ^{+}21, dSSE23]. Learning-based [CGK20, FFS^{+}22, LJH^{+}22, ZLS23]. Learning-driven [ZGTM24]. learning-enabled [Ala24]. least [PG20]. Ledgers [CAG^{+}23]. Letting [BKY21]. Level [IKK20, ALS23, ACHP22, CSS^{+}23, JB20, MGW24, RMJM21, SLZP24, VGTSG^{+}21, WZC^{+}20, Yaz23b]. levels [WYZ^{+}24]. Leveraging [CCSI21, MMESG^{+}21, JB20]. LFIB4 [Stp20]. library [BGDT22, VLCM^{+}20, PLBG21]. lifecycle [LLG^{+}24]. Lifespan [CLZ^{+}22]. Lifespan-based [CLZ^{+}22]. lifetime [GHD20, RMJM21, YS21]. light [VTT^{+}22]. light-weight [VTT^{+}22]. Lightweight [MGE20, AZF^{+}24, CW20, DWW^{+}21, PYX^{+}22, PSU^{+}21, SKB21, SC23]. like [SCZ24]. likelihoods [MAL^{+}23]. limitations [ACR23]. limits [SAY20]. line [JTV^{+}22]. line-of-sight [JTV^{+}22]. linear [GK24, KN24, MFAB23, SSH23, TV22, VLCM^{+}20, XZY^{+}23, ZAB20]. linkage [NPAM20]. linked [PMV20]. links [YZM23]. Linux [SWF^{+}22]. List [BCEH23]. lists [PMV20]. literature [IRLN23, RCVA22]. live [NTT^{+}23, PPN^{+}20]. Load [KSB^{+}20, MBM^{+}20, AM22, BAC22, DED^{+}20, FRAK23, FPdLS^{+}21, GGTSFD23, PSBB21, SOL22, TDCM21, VP20, ZZS^{+}21b]. load-balanced [SOL22]. loads [MDS20]. Local [LLX^{+}23, DLWF23, FHG^{+}20, PYYO22, WWL^{+}21]. locality [THPM22, XWJ^{+}23]. locality-optimized [XWJ^{+}23]. localization [KGPT21, LMG^{+}21, SM22]. Locally [ACR23]. location [LLR^{+}21, NYZ^{+}20]. LOCATOR [TAG23]. lock [AR21, ACHP22, MAR21]. lock-free [AR21, MAR21]. lock-freedom [ACHP22]. logic [PD21]. long [DLL^{+}21, MD20, MYS^{+}23]. long-running [MYS^{+}23]. long-term [DLL^{+}21]. LoRaWAN [IWS22]. LOSC [XWJ^{+}23]. loss [AA21, SZW^{+}22]. lossless [DKS21]. Low [TAG23, CKS22, CNFMA20, JHZ20, MAL^{+}23, XZH^{+}22, Yaz23a]. low-cost [CKS22]. Low-power [TAG23, Yaz23a]. low-rank [MAL^{+}23]. lower [SHI22]. lower-bounds [SHI22].

MaaS [BPT^{+}22]. Machine [DEFQO21, FTA^{+}22, SSA24a, Ala24, ACC^{+}23, BADP22, DW23, GJL^{+}24, GBEFBC24, GLL21, HFA20, HLBZ20, JH21, LCW^{+}21, NGS21, SSA24b, SBBP20, SPP^{+}23, WYH^{+}24, ZDL^{+}24, ZFL^{+}23]. Machines [XLL^{+}20, ALP21, CKS22]. MAD [KGPT21]. MAD-C [KGPT21]. MADRL [WCG^{+}24]. magnetohydrodynamics [ZT20]. make [KW20]. making [CNFMA20]. MalFCS [XLCL20]. malicious [HLL^{+}22]. Malware [DFP20, RAJ^{+}23, WYH^{+}21, XLCL20]. Management [AkBA^{+}20, Kur21, CWHC22, CGK20, CL22b, GF20, GQW^{+}21, HFP^{+}22, HZY^{+}21, JZS^{+}20, KBS^{+}21, KGTK20, LSZL20, LLF^{+}20, LHL21, NMPS20, RGB20, RWF^{+}21, SBBP20, SLFC22, TK23, TTD24, TNM^{+}22, WLZ20, Yaz23a, YTLF22, ZNX^{+}21, ZYW24]. manager [SSD^{+}20]. Managing [XB20]. manipulation [HSX^{+}21]. many [BV21, MZMM21, SBSB20]. many-core [BV21, SBSB20]. many-to-many [MZMM21]. manycore [KEK^{+}20]. manycores [PSU^{+}21, SC23]. map [AR21, HA21]. mapper [XDM^{+}22]. Mapping [GBC^{+}22, EL20, Jea22, LMG^{+}21, MRB20, SLZP24, WHL^{+}23]. MapReduce [BCM^{+}21, FW23]. March [Ano21-33, Ano22-28, Ano23-34, Ano24y]. Margolus [PAD22]. Marsa [Stp20]. Marsa-LFIB4 [Stp20]. mass [SHI22]. massively [NE23]. master [HLZ22]. master-copy-free [HLZ22]. MATAR [DMM^{+}21]. matching [DCM^{+}23, MLTT20, SALP20, TYOC24]. materials [GMS^{+}21, GMS^{+}21]. matrices [ATD20, AA20]. Matrix [ATD20, CIH^{+}23, CSS^{+}23, GXYH21, RCX^{+}21, SMT22, TLC20, WCLD21]. matrix-centric [CSS^{+}23]. matrix-matrix [RCX^{+}21]. matrix-vector [GXYH21, TLC20]. matroidal [ZZ24]. Matter [Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t]. matters [CL22a]. max [CSS21, PF22]. max-flow [PF22]. maximal [CTA20, FLFZTS20]. maximization [CLMH22, LCH^{+}21, YS21, ZZ22]. May [Ano21z, Ano22-33, Ano23-31, Ano24z]. MB [BPT^{+}22]. MB-MaaS [BPT^{+}22]. MDP [LYZ^{+}22b]. MDP-based [LYZ^{+}22b]. MDScale [BNOS21]. means [FBL^{+}21, LLM^{+}20]. MEC [CN22, SCZ^{+}23, WXZ^{+}23]. MEC-based [WXZ^{+}23]. MEC-enabled [SCZ^{+}23]. mechanism [DGMS20, GLF20, HA21, IRA20, PCC20, SM22, SS21, TDL^{+}22, TYM^{+}22, WYH^{+}21, YLL21, KK21]. meet [MMM22]. meets [LZD21]. memetic [QZW^{+}24]. memories [JB20]. Memory [EMSEMM20, AZF^{+}24, BBE^{+}21, BDRJ24, DDC^{+}24, DHF23, EMCE20, GPH^{+}22, HQL^{+}22, JSA21, LLG^{+}24, LGZZ23, MDS20, MBN22, MHdC^{+}23, MAR21, NPY^{+}23, NE23, NCR23, QM21, RLW^{+}24, SHI22, SZW^{+}22, SC23, TZZ^{+}20b, TGFPRA20, TV22, TLC20, URS21, WLL^{+}23a, WCTW22, YES22, ZTKL^{+}21, ZJW^{+}21, GM21]. memory-aware [TZZ^{+}20b]. memory-efficient [ZTKL^{+}21]. Mesh [KRK20, CGL^{+}22, DSZ^{+}21, DQH^{+}21, MRB20]. mesh-based [DSZ^{+}21, MRB20]. mesh-connected [DQH^{+}21]. message [DCM^{+}23, LLW^{+}20]. meta [HKTG20, IRA20]. meta-heuristic [HKTG20, IRA20]. metadata [ZTKL^{+}21]. metaheuristics [BG21]. meteorological [HWM^{+}23]. meter [DPEL24]. method [ABB22, BW22, CCC23, CaTZ^{+}24, DGA^{+}24, DZ24, GRZT22, HR23, HLZ22, JL23, LYZ^{+}22a, LJH^{+}22, SCZ^{+}23, SCZ24, SLZP24, SOS^{+}24, XWCJ22, ZFL^{+}23]. methodology [ACCN20, YES22]. Methods [TV22, BGDMF^{+}24, Mal21, MFAB23, MAR21, ZWCL21]. metric [GXJ^{+}24]. metropolitan [LZZ^{+}20]. MI [BGDMF^{+}24]. MI-based [BGDMF^{+}24]. micro [JB20, KJA^{+}22, MD20]. micro-core [JB20]. micro-object [MD20]. microgrid [DMKFJ20]. microscopy [BMK^{+}22]. microservice [LLD^{+}23]. MiCS [LZZ^{+}22]. MiCS-P [LZZ^{+}22]. middleware [LJW^{+}22]. MIDP [LYZ^{+}22b]. Migration [MLR^{+}23, AB22a, LZGL22, PC21, PPN^{+}20, WCR^{+}20, WYH^{+}24, YLL21, ZFL^{+}23]. millions [RSSP23]. min [CSS21, YNI^{+}22]. min-max [CSS21]. mini [PYYO22]. mini-batch [PYYO22]. MiniChain [CW20]. minimal [AKB^{+}20, RSSP23]. minimization [GSV21]. minimize [BCEH23, HSHT22]. minimizing [Che23]. Mining [BPT^{+}22, VP20]. Mining-as-a-Service [BPT^{+}22]. MIRAGE [PC21]. mismatch [CTA20]. Mitigating [MPR^{+}21]. Mixed [NPO^{+}23, AT24, DMSB20, GVI24, XZY^{+}23, YTLF22]. mixed-consistency [DMSB20]. mixture [XWL^{+}20]. ML [BDRJ24]. ML-driven [BDRJ24]. MLCN [dRBB21]. MLLess [SSA24a, SSA24b]. mmWave [JTV^{+}22]. Mobile [BPT^{+}22, CDY23, DFP20, DLL^{+}21, IRA20, JZWX20, KMS22, LZGL22, LLYZ23, LWD^{+}20, MRC21, SM22, TYM^{+}22, WYW^{+}20, WCR^{+}20, WCG^{+}24, ZLLC22]. mobility [CL22b, EBV22, GVC^{+}22, JTV^{+}22]. mobility-based [GVC^{+}22]. Möbius [PCC20]. mode [MHdC^{+}23]. Model [NRGL22, ACC^{+}23, BkB^{+}23, BDRJ24, DFL^{+}23, DK24, DLWF23, DW23, EBV22, HZL^{+}20, JJ22, KPS^{+}22, LCK23, Li22a, LXC^{+}22, LLD^{+}23, AAAS24, NVE^{+}21, PYX^{+}22, PLBG21, QM21, RWF^{+}21, RPM24, SKK21, SK21, SLSN24, SFML21, SCZ24, Sok21, VP22, WC22, ZJW22, ZLS23, ZLLC22, ZLCL24, HA21]. Model-based [NRGL22]. model-parallel [RPM24]. Modeling [XWL^{+}20, LLX^{+}23, LCC20, PK21a, TlW23]. Modelling [BDFG21]. models [DDC^{+}24, LCK23, LCW^{+}21, MS20, SSH23, SFML21, TLC20]. modern [EHH^{+}23, HCC^{+}20]. modes [WDL22]. Modified [CH24, SOS^{+}24]. modular [DF21, MD20, XZPL24]. module [QBS21]. module-based [QBS21]. moldable [KL22, PKSR24]. molecular [BNOS21, VFB^{+}24]. molecule [CZR^{+}24]. moments [JL23]. Momentum [ZJW22]. Momentum-driven [ZJW22]. Monero [LWL^{+}22]. monitoring [DMKFJ20, DPEL24]. Monotonic [Uhl20]. Monte [DPSD21]. moth [SS21]. moth-flame [SS21]. MPI [ACC^{+}21, CL22a, HCC^{+}20, KSB^{+}20, MBN22, NRGL22, QM21, ZT20]. MPI-3 [QM21]. MPI Sort [ZBF^{+}24]. MPSoC [NHR22]. MSHGN [WCMS24]. MSLShard [TTD24]. MT [VGMG20]. Multi [CLLM23, DED^{+}20, EMSEMM20, GM21, KK22, KGPT21, PKSR24, SCZ^{+}23, VGMG20, WCMS24, WXZ^{+}23, ZZS^{+}21b, AGMG23, BCM^{+}21, BNOS21, BMC^{+}24, CLT^{+}20, CMR20, DGA^{+}24, FBL^{+}21, GRZT22, GS20, HZL^{+}20, HMCG20, KD22, KuR24, KSS23, LMG^{+}21, Li22a, LLGC22, LQX^{+}20, LFRBGV^{+}21, MBN22, NMPS20, NCR23, NRGL22, OZ22, PD21, SS21, SM23, SOS^{+}24, TB22, TY23, TLC20, VTT^{+}22, WYA^{+}21, WLL^{+}23b, WCG^{+}24, XB20, XZPL24, Zha23, ZGTM24, dRBB21, WDL22]. multi-access [WCG^{+}24]. Multi-agent [WXZ^{+}23, BCM^{+}21, CMR20, TB22]. multi-bit [PD21]. multi-block [GS20]. multi-channel [VTT^{+}22]. Multi-Cloud [ZZS^{+}21b, XB20]. multi-cluster [BMC^{+}24]. multi-core [HMCG20, NMPS20, NRGL22, OZ22, SOS^{+}24, TLC20, XZPL24]. Multi-directional [CLLM23]. multi-domain [MBN22]. Multi-GPU [GM21, BNOS21, LFRBGV^{+}21]. multi-grid [Li22a]. multi-lane [dRBB21]. multi-layer [DGA^{+}24]. Multi-Objective [DED^{+}20, KK22, SCZ^{+}23, AGMG23, GRZT22, KuR24, KSS23, SS21, SM23]. multi-party [CLT^{+}20, FBL^{+}21]. multi-protection [LLGC22]. Multi-Rail [WDL22]. Multi-requestor [EMSEMM20]. Multi-resource [PKSR24]. multi-robot [LMG^{+}21]. multi-scale [WYA^{+}21]. Multi-scenario [WCMS24]. multi-source [LQX^{+}20, Zha23]. Multi-stage [KGPT21]. Multi-Tenant [VGMG20, HZL^{+}20]. multi-threaded [WLL^{+}23b]. multi-threading [KD22]. multi-type [ZGTM24]. multi-user [TY23]. multi-versioned [NCR23]. multi-zone [GM21]. multicast [Yaz23b, Zha23]. Multicore [KEK^{+}20, BDSQO22, BGDT22, BSWO23, LLM^{+}20, SFML21, SPK^{+}23]. multicores [SBB21]. MultiFog [DVS24]. MultiFog-Cloud [DVS24]. multigrid [IOG20]. Multilevel [VJR20, JSA21]. multimedia [SFT^{+}21]. multiphase [KGPT23]. Multiple [SALP20, BLB^{+}20, GK21, GSV21, IKK20, KKH^{+}23, Li22a, Li24, LWD^{+}20, WYZ^{+}24, YLDY23]. multiple-cell [Li22a]. multiple-precision [IKK20]. multiplication [ATD20, GXYH21, RCX^{+}21, SMT22, TLC20, WCLD21, ZD22]. multiplicity [CP24]. multiprocessor [QGP24, TZZ^{+}20a]. multiprocessors [QZW^{+}24]. multiprogrammed [WXAL22]. multitask [SBB21]. multitasking [WZC^{+}20]. multithreading [WXAL22]. multitiered [RH20]. Multiview [LLC20]. Mutation [LWHF22, GMA^{+}22]. Mutual [BCM23, ABMPL22, DGFR21, LZZ^{+}22]. mutual-information [LZZ^{+}22].

NAMD [MM24]. NAS [GM21]. Nash [MZZW21]. navigation [SLSN24]. near [CMR20]. nearest [Gow21]. needs [SZQ^{+}23]. negative [BkB^{+}23]. neighbor [Gow21, GPC23]. Neighborhood [Uhl20, PAD22]. neighbours [CMR20]. nested [LW20]. net [BLNP23, MLB21]. Nets [EMCE20, SLFC22]. Network [AJH^{+}20, WLZ20, AM20, AGMG23, BDSQO22, BV21, CGL^{+}22, DQZZ21, DZ24, DLWF23, FWZ^{+}20, HA21, HLS^{+}23, HLZ22, JJJ21, KK22, KD21, KGP^{+}21, KKT^{+}22, LLX^{+}23, LTBY20, MMA22, NVE^{+}21, ODXX21, QDD^{+}22, RK24, SCZ24, SALP20, SOL22, WZO^{+}21, WYA^{+}21, WCMS24, Yaz23b, ZWS^{+}20, Zha23, ZLLC22, ZDL^{+}24]. network-based [HA21]. network-on [KK22]. network-on-chip [BV21, RK24, SOL22, Yaz23b]. network-on-chips [AM20, ZDL^{+}24]. Networked [CTGJ22]. networking [YPD^{+}20]. Networks [AS21, ABB22, CGDS20, JHML21, KRK20, LAPB20, RGB20, ZZS^{+}21a, AMM21a, AB22a, AKB^{+}20, Amm21b, BSY24, BT20, BDRJ24, Che23, DF22, DFL^{+}23, GHD20, GA21, GPC23, JYH22, JTV^{+}22, KKH^{+}23, KYZ^{+}20, KTMB22, LSWY20, LCZL21, LLGC22, LLX^{+}23, LXC^{+}22, LLW^{+}20, LZZ^{+}20, LZWZ22, LFC^{+}24, MZZW21, NPY^{+}23, OMCW23, PW21, PG20, SM22, SAATK21, SADM24, SWY^{+}21, SATJ^{+}20, SFZ23, WHY^{+}21, XLCL20, YZC22, YSZL23, ZSL^{+}23, ZYW24, ZZ24, ZZLM22, dRBB21]. Networks-on-Chip [JHML21]. Neural [CGDS20, JHML21, ZZS^{+}21a, AGMG23, BDSQO22, BDRJ24, CGL^{+}22, DFL^{+}23, GA21, HLZ22, KTMB22, LXC^{+}22, LTBY20, ODXX21, SCZ24, WZO^{+}21, XLCL20, YZC22, YSZL23, ZLLC22]. neuron [VP22]. neutral [MS23]. new-generation [CIH^{+}23]. newborn [TZDC21]. next [MGW24]. NLSTM [YZC22]. NoC [NHR22, PSU^{+}21, QZW^{+}24]. NoC-based [NHR22, PSU^{+}21, QZW^{+}24]. NoCs [MCC20]. node [Lai21, LLX^{+}23, MKP22, WDL22]. node-disjoint [Lai21]. noisy [BSY24]. Non [CDY23, ASHO20, Ati20, BkB^{+}23, CCAACS21, FSL^{+}21, MAL^{+}23, OMCW23, SSH23]. non-asymptotic [Ati20]. Non-cooperative [CDY23]. non-functional [ASHO20]. non-Gaussian [MAL^{+}23, SSH23]. non-hierarchical [CCAACS21]. non-negative [BkB^{+}23]. non-SDN [OMCW23]. non-uniform [FSL^{+}21]. nonblocking [DHF23]. nonlinear [SLSN24]. NoSQL [BYW^{+}22]. note [Mal21]. Novel [BMC^{+}24, LZZJ23, AR21, CSY^{+}24, CGL^{+}22, GBEFBC24, HLL^{+}22, LZZ^{+}20, QM21, SLZP24, TK23, TDCM21, TYM^{+}22, XZH^{+}22, Yaz23a, ZZS^{+}21b, ZZZG21, ZCY^{+}21]. November [Ano20-30, Ano21-36, Ano22-34, Ano23z]. NPUs [SPK^{+}23]. NUMA [CIH^{+}23, EMCE20, THPM22]. number [Mal22, PD21, Stp20]. numbers [GMLW24]. numerical [Sok21]. NVIDIA [TMB^{+}21].

O [PB20, BHP^{+}24, CBO^{+}23, KLL^{+}21, XWL^{+}20, ZCD^{+}21]. object [CTFW22, CML^{+}24, DGWD21, ESA24, MD20]. object-based [DGWD21]. Objective [DED^{+}20, AGMG23, GRZT22, KK22, KuR24, KSS23, LWHF22, SCZ^{+}23, SS21, SM23]. objective-based [LWHF22]. oblivious [CLT^{+}20, SAY20, WCT^{+}22]. Observer [WYH^{+}21]. obstacle [KGPT21]. October [Ano20-31, Ano21-31, Ano22-27, Ano23-27]. off [JJ22]. off-chain [JJ22]. offline [IMP^{+}23]. offloading [CDY23, CN22, CH23, LMG^{+}21, LCH^{+}21, Li22b, LYF^{+}24, LLYZ23, TYM^{+}22, WCG^{+}24]. offs [MDS20]. omics [SHI22]. omnidirectional [BCM23]. OmpSs [VLCM^{+}20]. On-demand [YSZL23, GHD20]. On-GPU [LW20]. one [CDA20]. Online [HSHT22, LLYZ23, RGB20, SFT^{+}21, ZDZ^{+}21, ZdCLT22, TSTY22]. only [CSY^{+}24, ZWSL22]. onsensus [CGG^{+}23]. ontology [ASHO20]. open [FTA^{+}22]. OpenACC [ACC^{+}21, Stp20, XJR21]. OpenCL [MKP22]. OpenFlow [LSC22]. OpenMP [ACC^{+}21, KSB^{+}20, SBB20, Stp20, VLCM^{+}20, WLL^{+}23b, ZT20]. operating [PSU^{+}21, SC23]. operation [DMKFJ20]. operations [GHNS22, TAGEL23, ZKL21]. operator [CLLM23, CW21b, GMA^{+}22, KNK^{+}23]. operators [BDSQO22]. opportunistic [AKB^{+}20, LZS^{+}24]. Opportunities [TMB^{+}21, HQL^{+}22, IOG20, LZD21, MBRR24]. Oppositional [PJV^{+}22]. optical [AM20]. Optically [GPH^{+}22]. Optimal [BAC22, CGG^{+}23, IRLN23, Lai21, MRPH20, OS20, DEFQO21, NRGL22, PW21, TZZ^{+}20a, YS21, CH24]. optimally [GQX20]. Optimistic [KYGG20]. Optimization [DED^{+}20, TMB^{+}21, ZD22, BS24, CDTS24, CaTZ^{+}24, CL22a, DVS24, DLL^{+}21, DZ24, GK21, GHD20, KK22, KuR24, Li22b, LJH^{+}22, MD20, MMA22, NT20, ODXX21, PJV^{+}22, PWL^{+}22, RK24, SCZ^{+}23, SSG24, SS21, SCZ24, SWY^{+}21, SPK^{+}23, Stp20, XZY^{+}23, YNI^{+}22, YPD^{+}20, YHKR20, ZYL^{+}21, ZFL^{+}23, TSTY22]. optimizations [CSS^{+}23, ZCD^{+}21]. Optimized [XPW^{+}22, PD21, TB22, VFB^{+}24, XWJ^{+}23]. optimizer [KSS23, ZDD^{+}24]. Optimizing [AAAS24, SPP^{+}23, CWHC22, SWM23, dSSE23]. option [MS23]. optoelectronic [AAZMS20]. ORB [TAG23]. orchestrated [KKG^{+}23]. orchestration [YY22]. order [CW21b, DDC^{+}24, JL23]. order-independency [CW21b]. ordering [GSMÖ23]. organic [KG20]. Organizing [HA21]. orientation [CSS21]. oriented [DMM^{+}21, XZH^{+}22]. OTIS [AAZMS20]. out-of-core [XWJ^{+}23]. output [BSY24]. outsourced [FLZ^{+}20, OK21]. outsourcing [JSJC22, TY23, WCT^{+}22]. overlapped [AA20]. overlapping [TAGEL23].

P [LZZ^{+}22]. P2P [MZZW21]. PAASH [OK21]. packet [LSC22]. Packing [WHC21, FLFZTS20]. PackStealLB [FPdLS^{+}21]. Page [Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, SLZP24, TGFPRA20]. page-fault [TGFPRA20]. page-level [SLZP24]. Pages [Ano20z, Ano20-29, Ano20y, Ano20-27, Ano20-30, Ano20-31, Ano20-28, Ano21-32, Ano21-29, Ano21-34, Ano21-27, Ano21-35, Ano21-30, Ano21-28, Ano21-33, Ano21z, Ano21-36, Ano21-31, Ano21y, Ano22z, Ano22-30, Ano22-36, Ano22y, Ano22-35, Ano22-31, Ano22-29, Ano22-28, Ano22-33, Ano22-34, Ano22-27, Ano22-32, Ano23-30, Ano23-28, Ano23-32, Ano23-33, Ano23-29, Ano23-35, Ano23-34, Ano23-31]. PAGroup [CLW^{+}23b]. pair [CTA20]. Paired [ZLCL24]. pancyclic [KBW20]. paradigm [HZL^{+}20, SKS21]. ParaLiNGAM [SSH23]. Parallel [AAZMS20, AMPT23, DKS21, GK21, Gow21, KSB^{+}20, MS23, PB20, SKK21, SSA24a, SSH23, SBB20, VJR20, YWF23, ZAB20, ZWCL21, Ati20, BADP22, BCEH23, BSWO23, CDTS24, CTKdS21, CIH^{+}23, Che23, CLZ20, CTA20, DF22, GHKKL23, GSA21, GMS^{+}21, GSMÖ23, GM21, GBC^{+}22, GQW^{+}21, HMCG20, JH21, JP22, KKH^{+}23, KGPT23, LLM^{+}20, LNW21, LKAB^{+}22, MM24, Mal22, MS20, Mar20, NE23, NMA^{+}24, PF22, PGB^{+}22, QDD^{+}22, RPM24, ŠK23, Sok21, SBSB20, SFZ23, SOS^{+}24, TLD^{+}23, TYOC24, VGTSG^{+}21, WCLD21, WCTW22, YSMB21, ZZZG21, ZLS23, ZZZ^{+}20, ZT20, BGDT22, LZZ^{+}22]. Parallel-FST [BGDT22]. parallel/distributed [ZZZG21]. parallelisation [Li22a]. parallelism [LLC20, MFAB23, AAAS24, WLL^{+}23a]. parallelizable [FW23]. Parallelization [YWF21, YWF23, ACC^{+}21, BG21, CZR^{+}24, CW21b, KEK^{+}20, MBN22, Sch24, ZT20]. Parameter [DZ24, BG21, IWS22, LKAB^{+}22, SCZ24]. PARIS [GLL21]. parity [ZWSL22]. parity-only [ZWSL22]. PARMA [KGPT23]. PARMA-CC [KGPT23]. partial [LMG^{+}21, SBB21]. partial-ISA [SBB21]. Particle [DED^{+}20, ZZZG21, DSZ^{+}21]. Particle-In-Cell [DSZ^{+}21]. partition [DWWX22, DW23, SFT^{+}21]. partitioned [ZLCL24]. partitioner [SKB21]. partitioning [BMK^{+}22, GS20, JSA21, NMPS20]. party [CLT^{+}20, FBL^{+}21, HR23]. passing [DCM^{+}23]. Password [AKS^{+}20]. patch [MM24]. Path [SM22, GMA^{+}22, PAD22, PW21, RBS21, YLDY23, ZLCL24]. paths [Lai21, LZZJ23, LFC^{+}24, MZMM21, CH24]. pattern [GBEFBC24, SALP20]. PBBFMM3D [WCLD21]. PBFT [XBX^{+}22]. PCH [LWW^{+}23]. PCH-based [LWW^{+}23]. PDC [Ada21]. peaks [XWCJ22]. pedagogical [GMS^{+}21]. Peeking [CBO^{+}23]. Penalized [HA21]. Perfect [ZTKL^{+}21, DCM^{+}23]. Performance [BDL22, BPBD23, BV21, CYWL21, CNFMA20, IOG20, LLM^{+}20, TMB^{+}21, TlW23, ZYL^{+}21, dARR21, BKY21, CLW^{+}23b, CCC23, CLZ^{+}22, CRS22, CMR20, DDC^{+}24, DQZZ21, DLL^{+}21, DWR^{+}23, DMM^{+}21, EMCE20, EHH^{+}23, FHN^{+}22, HMS20, KVMR23, LCK23, LJW^{+}22, Mar20, MFAB23, AAAS24, MKP22, MHdC^{+}23, PGB^{+}22, RCX^{+}21, SPP^{+}23, SFML21, SPBR20, SMT22, SBB21, TLC20, XZH^{+}22, XWL^{+}20, YTLF22, ZKL21, ZFL^{+}23, ZdCLT22]. performance-aware [ZFL^{+}23]. performance-driven [CCC23]. performance-related [BKY21]. periodic [YDX^{+}22]. permission [YQZ^{+}20]. permutated [YNI^{+}22]. persistent [OGR^{+}24]. personalized [PW21]. personalized-exchange [PW21]. perspective [DBAC^{+}22]. Petri [MLB21, SLFC22]. PfTouch [TGFPRA20]. phase [CRS22, IMP^{+}23]. phased [MHdC^{+}23]. PHY [CCC23]. physical [NVE^{+}21]. physics [SPBR20]. PIM [ZJW^{+}21]. pipeline [AAAS24, WYH^{+}23, LCK23]. Pipelined [TV22]. placement [FWZ^{+}20, IRLN23, KYZ^{+}20, KHO22, LLR^{+}21, LZL22, LZS^{+}24, MSRB20, MRPH20, SM23, SWM23, SWY^{+}21, VBB22]. planning [GMA^{+}22, LJH^{+}22, NGS21, SM22]. plants [BG21]. plasma [AM22, KNK^{+}23]. plasticity [CGW23]. platform [CSY^{+}24, LZL22]. Platforms [Kur21, CBO^{+}23, CNFMA20, GBC^{+}22, KVMR23, KuR24, NHR22, SFT^{+}21]. plume [AM22]. PMSMC [EMSEMM20]. POCache [ZWSL22]. point [JTV^{+}22, LMG^{+}21, XMJG22]. point-to-all [XMJG22]. point-to-point [XMJG22]. Poisson [AR20, SOS^{+}24]. policies [CTFW22, GSA21, YTLF22]. Policy [OMCW23, HCY^{+}21, WYW^{+}20]. polynomial [Mal22]. pooled [GHKKL23]. pooling [SNT^{+}20]. popularity [CPZ^{+}20]. popularity-aware [CPZ^{+}20]. portability [DDC^{+}24, DMM^{+}21, HMS20]. portable [BDSQO22]. Power [Ato23, CNFMA20, HSHT22, IMP^{+}23, JYH22, PK21b, TAG23, Yaz23a, ZJW^{+}21]. Power-aware [Ato23, IMP^{+}23]. PPMCK [FBL^{+}21]. Practical [ACC^{+}21, PMV20, KSV^{+}20a]. practice [Mar20]. pre [SOL22]. pre-avoidance [SOL22]. precedence [ALP21, Mal21]. precise [Ati20]. precision [GC23, IKK20, NPO^{+}23, WZO^{+}21]. precisions [ATD20]. Preconditioned [TV22]. preconditioners [AR20]. predict [YLDY23]. Predicting [GLL21, JHML21, WZC^{+}20]. Prediction [CYWL21, AKL22, CJZ^{+}22, CHJ^{+}24, HFA20, HLBZ20, KGP^{+}21, QDD^{+}22, TLD^{+}23, WCMS24, WC22, XRBT21, ZDZ^{+}21]. prediction-based [AKL22]. predictions [ACC^{+}23, MAL^{+}23]. predictive [SLSN24, SFML21, ZQL^{+}21]. preemption [LAPB20]. Preemptive [ALP21]. Preface [LBHW20, dARR21]. prefetching [LGZZ23]. Preparing [AKB^{+}20]. prescribed [AMPT23]. presence [JSJC22, KKW23, Mal21, Pou20]. preservation [Rub22]. preserving [BKL^{+}20, BLNP23, FHG^{+}20, FBL^{+}21, GLY^{+}21, GPK21, LWL^{+}22, LWW^{+}23, LQX^{+}20, LLD^{+}23, OK21, UHAH^{+}24, WCG^{+}24]. prevention [HA21, SKS21]. PRI [LWW^{+}23]. prices [MS23]. Pricing [KK21]. Prime [KSV20b]. primitives [ZKL21]. Priority [EMSEMM20, ZYW24]. Priority-based [EMSEMM20]. Privacy [CLW^{+}23b, FHG^{+}20, FBL^{+}21, GLY^{+}21, LQX^{+}20, WCG^{+}24, BKL^{+}20, FLZ^{+}20, GPK21, HNN^{+}20, JJ22, KSV^{+}20a, LWL^{+}22, LWW^{+}23, LLD^{+}23, LBHW20, MRC21, NT20, OK21, RH20, Rub22, UHAH^{+}24, URC20, WMJ^{+}20, WYZ^{+}24]. Privacy-aware [CLW^{+}23b]. privacy-enhanced [HNN^{+}20]. Privacy-preserving [FBL^{+}21, GLY^{+}21, LQX^{+}20, WCG^{+}24, LWL^{+}22, LWW^{+}23, LLD^{+}23, OK21, UHAH^{+}24]. private [URS21, WCT^{+}22]. Privbus [HNN^{+}20]. Proactive [DK24, GA21]. proactively [KSLN24]. Probabilistic [GKP21, PB21]. probe [BRK^{+}21]. problem [AMM21a, CLMH22, GSMÖ23, JSJC22, JH21, QMB21, SPP^{+}23, VBB22, ZTKL^{+}21]. problem-based [QMB21]. problems [FW23, GK21, GKP21]. Process [Jea22, CJZ^{+}22, LJH^{+}22]. processing [ALS23, CSS^{+}23, FQL^{+}23, HCC^{+}20, IKK20, KG20, LoKS24, LYZ^{+}22b, MDS20, MTG^{+}24, VTT^{+}22, WZO^{+}21, WLL^{+}23a, XPW^{+}22, XWJ^{+}23, ZDD^{+}24, ZJW^{+}21]. processor [AMC20, KVMR23, MPR^{+}21, SBSB20, ZYL^{+}21, RSGA20]. processors [BDSQO22, BV21, CRS22, HMCG20, YES22]. procurement [HLL^{+}21]. production [YLDY23]. productive [CMFV^{+}20]. productivity [DMM^{+}21]. Profit [LCH^{+}21, MSRB20, CLMH22]. Profit-aware [MSRB20]. programmers [BKY21]. Programming [CIH^{+}23, AS20, CMFV^{+}20, DDC^{+}24, DMSB20, FFGEL21, JB20, LNW21, MS20, PLBG21, SBL20, SEM20, TAGEL23, XZY^{+}23, YSMB21]. programs [Mal22]. project [YSMB21]. prone [DGFR21]. Proof [Pou20, TZDC21, FF22, NZ23]. proof-labeling [FF22]. Proof-of-Space [TZDC21]. propagation [DZ24, SCZ24]. properties [GXJ^{+}24]. property [RBS21]. protection [JZWX20, KSV^{+}20a, LLGC22, PCC20]. protein [MM24, WC22]. protocol [ABMPL22, AKS^{+}20, CW20, JZWX20, KBS^{+}21, LCZL21, Li23, PJV^{+}22, Rub22, SATJ^{+}20, TB22, XLL^{+}21, ZWS^{+}20]. provenance [CTGJ22]. provision [CCAACS21, QDD^{+}22]. provisioning [JKK^{+}23, LWD^{+}20, MYS^{+}23, PS22, ZZS^{+}21b]. ProvNet [CTGJ22]. proxy [XZH^{+}22]. proxy-oriented [XZH^{+}22]. pruning [DGMS20]. pseudorandom [Stp20]. PTTS [Ato23]. pub [HZL^{+}20]. pub/sub [HZL^{+}20]. Public [CML^{+}24, CW20, GPK21, LQX^{+}20, XZH^{+}22]. publication [LCW^{+}21]. PUF [ABMPL22]. PUF-based [ABMPL22]. PUMIPic [DSZ^{+}21]. punishment [TY23]. purpose [JKM^{+}22, MTG^{+}24, SSD^{+}20]. Python [SZW^{+}22].

Q [WCR^{+}20]. Q-learning [WCR^{+}20]. QoS [CCAACS21, NMPS20, QDD^{+}22, SSG24]. QoS-driven [SSG24]. quadratic [SBL20, SEM20, YNI^{+}22]. quality [CGL^{+}22, GQW^{+}21]. quantised [HLZ22]. quantum [HWR^{+}24, WC22]. quantum-inspired [HWR^{+}24]. queries [RBS21, XMJG22]. query [LQX^{+}20]. querying [BYW^{+}22]. queue [ZYW24]. queuing [RGESG^{+}21, WXZ^{+}23]. QuickDedup [SSG^{+}20]. quicksort [AAZMS20]. Quorums [DO22].

Rabin [GMLW24]. race [TA24, VBB22]. radio [SATJ^{+}20]. RAID2.0 [LLXX24]. Rail [WDL22]. RAM [CLT^{+}20, WCT^{+}22]. random [MAL^{+}23, PD21, ZCY^{+}21]. Randomization [VGMG20, JYH22]. Randomized [BBE^{+}21]. range [BNOS21, LQX^{+}20]. ranite [RBS21]. rank [MAL^{+}23]. RASM [MLR^{+}23]. rates [BRK^{+}21]. RATS [LZY23]. rCUDA [SPBR20]. RD [KST^{+}23]. RD-FCA [KST^{+}23]. RDIC [WHY^{+}21]. RDMA [WLL^{+}23a]. RDMA-based [WLL^{+}23a]. Reachability [Mal22]. reaction [DZ24]. reaction-diffusion [DZ24]. Reactive [KSB^{+}20]. Read [CP24, DGFR21, GHNS22, XDM^{+}22]. Read/write [CP24, DGFR21]. Real [DMKFJ20, BMC^{+}24, CH23, MCD^{+}21, SUD^{+}22, TV22, TNM^{+}22, WHL^{+}23, XZPL24]. Real-time [DMKFJ20, BMC^{+}24, CH23, SUD^{+}22, TNM^{+}22, WHL^{+}23, XZPL24]. real-world [MCD^{+}21]. Realizing [YY22]. reallocation [BCM^{+}21]. Rebeca [MPAS24]. reclamation [MAR21]. recognition [GBEFBC24]. Recommendation [XLL^{+}20, LLD^{+}23, TZC^{+}24, YWF21, YWF23]. recommender [LLD^{+}23]. reconfiguration [DQZZ21, QZW^{+}24]. reconfiguring [DQH^{+}21]. reconstruction [CPZ^{+}20, GHKKL23]. record [NPAM20]. records [LQX^{+}20]. recovery [LLXX24]. Redactable [WDY^{+}24, ZNX^{+}21]. Reducing [BSWO23, ZR22, XWCJ22]. redundancy [XZPL24]. registration [BHB^{+}21, FFGEL21, LMG^{+}21]. regression [MBS^{+}20]. regular [Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, KBW20, ZZ24]. regulatory [LZY23]. Reinforcement [HLX^{+}22, MLR^{+}23, KK22, LLYZ23, SSG24, YHKR20]. reinforcing [SAATK21]. reining [WZC^{+}20]. related [BKY21]. Reliability [GRZT22, LLM24, ZXY21, ZZLM22, CLZ^{+}22, GBC^{+}22, JHZ20, KK22, LSZL20, LJW^{+}22, NPO^{+}23, WHL^{+}23]. Reliable [GKSS24, HZL^{+}20]. remapping [LW20]. remote [SLSN24, WHY^{+}21]. remoting [SM23]. remoting-based [SM23]. renaming [BBE^{+}21]. Rendezvous [PK21a]. renewable [HSHT22, XB20, ZZ22]. Reordering [AA20]. repair [MRPH20]. repair-efficient [MRPH20]. replacement [LZL22, NTT^{+}23, WYW^{+}20]. replica [LSZL20]. replicated [LRV20]. replication [BADP22, SNSK20]. representation [RAJ^{+}23, SBB20]. Reputation [TK23]. request [LLXG21, SWY^{+}21, XWCJ22]. requestor [EMSEMM20]. rerouting [DQH^{+}21, DQZZ21]. resampling [ZZZG21]. RESCAL [BkB^{+}23]. Research [ZZZ^{+}23, SSG21]. reshaping [ALS23]. residual [WYA^{+}21]. resilience [GLL21]. resilient [KST^{+}23]. Resisting [TZDC21]. resize [CDA20]. ResNet [NVE^{+}21]. resolution [WYA^{+}21, WYH^{+}24]. Resource [KK21, MLR^{+}23, dSSE23, CWHC22, CN22, DVS24, HLL^{+}21, JKK^{+}23, LWD^{+}20, MB21, MYS^{+}23, NGS21, PS22, PKSR24, XTGJ21, YY22, ZZS^{+}21b, ZZZ^{+}23]. Resource-Aware [MLR^{+}23]. resources [YTLF22]. responding [Ada21]. response [GSV21, ZR22]. REST [HZY^{+}21]. restricted [TGFPRA20, ZZLM22]. results [CHJ^{+}24]. retrieval [SFT^{+}21]. reusability [LWW^{+}23]. reuse [KSLN24]. review [BJ23, IRLN23, MB21, RCVA22, SSG21, dSSE23]. Reviewer [Ano24-27]. Revisiting [BHP^{+}24, GJL^{+}24]. revocation [GPK21]. Reward [EMCE20, TY23]. RF [ZCY^{+}21]. RF-RISA [ZCY^{+}21]. rings [ADD^{+}20]. RISA [ZCY^{+}21]. risk [BDRJ24, HFA20, MS23]. risk-neutral [MS23]. Rival [HA21]. Rival-Model [HA21]. RL [DWWX22]. RLS [DKS21]. RNA [PB20]. road [GHT^{+}21, IOG20]. Roadside [XLL^{+}21]. robot [LMG^{+}21]. robotic [GMA^{+}22, LJW^{+}22]. robots [BCM23, GMMP24, KMS22]. Robust [AS21, BSY24, HFP^{+}22, SAATK21, WYH^{+}24, ZWSL22]. robustness [DGMS20]. Rotorcraft [MBS^{+}20]. routine [SMT22]. routing [ABB22, CW21a, DPSD21, KBS^{+}21, LLGC22, LLW^{+}20, LZZ^{+}20, MMESG^{+}21, MCC20, MST24, PCC20, PW21, RK24, RGESG^{+}21, SSG24, SWY^{+}21, SATJ^{+}20, YS21, YPD^{+}20]. row [DQH^{+}21, DQZZ21]. RSA [SAY20]. Rule [LD21]. Rule-based [LD21]. running [MYS^{+}23, YLDY23]. runtime [GXJ^{+}24, MPAS24, SMMG22, WLL^{+}23b, YLDY23, ZdCLT22].

safe [DLWF23]. safety [ZDZ^{+}21]. safety-aware [ZDZ^{+}21]. sampling [AMPT23]. Santos [CBO^{+}23]. satellite [LD21]. satisfaction [CLMH22]. save [NMPS20]. saving [MZZW21, MBRR24, XZY^{+}23]. Sboing4Real [TNM^{+}22]. SCAB [VD21]. scalability [LKAB^{+}22, LWW^{+}23, Sok21, dRBB21]. Scalable [BNOS21, BRK^{+}21, JJ22, OGR^{+}24, RdCR^{+}24, SMMG20, SBSB20, XDM^{+}22, CGW23, DWWX22, FPdLS^{+}21, KBS^{+}21, NCR23, NMA^{+}24, RBS21, RLW^{+}24, SÖAOA20, WCTW22, YNI^{+}22, ZZS^{+}21a]. scalar [ZD22]. scale [CDTS24, GKSS24, KGP^{+}21, MAL^{+}23, PK21a, SBL20, STG^{+}20, WYA^{+}21, YZM23]. scales [LXC^{+}22]. scaling [BSWO23, DK24, GA21, MTR22, ZGTM24]. scan [PAD22]. scenario [WCMS24]. scenarios [FQL^{+}23, SZQ^{+}23]. Schedulability [LAPB20]. schedule [Mal21]. scheduled [LAPB20]. Scheduler [EMSEMM20]. schedulers [ZdCLT22]. schedules [BCEH23]. Scheduling [Che23, KRK20, Li24, TZZ^{+}20a, ALP21, BADP22, BS24, CW21a, GSA21, GSV21, HWM^{+}23, HSHT22, HKTG20, HWR^{+}24, IMP^{+}23, JH21, KYGG20, KD22, KL22, KuR24, KPS^{+}22, LWHF22, LCC20, LYC23, NHR22, OS20, PC21, PKSR24, SNSK20, TZZ^{+}20b, TLD^{+}23, TDCM21, WWH^{+}21, WXAL22, WYZ^{+}24, WHL^{+}23, XZPL24, ZYW24, ZQL^{+}21, ZD22, ZR22]. Scheme [KRK20, AB22a, DQH^{+}21, DQZZ21, FLZ^{+}20, FQL^{+}23, GKB^{+}20, KSV^{+}20a, LSWY20, LLGC22, LYZ^{+}22b, MBB22, MBN22, TY23, TDCM21, WHY^{+}21, WCG^{+}24, XTGJ21, XZH^{+}22, XWJ^{+}23, Yaz23a, YQZ^{+}20, ZNX^{+}21]. schemes [FF22, JH21, LZZJ23, RGESG^{+}21]. science [BMC^{+}24, LCW^{+}21]. scientific [ACC^{+}21, GM21, JZS^{+}20]. SCIPIS [OGR^{+}24]. screening [VFB^{+}24]. scrubbing [JHZ20]. SD [AJH^{+}20]. SD-WAN [AJH^{+}20]. SDN [Ala24, HZY^{+}21, OMCW23, SSG24, CH24]. SDN-enabled [HZY^{+}21]. SDWAN [AJH^{+}20]. SEAPP [HZY^{+}21]. Search [AGMG23, KL22, OGR^{+}24, SPP^{+}23, SW22, YQZ^{+}20, DPSD21]. searches [Gow21]. SECDA [HGC^{+}23]. SECDA-TFLite [HGC^{+}23]. section [CZR^{+}24, OZ22]. sector [Rub22]. Secure [JSJC22, NVE^{+}21, VD21, AKY20, AMP20, AKS^{+}20, CGC21, DWW^{+}21, FLZ^{+}20, HZY^{+}21, Jar20, KKG^{+}23, LLGC22, Li23, MBB22, NYZ^{+}20, OK21, PCC20, SMHK21, WLZ20, XBX^{+}22, YPD^{+}20, YQZ^{+}20]. Secured [BL23, KTM^{+}21]. Securing [Ala24, LLF^{+}20, LSWY20]. Security [LBHW20, WWH^{+}21, ABMPL22, AZF^{+}24, FTA^{+}22, GLF20, HFP^{+}22, HCY^{+}21, JJ22, KTM^{+}21, LCH^{+}21, LLW^{+}20, RH20, SALP20, ZLD22]. Security-aware [WWH^{+}21, LCH^{+}21]. SED [PYX^{+}22]. seed [RSSP23]. seer [ESA24]. Selection [VJR20, BGDT22, BGDMF^{+}24, BkB^{+}23, CLW^{+}23a, DEFQO21, NRGL22, SEM20]. Selective [EL20]. Self [HA21, SDLM20, UHAH^{+}24, BT20, BDRJ24]. Self-Organizing [HA21]. Self-sovereign [UHAH^{+}24]. Self-stabilizing [SDLM20, BT20]. self-supervised [BDRJ24]. Semantic [SBBP20, ACCN20]. semantic-based [ACCN20]. Semantic-centric [SBBP20]. semiconductor [SKK21]. sensed [AKB^{+}20]. sensibility [CL22a]. Sensitive [LAPB20, FF22]. sensitivity [RPM24]. Sensor [KRK20, AMM21a, Amm21b, GHD20, PG20, SM22, SAATK21, SATJ^{+}20, WMJ^{+}20, ZYW24]. sensor-cloud [WMJ^{+}20]. sensors [MBS^{+}20]. Separating [ACHP22]. separation [KLL^{+}21]. September [Ano20-28, Ano21y, Ano22-32, Ano23-36, Ano24-28]. sequence [KEK^{+}20]. sequences [LLXG21]. Sequential [CTA20, GSMÖ23, TA24]. series [GBC^{+}22, QGP24]. series-parallel [GBC^{+}22]. serious [MMM21]. server [CRS22, LLR^{+}21, XZY^{+}23]. server-class [CRS22]. Serverless [SSA24a, DAG24, ESA24, FPGLSA24, NRdA^{+}20, SSA24b]. servers [Li24]. Service [BPT^{+}22, Kur21, MLR^{+}23, HNN^{+}20, HCY^{+}21, HLX^{+}22, IRA20, KYGG20, LZGL22, PS22, TZC^{+}24, WCR^{+}20, XWCJ22, YLL21, ZR22]. service-based [PS22]. services [ASHO20, AAD^{+}20, HLK^{+}22, LZD21, NGS21, ZGTM24]. session [Zha23]. set [FLFZTS20]. SG [XBX^{+}22]. SG-PBFT [XBX^{+}22]. SGD [BSY24, CLZ20, YDX^{+}22]. Shapley [PK21b]. sharding [TTD24]. sharding-based [TTD24]. shared [BBE^{+}21, DDC^{+}24, GPK21, JSA21, LZS^{+}24, MYS^{+}23, QM21, TZDC21]. shared-memory [DDC^{+}24, JSA21]. sharing [BHP^{+}24, CTGJ22, FLZ^{+}20, SMHK21, SC23, TY23]. shelf [BCEH23]. Short [DLL^{+}21, MD20, BNOS21]. Short- [DLL^{+}21]. short-range [BNOS21]. Shortest [PW21]. Shortest-path [PW21]. Shubik [PK21b]. shuffle [RWF^{+}21]. shuffling [ESA24]. siamese [JJJ21]. sidechain [LWL^{+}22]. sidechain-based [LWL^{+}22]. sided [Ato23]. sight [JTV^{+}22]. Sigmoid [PSBB21]. signature [KSV^{+}20a, MLTT20]. silent [GMMP24]. silhouette [ŠK23]. SIMD [AS20]. SimGQ [XMJG22]. similar [GBEFBC24]. similarity [SWF^{+}22, TZC^{+}24, WWL^{+}21]. simple [AMPT23]. simplex [MN24]. Simplifying [LCW^{+}21, RH20]. simulated [VBB22]. Simulating [CGW23]. Simulation [CL22a, AJH^{+}20, CTKdS21, DF22, FSL^{+}21, KD22, PK21a, TLC20]. Simulation-based [CL22a]. simulations [AM22, CMR20, DWR^{+}23, KNK^{+}23]. simultaneous [LMG^{+}21]. Simultaneously [XMJG22]. single [MLTT20, MKP22]. situ [DWR^{+}23]. sizes [SPP^{+}23]. Sketch [DGA^{+}24]. Sketch-fusion [DGA^{+}24]. skills [YSMB21]. skinny [RCX^{+}21]. Skip [HNKÖ21, THPM22]. SLA [FWZ^{+}20, TDCM21]. slack [CRS22]. sLASs [VLCM^{+}20]. slicing [Amm21b, SPK^{+}23]. slim [BCM23]. slowdown [WZC^{+}20]. small [ATD20, ZTKL^{+}21]. Smart [ARL20, CL22b, YLL21, Ala24, AKB^{+}20, CJZ^{+}22, CGC21, DPEL24, GF20, GLY^{+}21, JJJ21, KPS^{+}22, KSV^{+}20a, LZZ^{+}20, MLTT20, OK21, Pou20, WLZ20, XTGJ21]. SMDP [NHR22]. SMDP-based [NHR22]. smell [CH24]. SMS [WYH^{+}21]. SMS-based [WYH^{+}21]. SMT [FRAK23]. SNARK [NZ23]. Sobel [CLLM23]. Social [RGB20, KYZ^{+}20, LBHW20]. SocialBlock [GF20]. soft [RPM24, SUD^{+}22, YSMB21]. soft-error [RPM24]. software [AB22a, HA21, LCC20, NCR23, SSD^{+}20, YES22, YPD^{+}20, AJH^{+}20]. software-defined [AB22a, AJH^{+}20]. solar [HWM^{+}23]. solution [LWL^{+}22, LLC20]. Solutions [FTA^{+}22, KGTK20, SZQ^{+}23]. solvable [ACR23]. solve [DGFR21]. solver [AR20, DDC^{+}24, HMS20, KN24, NE23, YNI^{+}22]. solvers [DEFQO21, KNK^{+}23, PG20]. solving [AMM21a, GK21, GKP21, SOS^{+}24]. sort [KKH^{+}23, TYOC24]. sort-based [TYOC24]. sorting [SBSB20]. Soundness [BLNP23]. Soundness-preserving [BLNP23]. source [LQX^{+}20, SW22, Zha23]. sovereign [UHAH^{+}24]. Space [TZDC21, VGMG20, KL22, MAL^{+}23, OS20, SPP^{+}23, SDLM20, ZJW^{+}21]. space-time [MAL^{+}23]. spanning [KKH^{+}23]. spare [ZJW^{+}21]. Spark [YWF23, CYWL21, GBEFBC24, LZL22, LZZ^{+}22, TZZ^{+}20b, YWF21]. SparkDQ [GQW^{+}21]. Sparse [AS21, AA20, DEFQO21, GXYH21, KNK^{+}23, SZW^{+}22, TLC20]. sparsity [Ato23]. Spartan [AS21]. Spatial [AHAB23, WCMS24, WZC^{+}20]. Spatiotemporal [SCZ24]. Special [Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t, BDL22, GDF^{+}23, MMM22, dARR21, MS20]. specific [CSY^{+}24]. specification [SLFC22]. spectral [HMCG20]. spectrometry [SHI22]. spectrum [SATJ^{+}20]. speculation [LRV20]. Speculative [FRAK23]. Speed [PD21, FHN^{+}22, GVI24, PAD22]. Speed-area [PD21, PAD22]. speeds [BCEH23]. Speedup [Sch24]. SPH [MBN22]. spherical [Li22a]. Spiking [JHML21, WZO^{+}21]. spine [CW21a]. SpMV [GJL^{+}24, ZYL^{+}21]. spread [BMC^{+}24]. SpTV [WYH^{+}23]. square [NA23]. squares [NA23, PG20]. SSD [CLZ^{+}22]. SSDs [SLZP24]. SSSP [CHJ^{+}24]. Stab [SADM24]. Stab-FD [SADM24]. stabbing [RSGA20]. stabilization [HNKÖ21]. stabilizing [BT20, SDLM20]. stack [SWF^{+}22]. stage [HFA20, KGPT21]. staging [JZS^{+}20]. staging-based [JZS^{+}20]. Starlight [ZDD^{+}24]. state [BADP22, OS20]. state-space [OS20]. static [KL22]. Stealing [GKTW21, CP24, FPdLS^{+}21]. Steiner [RSSP23, SK21, WHC21]. steroids [BADP22]. Stochastic [EMCE20, DBAC^{+}22, EBV22, PXY^{+}20, YDX^{+}22, TSTY22]. storage [BW22, CPZ^{+}20, CTFW22, CML^{+}24, ESA24, GHNS22, GPK21, JJ22, KJA^{+}22, MRPH20, SÖAOA20, WCT^{+}22, YLZ^{+}20, ZYL^{+}21, ZCD^{+}21]. store [FRAK23]. store-to-load [FRAK23]. stores [LRV20]. Storing [BYW^{+}22]. straggler [ZWSL22]. Strategic [DBAC^{+}22]. strategies [BHP^{+}24, ODXX21, WHL^{+}23]. strategy [GVC^{+}22, LLM^{+}20, LMG^{+}21, LZL22, MZZW21, MM24, SUD^{+}22, TZZ^{+}20a, ZZS^{+}21b, ZYW24]. stream [AMC20, CKS22, KG20, MTG^{+}24, VTT^{+}22, XPW^{+}22]. streaming [GBC^{+}22, JSA21]. strength [AT24]. stretching [Amm21b]. stripe [BW22]. stripe-based [BW22]. structural [CGW23, WWL^{+}21]. structure [ASHO20, SSH23, WC22]. Structured [GS20, XJR21]. structures [DHF23, JKM^{+}22]. studies [SZQ^{+}23]. Study [KGP^{+}21, Ada21, KLL^{+}21, SMMG22, TLC20]. SUARA [NMA^{+}24]. sub [HZL^{+}20]. subarrays [DQH^{+}21]. subcubes [WFS^{+}22]. Subgraph [GQX20, FDT^{+}24, XWJ^{+}23]. subgraphs [KBW20]. substrings [CTA20]. Subversion [GPC23]. succinct [XDM^{+}22]. suite [MKP22]. SunWay [XDM^{+}22, LFJ^{+}20]. super [WYA^{+}21]. super-resolution [WYA^{+}21]. supercomputer [CDTS24, LFJ^{+}20, RMJM21]. supercomputers [YLDY23]. supervised [BDRJ24]. support [BMC^{+}24, LAPB20, NPO^{+}23, PD21, Yaz23b]. Supporting [TAGEL23, DHF23, KK21, TY23]. surplus [YTLF22]. survey [DM20, LCC20, ODXX21, SZQ^{+}23, ZDL^{+}24]. sustainability [BJ23]. Swap [KW20]. Swarm [DED^{+}20]. switch [AB22a, CCAACS21]. switching [AMPT23]. switchless [NPY^{+}23]. synchronization [OZ22, PPN^{+}20, PMV20, ZJW22]. synchronizing [KEK^{+}20]. synchronous [DPEL24]. synchronously [BLNP23]. synthesis [JJJ21]. System [EMSEMM20, AGMG23, BCM^{+}21, BYW^{+}22, BMC^{+}24, CWHC22, CLZ20, CDA20, DAG24, DPEL24, DFL^{+}23, DZ24, DEFQO21, FDT^{+}24, FFS^{+}22, GMS^{+}21, HSX^{+}21, KTM^{+}21, KGP^{+}21, KKT^{+}22, KKG^{+}23, LMG^{+}21, LZY23, MD20, MPAS24, PSU^{+}21, PK21b, RLW^{+}24, RKAA20, SMMG22, TNM^{+}22, WLL^{+}23a, XRBT21, ZCD^{+}21]. Systematic [KL22, BJ23, IRLN23, LFJ^{+}20, RCVA22, SSG21, dSSE23]. systems [BW22, BMK^{+}22, BBE^{+}21, BDFG21, BSWO23, CZR^{+}24, CPZ^{+}20, DGFR21, DGMS20, DHF23, GM21, HQL^{+}22, KJA^{+}22, KKW23, KLL^{+}21, KSV20b, MCD^{+}21, MDS20, AAAS24, MLB21, MAL^{+}23, MBRR24, NMPS20, NVE^{+}21, OGR^{+}24, OZ22, PJV^{+}22, PSBB21, Sok21, SMT22, SLFC22, SC23, TZZ^{+}20a, TDL^{+}22, TV22, VGMG20, WMJ^{+}20, XWL^{+}20, XZPL24, Yaz23a, YY22].

TaihuLight [LFJ^{+}20, XDM^{+}22]. Taking [MGW24]. tall [RCX^{+}21]. tall-and-skinny [RCX^{+}21]. Targeting [VTT^{+}22]. Task [BS24, KSB^{+}20, YTLF22, BCM^{+}21, DGMS20, FQL^{+}23, HWM^{+}23, KD22, LCH^{+}21, NHR22, OS20, PLBG21, TYM^{+}22, WWH^{+}21]. task-based [PLBG21]. Task-Parallel [KSB^{+}20]. tasks [ACR23, BKY21, BCEH23, HKTG20, KL22, Li24, Mal21, SC23]. taxi [XRBT21]. taxonomy [SSG21]. Teaching [CTKdS21, GHT^{+}21, MMM21, PGB^{+}22, CCSI21]. technique [CPZ^{+}20, DK24]. Techniques [DGWD21, ACC^{+}23, LYC23]. technological [Ada21]. Technologies [BGA^{+}21, MMM22]. technology [PGB^{+}22, URC20, ZZZ^{+}23]. temporal [GXJ^{+}24, GA21, GKP21, RBS21].

Tenant [VGMG20, HZL^{+}20]. Tensor [SPK^{+}23, Ato23, SZW^{+}22, ZKL21]. term [DLL^{+}21, MD20]. terminal [XZH^{+}22]. terms [LZWZ22, ZXY21, ZZLM22, YTLF22]. ternary [MZMM21]. tessellation [NA23]. testable [PD21]. TFLite [HGC^{+}23]. Thank [Ano24-27]. theft [JJJ21]. their [JKM^{+}22]. theoretic [GHKKL23]. theory [DZ24, Mar20, WXZ^{+}23, XWCJ22]. Thermal [MCC20, NHR22]. Thermal-aware [MCC20, NHR22]. Things [Alm20, AGMG23, CJZ^{+}22, FTA^{+}22, FFS^{+}22, GLF20, GDF^{+}23, YS21]. third [HR23]. thread [DGWD21, FRAK23, GJL^{+}24, LW20]. thread-data [LW20]. threaded [WLL^{+}23b]. threading [KD22]. threads [Mal22]. three [CC23, GHNS22, Li23, MBB22]. three-factor [Li23, MBB22]. three-way [CC23]. thresholding [DT21]. throttle [KGP^{+}21]. throttling [MPR^{+}21]. throughput [SMMG22]. THz [SMMG20]. THz-band [SMMG20]. TIGRE [BLB^{+}20]. Tile [MAL^{+}23]. tiled [PB20]. Time [LAPB20, MDS20, MMA22, ACC^{+}23, BMC^{+}24, Che23, CHJ^{+}24, CH23, DMKFJ20, GSV21, LYZ^{+}22a, Mal21, MAL^{+}23, MPAS24, QGP24, RKAA20, SUD^{+}22, SC23, TZC^{+}24, TNM^{+}22, WHL^{+}23, XZPL24, ZSL^{+}23]. time-aware [TZC^{+}24]. time-delay [ZSL^{+}23]. time-dependent [LYZ^{+}22a]. Time-energy [MDS20]. time-evolving [CHJ^{+}24]. Time-Sensitive [LAPB20]. time-sharing [SC23]. Time-varying [MMA22]. Tiny [MPAS24]. Title [Ano20x, Ano20w, Ano24m, Ano24u, Ano24n, Ano24o, Ano24p, Ano24q, Ano24r, Ano24s, Ano24t]. token [SDLM20, URM23]. tolerability [WFS^{+}22]. tolerance [GQX20, YZM23, ZWSL22]. tolerant [AM20, DLWF23, GSV21, LFC^{+}24, MBRR24, PG20, SNSK20]. tomography [BLB^{+}20]. Tool [CFRGPM22]. toolbox [BLB^{+}20]. toolkit [HGC^{+}23]. tools [Ada21, DGWD21]. top [RMJM21]. top-level [RMJM21]. topics [DF21, GMS^{+}21]. topologies [RGESG^{+}21]. topology [Ati20, Jea22, KD21, MST24, QZW^{+}24]. TopoMatch [Jea22]. torus [MST24]. total [Che23]. TPU [HLS^{+}23]. trace [SWF^{+}22]. tracing [BPBD23]. trade [MDS20]. trade-offs [MDS20]. tradeoffs [LZGL22]. Trading [ZJW^{+}21, GLY^{+}21]. traditional [OMCW23]. traffic [EBV22, GVI24, HLL^{+}22, JHML21, LAPB20, TNM^{+}22, TLC20, URM23]. Training [CGDS20, SSA24a, BDSQO22, CCC23, CLZ20, DGA^{+}24, HLZ22, LKAB^{+}22, AAAS24, ODXX21, SSA24b, SUD^{+}22, ZZS^{+}21a, ZJW22]. trajectory [CZR^{+}24]. transaction [LZY23]. transactional [LRV20, LLG^{+}24, MHdC^{+}23, NCR23, TGFPRA20]. Transactions [AMC20, KJA^{+}22, SSD^{+}20]. transceiver [Yaz23a]. transfer [RAJ^{+}23]. transfers [AA21]. transient [AGC^{+}21]. transition [MPAS24, MHdC^{+}23]. transmission [KKG^{+}23, LLGC22, PJV^{+}22, QDD^{+}22]. transmissions [LSWY20]. Transparent [LRV20]. transport [SKK21]. transportation [MN24]. transpose [AR20]. transpose-free [AR20]. traversal [MLGC^{+}21]. Tree [DPSD21, RGESG^{+}21, Ati20, LD21, LSC22, MLGC^{+}21, SK21, ŠK23, WDL22]. trees [GS20, KKH^{+}23, RSSP23, SDLM20, WHC21]. Trends [MMM22, SSG21]. triangular [DEFQO21]. trie [AR21, MAR21]. triple [XZPL24]. Trust [AkBA^{+}20, KBS^{+}21, LHL21, PJV^{+}22, TK23, TZC^{+}24, TTD24, TYM^{+}22]. trusted [HCY^{+}21, ZWS^{+}20, ZZZ^{+}23]. trustworthy [ZNX^{+}21]. Trustzone [DWW^{+}21]. Trustzone-based [DWW^{+}21]. TSM2X [RCX^{+}21]. TSpoon [AMC20]. TU [SAY20]. Tukey [MAL^{+}23]. tuned [ESA24, PSBB21, VLCM^{+}20]. tuning [BG21, DZZ^{+}23]. TurBO [DZZ^{+}23]. turbulent [NE23]. Twins [MPAS24]. Two [KW20, Amm21b, Ato23, IMP^{+}23, TZC^{+}24, XTGJ21, Yaz23b]. two-dimensional [Amm21b, TZC^{+}24]. two-layer [XTGJ21]. two-level [Yaz23b]. two-phase [IMP^{+}23]. two-sided [Ato23]. type [AR20, MZMM21, ZGTM24].

ULSED [PYX^{+}22]. ultra [PYX^{+}22]. ultra-lightweight [PYX^{+}22]. unbalanced [CLZ20, FPGLSA24]. unbalanced-workload [CLZ20]. uncertainty [CNFMA20]. unconstrained [YNI^{+}22]. Uncovering [CBO^{+}23]. undergrads [CCSI21]. underlying [LCK23]. Understanding [LXC^{+}22, RMJM21, WDL22, GSA21, VGTSG^{+}21]. unicast [DLWF23]. unidirectional [ADD^{+}20]. unified [LGZZ23, NTT^{+}23, RLW^{+}24, GM21]. uniform [AMPT23, FSL^{+}21]. unifying [Sch24]. Unit [XLL^{+}21, Mal21]. Unit-assisted [XLL^{+}21]. units [IKK20, LoKS24, WZO^{+}21]. universal [NMA^{+}24]. unlabeled [ZDZ^{+}21]. unmanned [ZSL^{+}23]. unrelated [ALP21]. unreliable [SWY^{+}21]. unrolled [PMV20]. unstructured [DDC^{+}24, DSZ^{+}21, NE23]. untrusted [YPD^{+}20]. update [BW22, LKAB^{+}22, YQZ^{+}20]. updates [LSC22]. urban [EBV22]. usage [SC23, ZZ22]. use [LCW^{+}21, dSSE23]. user [CN22, DLL^{+}21, GF20, GPK21, MRC21, TY23]. user-centric [GF20]. user-collaborative [CN22]. users [JZWX20, LWD^{+}20]. Using [FHN^{+}22, THPM22, ACC^{+}23, AS20, Amm21b, AGMG23, Ato23, BYW^{+}22, BLB^{+}20, BG21, BSWO23, CZR^{+}24, CJZ^{+}22, DMKFJ20, DDC^{+}24, DT21, DK24, DLWF23, DPSD21, EMCE20, FSL^{+}21, GBEFBC24, GS20, GA21, GLL21, HSHT22, HR23, HKTG20, IRA20, Jar20, JJ22, JTV^{+}22, JP22, KMS22, KK21, LoKS24, LMG^{+}21, LLXG21, LD21, MBM^{+}20, MM24, MKP22, MGE20, MPAS24, NA23, NGS21, NVE^{+}21, PMV20, RAJ^{+}23, SSG24, SMMG20, SNSK20, STG^{+}20, ŠK23, SMHK21, SZW^{+}22, SUD^{+}22, SBB20, Stp20, VD21, WZO^{+}21, XWL^{+}20, XJR21, YTLF22, ZAB20, ZD22, ZdCLT22]. utility [KJA^{+}22]. Utilization [AKL22, HLBZ20, SFML21, WCMS24]. Utilization-prediction-aware [HLBZ20]. UTXO [CW20].

vaccine [BMC^{+}24]. valency [ACR23]. Vampire [DPEL24]. VANET [KBS^{+}21]. VANETs [ABB22, LHL21]. Variability [CL22a, XWL^{+}20]. variable [WZO^{+}21]. variables [CaTZ^{+}24, SFML21]. variants [WFS^{+}22]. Variational [NT20, XXZ^{+}24]. variegated [PAD22]. varying [MMA22]. vector [AS20, GXYH21, GRZT22, HLL^{+}22, KSV20b, TLC20, WCLD21]. vectorization [SALP20]. vehicle [GVI24, QDD^{+}22, ZSL^{+}23]. vehicle-based [GVI24]. Vehicles [MMM22, XLL^{+}21, XBX^{+}22]. vehicular [AB22a, LYZ^{+}22b, ABB22]. verifiable [CTGJ22, WDY^{+}24]. Verification [AA21, GXJ^{+}24, LZZ^{+}20, NYZ^{+}20, YLZ^{+}20]. verified [LLW^{+}20]. versioned [NCR23]. vertex [WWL^{+}21]. vertices [RSSP23]. vGPU [SM23]. via [HNN^{+}20, JHZ20, LLG^{+}24, LTBY20, MFAB23, MBS^{+}20, OZ22, SC23, TZDC21, WYA^{+}21, WYH^{+}23]. Video [LLC20, KTM^{+}21, WYW^{+}20]. videos [GSA21]. Virtual [GM21, XLL^{+}20, HLBZ20, LGZZ23, MBS^{+}20, SBBP20, VFB^{+}24, WYH^{+}24, ZFL^{+}23]. virtualization [MB21, NRdA^{+}20, SM23]. virtualized [PC21]. visibility [BCM23]. Visual [GSA21]. visually [DGWD21]. VLSI [DQH^{+}21, DQZZ21, PD21]. VM [AKL22, PPN^{+}20, SSG^{+}20, SM23, ZFL^{+}23]. voltage [BSWO23, MTR22]. voltage-frequency [MTR22]. volumes [SKB21]. VSIM [WWL^{+}21]. Vulnerability [AM20].

wait [ACHP22]. wait-freedom [ACHP22]. wallet [DWW^{+}21]. WAN [AJH^{+}20]. Warehouses [AHAB23]. warp [FDT^{+}24]. warp-centric [FDT^{+}24]. water [SFZ23]. WAVEWATCH [Li22a]. Waving [YM21]. way [CC23]. weakest [DGFR21]. web [DK24]. weight [GRZT22, VTT^{+}22]. Weighted [CLZ20, Che23]. weights [WZO^{+}21]. whitepapers [LTBY20]. Wide [AJH^{+}20, SADM24]. WiFi [JTV^{+}22]. wind [MBM^{+}20]. WindFlow [MTG^{+}24]. windows [QM21]. Wireless [KRK20, Amm21b, CL22b, GHD20, LSWY20, LCZL21, PG20, SM22, SAATK21, SMMG20, SOL22, ZYW24]. wise [CaTZ^{+}24]. within [PB21]. without [Ati20, BL23]. witness [Pou20]. WNoC [Yaz23a]. wolf [KuR24]. Work [GKTW21, CP24, FPdLS^{+}21, SFT^{+}21]. work-stealing [CP24]. Workflow [KD22, BLNP23, BMC^{+}24, HWR^{+}24, LWHF22, SNSK20, WYZ^{+}24]. workflows [JZS^{+}20, PKSR24, WHL^{+}23]. Workload [CTFW22, CWHC22, CLZ20, FPdLS^{+}21, GBEFBC24, HSHT22, LYZ^{+}22a, TLD^{+}23, WLL^{+}23a]. Workload-aware [CTFW22]. workloads [FPGLSA24, YTLF22, ZPN^{+}21, ZBF^{+}24]. works [HDJ21]. world [MCD^{+}21]. WP [CLZ20]. WP-SGD [CLZ20]. WRENCH [CTKdS21]. write [CP24, DGFR21].

X [HCC^{+}20]. Xilinx [MKP22]. XML [BYW^{+}22]. XML2HBase [BYW^{+}22].

yzantine [CGG^{+}23].

Zernike [JL23]. Zero [HCC^{+}20, NZ23]. Zero-copy [HCC^{+}20]. ZeroCross [LWL^{+}22]. zk [NZ23]. zk-SNARK [NZ23]. zone [GM21]. Zoro [LJW^{+}22].

References

Acer:2020:RSM

[AA20] Seher Acer and Cevdet Aykanat. Reordering sparse matrices into block-diagonal column-overlapped form. Journal of Parallel and Distributed Computing, 140(??):99–109, June 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303885.

Alhussen:2021:ADL

[AA21] Ahmed Alhussen and Engin Arslan. Avoiding data loss and corruption for file transfers with fast integrity verification. Journal of Parallel and Distributed Computing, 152(??):33–44, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100023X.

Maruf:2024:ODT

[AAAS24] Md Al Maruf, Akramul Azim, Nitin Auluck, and Mansi Sahi. Optimizing DNN training with pipeline model parallelism for enhanced performance in embedded systems. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000546.

Anisetti:2020:CED

[AAD^{+}20] Marco Anisetti, Claudio A. Ardagna, Ernesto Damiani, Filippo Gaudenzi, and Gwanggil Jeon. Cost-effective deployment of certified cloud composite services. Journal of Parallel and Distributed Computing, 135(??):203–218, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305756.

Al-Adwan:2020:PQA

[AAZMS20] Aryaf Al-Adwan, Rawan Zaghloul, Basel A. Mahafzah, and Ahmad Sharieh. Parallel quicksort algorithm on OTIS hyper hexa-cell optoelectronic architecture. Journal of Parallel and Distributed Computing, 141(??):61–73, July 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307871.

Aljeri:2022:EHS

[AB22a] Noura Aljeri and Azzedine Boukerche. An efficient heuristic switch migration scheme for software-defined vehicular networks. Journal of Parallel and Distributed Computing, 164(??): 96–105, June 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000181.

Azadani:2022:DDI

[AB22b] Mozhgan Nasr Azadani and Azzedine Boukerche. DriverRep: Driver identification through driving behavior embeddings. Journal of Parallel and Distributed Computing, 162(??):105– 117, April 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000168.

Azhdari:2022:CBR

[ABB22] Mohammad Sadegh Azhdari, Ali Barati, and Hamid Barati. A cluster-based routing method with authentication capability in Vehicular Ad hoc Networks (VANETs). Journal of Parallel and Distributed Computing, 169(??):1–23, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001447.

Antoniadis:2023:LC

[ABD^{+}23] Karolos Antoniadis, Julien Benhaim, Antoine Desjardins, Elias Poroma, Vincent Gramoli, Rachid Guerraoui, Gauthier Voron, and Igor Zablotchi. Leaderless consensus. Journal of Parallel and Distributed Computing, 176(??):95–113, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000151.

Adeli:2022:CSP

[ABMPL22] Morteza Adeli, Nasour Bagheri, Honorio Martín, and Pedro Peris-Lopez. Challenging the security of “A PUF-based hardware mutual authentication protocol”. Journal of Parallel and Distributed Computing, 169(??):199–210, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001538.

Aldinucci:2021:PPS

[ACC^{+}21] Marco Aldinucci, Valentina Cesare, Iacopo Colonnelli, Alberto Riccardo Martinelli, Gianluca Mittone, Barbara Cantalupo, Carlo Cavazzoni, and Maurizio Drocco. Practical parallelization of scientific applications with OpenMP, OpenACC and MPI. Journal of Parallel and Distributed Computing, 157(??):13–29, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521001295.

Amaris:2023:EET

[ACC^{+}23] Marcos Amaris, Raphael Camargo, Daniel Cordeiro, Alfredo Goldman, and Denis Trystram. Evaluating execution time predictions on GPU kernels using an analytical model and machine learning techniques. Journal of Parallel and Distributed Computing, 171(??):66–78, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001903.

Amato:2020:SBM

[ACCN20] Flora Amato, Aniello Castiglione, Giovanni Cozzolino, and Fabio Narducci. A semantic-based methodology for digital forensics analysis. Journal of Parallel and Distributed Computing, 138(??):172–177, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519300644.

Attiya:2022:SLF

[ACHP22] Hagit Attiya, Armando Castañeda, Danny Hendler, and Matthieu Perrin. Separating lock-freedom from wait-freedom at every level of the consensus hierarchy. Journal of Parallel and Distributed Computing, 163(??):181–197, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000302.

Attiya:2023:LST

[ACR23] Hagit Attiya, Armando Castañeda, and Sergio Rajsbaum. Locally solvable tasks and the limitations of valency arguments.

Journal of Parallel and Distributed Computing, 176(??):28– 40, June 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000175.

Adams:2021:EPC

[Ada21] Joel C. Adams. Evolving PDC curriculum and tools: a study in responding to technological change. Journal of Parallel and Distributed Computing, 157(??):201–219, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001490.

Altisen:2020:EUR

[ADD^{+}20] Karine Altisen, Ajoy K. Datta, Stéphane Devismes, Anaïs Durand, and Lawrence L. Larmore. Election in unidirectional rings with homonyms. Journal of Parallel and Distributed Computing, 146(??):79–95, December 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303440.

Ariza:2021:IAA

[AGC^{+}21] Jairo Ariza, Kelly Garcés, Nicolás Cardozo, Juan Pablo Rodríguez Sánchez, and Fernando Jiménez Vargas. IoT architecture for adaptation to transient devices. Journal of Parallel and Distributed Computing, 148(??):14–30, February 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303737.

Asgharzadeh:2023:ABI

[AGMG23] Hossein Asgharzadeh, Ali Ghaffari, Mohammad Masdari, and Farhad Soleimanian Gharehchopogh. Anomaly-based intrusion detection system in the Internet of Things using a convolutional neural network and multi-objective enhanced Capuchin Search Algorithm. Journal of Parallel and Distributed Computing, 175(??):1–21, May 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002611.

AitErrami:2023:SBD

[AHAB23] Soukaina Ait Errami, Hicham Hajji, Kenza Ait El Kadi, and Hassan Badir. Spatial big data architecture: From data warehouses and data lakes to the LakeHouse. Journal of Parallel and Distributed Computing, 176(??):70–79, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000229.

Alwasel:2020:ISS

[AJH^{+}20] Khaled Alwasel, Devki Nandan Jha, Eduardo Hernandez, Deepak Puthal, Mutaz Barika, Blesson Varghese, Saurabh Kumar Garg, Philip James, Albert Zomaya, Graham Morgan, and Rajiv Ranjan. IoTSim-SDWAN: a simulation framework for interconnecting distributed datacenters over Software-Defined Wide Area Network (SD-WAN). Journal of Parallel and Distributed Computing, 143(??):17–35, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151930886X.

Amah:2020:PON

[AKB^{+}20] Tekenate E. Amah, Maznah Kamat, Kamalrulnizam Abu Bakar, Waldir Moreira, Antonio Oliveira, and Marcos A. Batista. Preparing opportunistic networks for smart cities: Collecting sensed data with minimal knowledge. Journal of Parallel and Distributed Computing, 135(??):21–55, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519306240.

Al-khafajiy:2020:CFC

[AkBA^{+}20] Mohammed Al-khafajiy, Thar Baker, Muhammad Asim, Zehua Guo, Rajiv Ranjan, Antonella Longo, Deepak Puthal, and Mark Taylor. COMITMENT: a fog computing trust management approach. Journal of Parallel and Distributed Computing, 137(??): 1–16, March 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519302965.

Awad:2022:UPB

[AKL22] Mirna Awad, Nadjia Kara, and Aris Leivadeas. Utilization prediction-based VM consolidation approach. Journal of Parallel and Distributed Computing, 170(??):24–38, December 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001782.

Amin:2020:CPB

[AKS^{+}20] Ruhul Amin, Sourav Kunal, Arijit Saha, Debasis Das, and Atif Alamri. CFSec: Password based secure communication protocol in cloud–fog environment. Journal of Parallel and Distributed Computing, 140(??):52–62, June 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520301076.

Alabdulatif:2020:TSB

[AKY20] Abdulatif Alabdulatif, Ibrahim Khalil, and Xun Yi. Towards secure big data analytic for cloud-enabled applications with fully homomorphic encryption. Journal of Parallel and Distributed Computing, 137(??):192–204, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300887.

Alani:2024:HSS

[Ala24] Mohammed M. Alani. HoneyTwin: Securing smart cities with machine learning-enabled SDN edge and cloud-based honeypots. Journal of Parallel and Distributed Computing, 188(??): ??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731524000303.

Almogren:2020:IDE

[Alm20] Ahmad S. Almogren. Intrusion detection in Edge-of-Things computing. Journal of Parallel and Distributed Computing, 137(??):259–265, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373151930872X.

Aggarwal:2021:PSU

[ALP21] Vaneet Aggarwal, Tian Lan, and Dheeraj Peddireddy. Preemptive scheduling on unrelated machines with fractional precedence constraints. Journal of Parallel and Distributed Computing, 157(??):280–286, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521001568.

Alamro:2023:FDC

[ALS23] Sultan Alamro, Tian Lan, and Suresh Subramaniam. Forseti: Dynamic chunk-level reshaping for data processing on heterogeneous clusters. Journal of Parallel and Distributed Computing, 171(??):14–23, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001915.

Abdollahi:2020:VAF

[AM20] Meisam Abdollahi and Siamak Mohammadi. Vulnerability assessment of fault-tolerant optical network-on-chips. Journal of Parallel and Distributed Computing, 145(??):140–159, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303294.

Araki:2022:DLB

[AM22] Samuel J. Araki and Robert S. Martin. Dynamic load balancing with over decomposition in plasma plume simulations. Journal of Parallel and Distributed Computing, 163(??):136–146, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000296.

Affetti:2020:TTS

[AMC20] Lorenzo Affetti, Alessandro Margara, and Gianpaolo Cugola. TSpoon: Transactions on a stream processor. Journal of Parallel and Distributed Computing, 140(??):65–79, June 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305082.

Alibeiki:2021:NGB

[AMM21a] Abolghasem Alibeiki, Homayun Motameni, and Hosein Mohamadi. A new genetic-based approach for solving k-coverage problem in directional sensor networks. Journal of Parallel and Distributed Computing, 154(??):16–26, August 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000599.

Ammari:2021:CCT

[Amm21b] Habib M. Ammari. Connected k-coverage in two-dimensional wireless sensor networks using hexagonal slicing and area stretching. Journal of Parallel and Distributed Computing, 153(??):89–109, July 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373152030424X.

Aldwairi:2020:ESF

[AMP20] Monther Aldwairi, Suaad Mohammed, and Megana Lakshmi Padmanabhan. Efficient and secure flash-based gaming CAPTCHA. Journal of Parallel and Distributed Computing, 142(??):27–35, August 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519308858.

Allendorf:2023:PGE

[AMPT23] Daniel Allendorf, Ulrich Meyer, Manuel Penschuck, and Hung Tran. Parallel global edge switching for the uniform sampling of simple graphs with prescribed degrees. Journal of Parallel and Distributed Computing, 174(??):118–129, April 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002623.

Anonymous:2020:EBa

[Ano20a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 135(??):ii, January 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519307208.

Anonymous:2020:EBb

[Ano20b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 136(??):i, February 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151930783X.

Anonymous:2020:EBc

[Ano20c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 136(??):ii, February 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519307841.

Anonymous:2020:EBd

[Ano20d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 137(??):i, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520300149.

Anonymous:2020:EBe

[Ano20e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 137(??):ii, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520300150.

Anonymous:2020:EBf

[Ano20f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 138(??):i, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520300824.

Anonymous:2020:EBg

[Ano20g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 138(??):ii, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520300836.

Anonymous:2020:EBh

[Ano20h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 139(??):i, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520301623.

Anonymous:2020:EBi

[Ano20i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 139(??):ii, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520301635.

Anonymous:2020:EBj

[Ano20j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 140(??):i, June 2020. CODEN JPDCER.

ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520302264.

Anonymous:2020:EBk

[Ano20k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 140(??):ii, June 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520302276.

Anonymous:2020:EBl

[Ano20l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 141(??):i, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520302641.

Anonymous:2020:EBm

[Ano20m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 141(??):ii, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731520302653.

Anonymous:2020:EBn

[Ano20n] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 142(??):i, August 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S074373152030280X.

Anonymous:2020:EBo

[Ano20o] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 142(??):ii, August 2020. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731520302811.

Anonymous:2020:EBp

[Ano20p] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 143(??):i, September 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303026.

Anonymous:2020:EBq

[Ano20q] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 143(??):ii, September 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303038.

Anonymous:2020:EBr

[Ano20r] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 144(??):ii, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303257.

Anonymous:2020:EBs

[Ano20s] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 145(??):i, November 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303531.

Anonymous:2020:EBt

[Ano20t] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 145(??):ii, November 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303543.

Anonymous:2020:EBu

[Ano20u] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 146(??):i, December 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303786.

Anonymous:2020:EBv

[Ano20v] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 146(??):ii, December 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303798.

Anonymous:2020:FMF

[Ano20w] Anonymous. Front matter 1-full title page (regular issues)/ special issue title page (special issues). Journal of Parallel and Distributed Computing, 144(??):i, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303245.

Anonymous:2020:FCT

[Ano20x] Anonymous. Full [cover] title. Journal of Parallel and Distributed Computing, 135(??):i, January 2020. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731519307191.

Anonymous:2020:PJa

[Ano20y] Anonymous. Pages 1–110 (June 2020). Journal of Parallel and Distributed Computing, 140(??):1–110, June 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2020:PA

[Ano20z] Anonymous. Pages 1–118 (August 2020). Journal of Parallel and Distributed Computing, 142(??):1–118, August 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2020:PJb

[Ano20-27] Anonymous. Pages 1–174 (July 2020). Journal of Parallel and Distributed Computing, 141(??):1–174, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2020:PS

[Ano20-28] Anonymous. Pages 1–178 (September 2020). Journal of Parallel and Distributed Computing, 143(??):1–178, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2020:PD

[Ano20-29] Anonymous. Pages 1–212 (December 2020). Journal of Parallel and Distributed Computing, 146(??):1–212, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2020:PN

[Ano20-30] Anonymous. Pages 1–216 (November 2020). Journal of Parallel and Distributed Computing, 145(??):1–216, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2020:PO

[Ano20-31] Anonymous. Pages 1–294 (October 2020). Journal of Parallel and Distributed Computing, 144(??):1–294, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:EBa

[Ano21a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 147(??):i, January 2021. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731520303944.

Anonymous:2021:EBb

[Ano21b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 147(??):ii, January 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303956.

Anonymous:2021:EBc

[Ano21c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 148(??):i, February 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152030410X.

Anonymous:2021:EBd

[Ano21d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 148(??):ii, February 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304111.

Anonymous:2021:EBe

[Ano21e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 149(??):i, March 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000083.

Anonymous:2021:EBf

[Ano21f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 149(??):ii, March 2021. CODEN JPDCER.

ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000095.

Anonymous:2021:EBg

[Ano21g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 150(??):i, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000162.

Anonymous:2021:EBh

[Ano21h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 150(??):ii, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000174.

Anonymous:2021:EBi

[Ano21i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 151(??):i, May 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000502.

Anonymous:2021:EBj

[Ano21j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 151(??):ii, May 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000514.

Anonymous:2021:EBk

[Ano21k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 152(??):i, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000691.

Anonymous:2021:EBl

[Ano21l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 152(??):ii, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000708.

Anonymous:2021:EBm

[Ano21m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 153(??):i, July 2021. CODEN JPDCER.

ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000861.

Anonymous:2021:EBn

[Ano21n] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 153(??):ii, July 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521000873.

Anonymous:2021:EBo

[Ano21o] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 154(??):i, August 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731521001039.

Anonymous:2021:EBp

[Ano21p] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 154(??):ii, August 2021. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731521001040.

Anonymous:2021:EBq

[Ano21q] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 155(??):i, September 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001180.

Anonymous:2021:EBr

[Ano21r] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 155(??):ii, September 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001192.

Anonymous:2021:EBs

[Ano21s] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 156(??):i, October 2021. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731521001398.

Anonymous:2021:EBt

[Ano21t] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 156(??):ii, October 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001404.

Anonymous:2021:EBu

[Ano21u] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 157(??):i, November 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001738.

Anonymous:2021:EBv

[Ano21v] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 157(??):ii, November 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100174X.

Anonymous:2021:EBw

[Ano21w] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 158(??):i, December 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001878.

Anonymous:2021:EBx

[Ano21x] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 158(??):ii, December 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100188X.

Anonymous:2021:PS

[Ano21y] Anonymous. Pages 1–120 (September 2021). Journal of Parallel and Distributed Computing, 155(??):1–120, September 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PMb

[Ano21z] Anonymous. Pages 1–138 (May 2021). Journal of Parallel and Distributed Computing, 151(??):1–138, May 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PF

[Ano21-27] Anonymous. Pages 1–150 (February 2021). Journal of Parallel and Distributed Computing, 148(??):1–150, February 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PJc

[Ano21-28] Anonymous. Pages 1–160 (July 2021). Journal of Parallel and Distributed Computing, 153(??):1–160, July 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PAb

[Ano21-29] Anonymous. Pages 1–162 (August 2021). Journal of Parallel and Distributed Computing, 154(??):1–162, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PJb

[Ano21-30] Anonymous. Pages 1–176 (June 2021). Journal of Parallel and Distributed Computing, 152(??):1–176, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PO

[Ano21-31] Anonymous. Pages 1–184 (October 2021). Journal of Parallel and Distributed Computing, 156(??):1–184, October 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PAa

[Ano21-32] Anonymous. Pages 1–196 (April 2021). Journal of Parallel and Distributed Computing, 150(??):1–196, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PMa

[Ano21-33] Anonymous. Pages 1–206 (March 2021). Journal of Parallel and Distributed Computing, 149(??):1–206, March 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PD

[Ano21-34] Anonymous. Pages 1–226 (December 2021). Journal of Parallel and Distributed Computing, 158(??):1–226, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PJa

[Ano21-35] Anonymous. Pages 1–270 (January 2021). Journal of Parallel and Distributed Computing, 147(??):1–270, January 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2021:PN

[Ano21-36] Anonymous. Pages 1–330 (November 2021). Journal of Parallel and Distributed Computing, 157(??):1–330, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:EBa

[Ano22a] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 159(??):i, January 2022. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731521001982.

Anonymous:2022:EBb

[Ano22b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 159(??):ii, January 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001994.

Anonymous:2022:EBc

[Ano22c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 160(??):i, February 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002161.

Anonymous:2022:EBd

[Ano22d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 160(??):ii, February 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002173.

Anonymous:2022:EBe

[Ano22e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 161(??):i, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S074373152200003X.

Anonymous:2022:EBf

[Ano22f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 161(??):ii, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522000041.

Anonymous:2022:EBg

[Ano22g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 162(??):i, April 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522000429.

Anonymous:2022:EBh

[Ano22h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 162(??):ii, April 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522000430.

Anonymous:2022:EBi

[Ano22i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 163(??):i, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S074373152200051X.

Anonymous:2022:EBj

[Ano22j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 163(??):ii, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522000521.

Anonymous:2022:EBk

[Ano22k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 164(??):i, June 2022. CODEN JPDCER.

ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522000752.

Anonymous:2022:EBl

[Ano22l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 164(??):ii, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522000764.

Anonymous:2022:EBm

[Ano22m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 165(??):i, July 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522001058.

Anonymous:2022:EBn

[Ano22n] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 165(??):ii, July 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S074373152200106X.

Anonymous:2022:EBo

[Ano22o] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 166(??):i, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522001174.

Anonymous:2022:EBp

[Ano22p] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 166(??):ii, August 2022. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731522001186.

Anonymous:2022:EBq

[Ano22q] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 167(??):i, September 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001368.

Anonymous:2022:EBr

[Ano22r] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 167(??):ii, September 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200137X.

Anonymous:2022:EBs

[Ano22s] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 168(??):i, October 2022. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731522001629.

Anonymous:2022:EBt

[Ano22t] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 168(??):ii, October 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001630.

Anonymous:2022:EBu

[Ano22u] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 169(??):i, November 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001824.

Anonymous:2022:EBv

[Ano22v] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 169(??):ii, November 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001836.

Anonymous:2022:EBw

[Ano22w] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 170(??):i, December 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002052.

Anonymous:2022:EBx

[Ano22x] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 170(??):ii, December 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002064.

Anonymous:2022:PF

[Ano22y] Anonymous. Pages 1–110 (February 2022). Journal of Parallel and Distributed Computing, 160(??):1–110, February 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PAa

[Ano22z] Anonymous. Pages 1–118 (April 2022). Journal of Parallel and Distributed Computing, 162(??):1–118, April 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PO

[Ano22-27] Anonymous. Pages 1–138 (October 2022). Journal of Parallel and Distributed Computing, 168(??):1–138, October 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PMa

[Ano22-28] Anonymous. Pages 1–154 (March 2022). Journal of Parallel and Distributed Computing, 161(??):1–154, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PJc

[Ano22-29] Anonymous. Pages 1–166 (July 2022). Journal of Parallel and Distributed Computing, 165(??):1–166, July 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PAb

[Ano22-30] Anonymous. Pages 1–178 (August 2022). Journal of Parallel and Distributed Computing, 166(??):1–178, August 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PJb

[Ano22-31] Anonymous. Pages 1–190 (June 2022). Journal of Parallel and Distributed Computing, 164(??):1–190, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PS

[Ano22-32] Anonymous. Pages 1–254 (September 2022). Journal of Parallel and Distributed Computing, 167(??):1–254, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PMb

[Ano22-33] Anonymous. Pages 1–312 (May 2022). Journal of Parallel and Distributed Computing, 163(??):1–312, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PN

[Ano22-34] Anonymous. Pages 1–352 (November 2022). Journal of Parallel and Distributed Computing, 169(??):1–352, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PJa

[Ano22-35] Anonymous. Pages 1–84 (January 2022). Journal of Parallel and Distributed Computing, 159(??):1–84, January 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2022:PD

[Ano22-36] Anonymous. Pages 1–86 (December 2022). Journal of Parallel and Distributed Computing, 170(??):1–86, December 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:D

[Ano23a] Anonymous. December 2023. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:EBa

[Ano23b] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 171(??):i, January 2023. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731522002192.

Anonymous:2023:EBb

[Ano23c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 171(??):ii, January 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002209.

Anonymous:2023:EBc

[Ano23d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 172(??):i, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002374.

Anonymous:2023:EBd

[Ano23e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 172(??):ii, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002386.

Anonymous:2023:EBe

[Ano23f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 173(??):i, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731522002568.

Anonymous:2023:EBf

[Ano23g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 173(??):ii, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S074373152200257X.

Anonymous:2023:EBg

[Ano23h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 174(??):i, April 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523000096.

Anonymous:2023:EBh

[Ano23i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 174(??):ii, April 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523000102.

Anonymous:2023:EBi

[Ano23j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 175(??):i, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523000254.

Anonymous:2023:EBj

[Ano23k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 175(??):ii, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523000266.

Anonymous:2023:EBk

[Ano23l] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 176(??):i, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523000461.

Anonymous:2023:EBl

[Ano23m] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 176(??):ii, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523000473.

Anonymous:2023:EBm

[Ano23n] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 177(??):i, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523000618.

Anonymous:2023:EBn

[Ano23o] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 177(??):ii, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S074373152300062X.

Anonymous:2023:EBo

[Ano23p] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 178(??):i, August 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523000783.

Anonymous:2023:EBp

[Ano23q] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 178(??):ii, August 2023. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731523000795.

Anonymous:2023:EBq

[Ano23r] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000953.

Anonymous:2023:EBr

[Ano23s] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000965.

Anonymous:2023:EBs

[Ano23t] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001065.

Anonymous:2023:EBt

[Ano23u] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001077.

Anonymous:2023:EBu

[Ano23v] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001260.

Anonymous:2023:EBv

[Ano23w] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001272.

Anonymous:2023:EBw

[Ano23x] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001417.

Anonymous:2023:EBx

[Ano23y] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001429.

Anonymous:2023:N

[Ano23z] Anonymous. November 2023. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:O

[Ano23-27] Anonymous. October 2023. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:PAb

[Ano23-28] Anonymous. Pages 1–112 (August 2023). Journal of Parallel and Distributed Computing, 178(??):1–112, August 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:PJb

[Ano23-29] Anonymous. Pages 1–128 (June 2023). Journal of Parallel and Distributed Computing, 176(??):1–128, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:PAa

[Ano23-30] Anonymous. Pages 1–130 (April 2023). Journal of Parallel and Distributed Computing, 174(??):1–130, April 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:PMb

[Ano23-31] Anonymous. Pages 1–150 (May 2023). Journal of Parallel and Distributed Computing, 175(??):1–150, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:PF

[Ano23-32] Anonymous. Pages 1–158 (February 2023). Journal of Parallel and Distributed Computing, 172(??):1–158, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:PJa

[Ano23-33] Anonymous. Pages 1–162 (January 2023). Journal of Parallel and Distributed Computing, 171(??):1–162, January 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:PMa

[Ano23-34] Anonymous. Pages 1–168 (March 2023). Journal of Parallel and Distributed Computing, 173(??):1–168, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:PJc

[Ano23-35] Anonymous. Pages 1–192 (July 2023). Journal of Parallel and Distributed Computing, 177(??):1–192, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2023:S

[Ano23-36] Anonymous. September 2023. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:Aa

[Ano24a] Anonymous. April 2024. Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:Ab

[Ano24b] Anonymous. August 2024. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:EBa

[Ano24c] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001569.

Anonymous:2024:EBb

[Ano24d] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001776.

Anonymous:2024:EBc

[Ano24e] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.com/science/article/pii/S0743731523001958.

Anonymous:2024:EBd

[Ano24f] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S074373152400008X.

Anonymous:2024:EBe

[Ano24g] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731524000248.

Anonymous:2024:EBf

[Ano24h] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S074373152400039X.

Anonymous:2024:EBg

[Ano24i] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CODEN JPDCER.

ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731524000595.

Anonymous:2024:EBh

[Ano24j] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000765.

Anonymous:2024:EBi

[Ano24k] Anonymous. Editorial Board. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000959.

Anonymous:2024:F

[Ano24l] Anonymous. February 2024. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:FMFa

[Ano24m] Anonymous. Front matter 1 — full title page (regular issues)/special issue title page (special issues). Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001557.

Anonymous:2024:FMFc

[Ano24n] Anonymous. Front matter 1 — full title page (regular issues)/special issue title page (special issues). Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001946.

Anonymous:2024:FMFd

[Ano24o] Anonymous. Front matter 1 — full title page (regular issues)/special issue title page (special issues). Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000078.

Anonymous:2024:FMFe

[Ano24p] Anonymous. Front matter 1 — full title page (regular issues)/special issue title page (special issues). Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000236.

Anonymous:2024:FMFf

[Ano24q] Anonymous. Front matter 1 — full title page (regular issues)/special issue title page (special issues). Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000388.

Anonymous:2024:FMFg

[Ano24r] Anonymous. Front matter 1 — full title page (regular issues)/special issue title page (special issues). Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000583.

Anonymous:2024:FMFh

[Ano24s] Anonymous. Front matter 1 — full title page (regular issues)/special issue title page (special issues). Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000753.

Anonymous:2024:FMFi

[Ano24t] Anonymous. Front matter 1 — full title page (regular issues)/special issue title page (special issues). Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000947.

Anonymous:2024:FMFb

[Ano24u] Anonymous. Front matter [1-Full Title Page (regular issues)/Special Issue Title page (special issues)]. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001764.

Anonymous:2024:Ja

[Ano24v] Anonymous. January 2024. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:Jc

[Ano24w] Anonymous. July 2024. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:Jb

[Ano24x] Anonymous. June 2024. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:Ma

[Ano24y] Anonymous. March 2024. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:Mb

[Ano24z] Anonymous. May 2024. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Anonymous:2024:RTY

[Ano24-27] Anonymous. Reviewer thank you. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001788.

Anonymous:2024:S

[Ano24-28] Anonymous. September 2024. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Arbenz:2020:BTF

[AR20] Peter Arbenz and Lubomír Ríha. Batched transpose-free ADItype preconditioners for a Poisson solver on GPGPUs. Journal of Parallel and Distributed Computing, 137(??):148–159, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519307609.

Areias:2021:CEN

[AR21] Miguel Areias and Ricardo Rocha. On the correctness and efficiency of a novel lock-free hash trie map design. Journal of Parallel and Distributed Computing, 150(??):184–195, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000010.

Alcaraz:2020:BAA

[ARL20] Cristina Alcaraz, Juan E. Rubio, and Javier Lopez. Blockchainassisted access for federated smart grid domains: Coupling and features. Journal of Parallel and Distributed Computing, 144(??):124–135, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520302914.

Amiri:2020:SPU

[AS20] Hossein Amiri and Asadollah Shahbahrami. SIMD programming using Intel vector extensions. Journal of Parallel and Distributed Computing, 135(??):83–100, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830813X.

Augustine:2021:SSR

[AS21] John Augustine and Sumathi Sivasubramaniam. Spartan: Sparse robust addressable networks. Journal of Parallel and Distributed Computing, 150(??):121–138, April 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304299.

Al-Sayed:2020:COS

[ASHO20] Mustafa M. Al-Sayed, Hesham A. Hassan, and Fatma A. Omara. CloudFNF: an ontology structure for functional and non-functional features of cloud services. Journal of Parallel and Distributed Computing, 141(??):143–173, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300437.

Ansotegui:2024:ECH

[AT24] Carlos Ansótegui and Eduard Torres. Effectively computing high strength mixed covering arrays with constraints. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001612.

Abdelfattah:2020:MMB

[ATD20] Ahmad Abdelfattah, Stanimire Tomov, and Jack Dongarra. Matrix multiplication on batches of small matrices in half and half-complex precisions. Journal of Parallel and Distributed Computing, 145(??):188–201, November 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303300.

Atighehchi:2020:PNA

[Ati20] Kevin Atighehchi. A precise non-asymptotic complexity analysis of parallel hash functions without tree topology constraints. Journal of Parallel and Distributed Computing, 137(??):246– 251, March 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518308682.

Atoofian:2023:PPA

[Ato23] Ehsan Atoofian. PTTS: Power-aware tensor cores using twosided sparsity. Journal of Parallel and Distributed Computing, 173(??):70–82, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522002325.

Ajmi:2024:ELM

[AZF^{+}24] Hala Ajmi, Fakhreddine Zayer, Amira Hadj Fredj, Hamdi Belgacem, Baker Mohammad, Naoufel Werghi, and Jorge Dias. Efficient and lightweight in-memory computing architecture for hardware security. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000625.

Boulmier:2022:OLB

[BAC22] Anthony Boulmier, Nabil Abdennadher, and Bastien Chopard. Optimal load balancing and assessment of existing load balancing criteria. Journal of Parallel and Distributed Computing, 169(??):211–225, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373152200168X.

Batista:2022:ESS

[BADP22] Eliã Batista, Eduardo Alchieri, Fernando Dotti, and Fernando Pedone. Early scheduling on steroids: Boosting parallel state machine replication. Journal of Parallel and Distributed Computing, 163(??):269–282, May 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000375.

Berenbrink:2021:RRS

[BBE^{+}21] Petra Berenbrink, André Brinkmann, Robert Elsässer, Tom Friedetzky, and Lars Nagel. Randomized renaming in shared memory systems. Journal of Parallel and Distributed Computing, 150(??):112–120, April 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000022.

Benoit:2023:LSS

[BCEH23] Anne Benoit, Louis-Claude Canon, Redouane Elghazi, and Pierre-Cyrille Héam. List and shelf schedules for independent parallel tasks to minimize the energy consumption with discrete or continuous speeds. Journal of Parallel and Distributed Computing, 174(??):100–117, April 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002489.

Baert:2021:AMA

[BCM^{+}21] Quentin Baert, Anne-Cécile Caron, Maxime Morge, Jean-Christophe Routier, and Kostas Stathis. An adaptive multiagent system for task reallocation in a MapReduce job. Journal of Parallel and Distributed Computing, 153(??):75–88, July 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000617.

Bose:2023:MVF

[BCM23] Kaustav Bose, Abhinav Chakraborty, and Krishnendu Mukhopadhyaya. Mutual visibility by fat robots with slim omnidirectional camera. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523000862.

Bodei:2021:MAI

[BDFG21] Chiara Bodei, Pierpaolo Degano, Gian-Luigi Ferrari, and Letterio Galletta. Modelling and analysing IoT systems. Journal of Parallel and Distributed Computing, 157(??):233–242, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001507.

Barbosa:2022:SIC

[BDL22] Jorge G. Barbosa, Lúcia M. A. Drummond, and Laurent Lefèvre. Special issue on computer architecture and high-performance computing. Journal of Parallel and Distributed Computing, 168(??):137–138, October 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001484.

Breitenbach:2024:MDR

[BDRJ24] Tim Breitenbach, Shrikanth Malavalli Divakar, Lauritz Rasbach, and Patrick Jahnke. ML-driven risk estimation for memory failure in a data center environment with convolutional neural networks, self-supervised data labeling and distribution-based model drift determination. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001703.

Barrachina:2022:EPG

[BDSQO22] Sergio Barrachina, Manuel F. Dolz, Pablo San Juan, and Enrique S. Quintana-Ortí. Efficient and portable GEMMbased convolution operators for deep neural network training on multicore processors. Journal of Parallel and Distributed Computing, 167(??):240–254, September 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001241.

Borisenko:2021:EGP

[BG21] Andrey Borisenko and Sergei Gorlatch. Efficient GPUparallelization of batch plants design using metaheuristics with parameter tuning. Journal of Parallel and Distributed Computing, 154(??):74–81, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000654.

Baker:2021:ETE

[BGA^{+}21] Thar Baker, Zehua Guo, Ali Ismail Awad, Shangguang Wang, and Benjamin C. M. Fung. Enabling technologies for energy cloud. Journal of Parallel and Distributed Computing, 152(??):108–110, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521000411.

Beceiro:2024:CAM

[BGDMF^{+}24] Bieito Beceiro, Jorge González-Domínguez, Laura Morán-Fernández, Verónica Bolón-Canedo, and Juan Touriño. CUDA acceleration of MI-based feature selection methods. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000650.

Beceiro:2022:PFF

[BGDT22] Bieito Beceiro, Jorge González-Domínguez, and Juan Touriño. Parallel-FST: a feature selection library for multicore clusters. Journal of Parallel and Distributed Computing, 169 (??):106–116, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001472.

Brunn:2021:FGD

[BHB^{+}21] Malte Brunn, Naveen Himthani, George Biros, Miriam Mehl, and Andreas Mang. Fast GPU 3D diffeomorphic image registration. Journal of Parallel and Distributed Computing, 149(??):149–162, March 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152030407X.

Benoit:2024:RBS

[BHP^{+}24] Anne Benoit, Thomas Herault, Lucas Perotin, Yves Robert, and Frédéric Vivien. Revisiting I/O bandwidth-sharing strategies for HPC applications. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731524000273.

Bermejo:2023:ICE

[BJ23] Belen Bermejo and Carlos Juiz. Improving cloud/edge sustainability through artificial intelligence: a systematic review. Journal of Parallel and Distributed Computing, 176(??):41–54, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000217.

Bhattarai:2023:DNN

[BkB^{+}23] Manish Bhattarai, Namita kharat, Ismael Boureima, Erik Skau, Benjamin Nebgen, Hristo Djidjev, Sanjay Rajopadhye, James P. Smith, and Boian Alexandrov. Distributed non-negative RESCAL with automatic model selection for exascale data. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000710.

Belguith:2020:APP

[BKL^{+}20] Sana Belguith, Nesrine Kaaniche, Maryline Laurent, Abderrazak Jemai, and Rabah Attia. Accountable privacy preserving attribute based framework for authenticated encrypted access in clouds. Journal of Parallel and Distributed Computing, 135(??):1–20, January 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302175.

Bednarek:2021:LFP

[BKY21] David Bednárek, Martin Kruliš, and Jakub Yaghob. Letting future programmers experience performance-related tasks. Journal of Parallel and Distributed Computing, 155(??):74–86, September 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000976.

Barenboim:2023:SDA

[BL23] Leonid Barenboim and Harel Levin. Secured distributed algorithms without hardness assumptions. Journal of Parallel and Distributed Computing, 171(??):130–140, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002003.

Biguri:2020:ALT

[BLB^{+}20] Ander Biguri, Reuben Lindroos, Robert Bryll, Hossein Towsyfyan, Hans Deyhle, Ibrahim El khalil Harrane, Richard Boardman, Mark Mavrogordato, Manjit Dosanjh, Steven Hancock, and Thomas Blumensath. Arbitrarily large tomography with iterative algorithms on multiple GPUs using the TIGRE toolbox. Journal of Parallel and Distributed Computing, 146(??):52–63, December 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303336.

Bernardinello:2023:SPC

[BLNP23] Luca Bernardinello, Irina Lomazova, Roman Nesterov, and Lucia Pomello. Soundness-preserving composition of synchronously and asynchronously interacting workflow net components. Journal of Parallel and Distributed Computing, 179 (??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523000679.

Bhattacharya:2024:NMC

[BMC^{+}24] Parantapa Bhattacharya, Dustin Machi, Jiangzhuo Chen, Stefan Hoops, Bryan Lewis, Henning Mortveit, Srinivasan Venkatramanan, Mandy L. Wilson, Achla Marathe, Przemyslaw Porebski, Brian Klahn, Joseph Outten, Anil Vullikanti, Dawen Xie, Abhijin Adiga, Shawn Brown, Christopher Barrett, and Madhav Marathe. Novel multi-cluster workflow system to support real-time HPC-enabled epidemic science: Investigating the impact of vaccine acceptance on COVID-19 spread. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000637.

Barreiros:2022:EMI

[BMK^{+}22] Willian Barreiros, Alba C. M. A. Melo, Jun Kong, Renato Ferreira, Tahsin M. Kurc, Joel H. Saltz, and George Teodoro. Efficient microscopy image analysis on CPU–GPU systems with cost-aware irregular data partitioning. Journal of Parallel and Distributed Computing, 164(??):40–54, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000466.

Barreales:2021:MSM

[BNOS21] Gonzalo Nicolas Barreales, Marcos Novalbos, Miguel A. Otaduy, and Alberto Sanchez. MDScale: Scalable multi-GPU bonded and short-range molecular dynamics. Journal of Parallel and Distributed Computing, 157(??):243–255, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001520.

Belkhiri:2023:PAD

[BPBD23] Adel Belkhiri, Martin Pepin, Mike Bly, and Michel Dagenais. Performance analysis of DPDK-based applications through tracing. Journal of Parallel and Distributed Computing, 173(??): 1–19, March 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522002271.

Bhattacharya:2022:MMM

[BPT^{+}22] Pronaya Bhattacharya, Farnazbanu Patel, Sudeep Tanwar, Neeraj Kumar, and Ravi Sharma. MB-MaaS: Mobile blockchainbased mining-as-a-service for IIoT environments. Journal of Parallel and Distributed Computing, 168(??):1–16, October 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001228.

Bhattacharjee:2021:SFP

[BRK^{+}21] Raktim Bhattacharjee, R. Rajesh, K. R. Prasanna Kumar, Vinu Paul MV, G. Athithan, and A. V. Sahadevan. Scalable flow probe architecture for 100 Gbps+ rates on commodity hardware: Design considerations and approach. Journal of Parallel and Distributed Computing, 155(??):87–100, September 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000988.

Behera:2024:TSO

[BS24] Ipsita Behera and Srichandan Sobhanayak. Task scheduling optimization in heterogeneous cloud computing environments: a hybrid GA-GWO approach. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001363.

Bratek:2023:REC

[BSWO23] Pawel Bratek, Lukasz Szustak, Roman Wyrzykowski, and Tomasz Olas. Reducing energy consumption using heterogeneous voltage frequency scaling of data-parallel applications for multicore systems. Journal of Parallel and Distributed Computing, 175(??):121–133, May 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000059.

Barani:2024:DLB

[BSY24] Fatemeh Barani, Abdorreza Savadi, and Hadi Sadoghi Yazdi. A distributed learning based on robust diffusion SGD over adaptive networks with noisy output data. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000479.

Blin:2020:CSS

[BT20] Lélia Blin and Sébastien Tixeuil. Compact self-stabilizing leader election for general networks. Journal of Parallel and Distributed Computing, 144(??):278–294, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302987.

Bhaskar:2021:PAN

[BV21] A. Vijaya Bhaskar and T. G. Venkatesh. Performance analysis of network-on-chip in many-core processors. Journal of Parallel and Distributed Computing, 147(??):196–208, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303749.

Bao:2022:EES

[BW22] Han Bao and Yijie Wang. ESDU: an elastic stripebased delta update method for erasure-coded cross-data center storage systems. Journal of Parallel and Distributed Computing, 167(??):173–186, September 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001125.

Bao:2022:XSQ

[BYW^{+}22] Liang Bao, Jin Yang, Chase Q. Wu, Haiyang Qi, Xin Zhang, and Shunda Cai. XML2HBase: Storing and querying large collections of XML documents using a NoSQL database system. Journal of Parallel and Distributed Computing, 161(??): 83–99, March 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521002100.

Cholvi:2023:AAA

[CAG^{+}23] Vicent Cholvi, Antonio Fernández Anta, Chryssis Georgiou, Nicolas Nicolaou, Michel Raynal, and Antonio Russo. Atomic appends in asynchronous Byzantine distributed ledgers. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001181.

Chen:2024:DOL

[CaTZ^{+}24] Tian Chen, Yu an Tan, Zheng Zhang, Nan Luo, Bin Li, and Yuanzhang Li. Dataflow optimization with layer-wise design variables estimation method for enflame CNN accelerators. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000339.

Carneiro:2023:UDH

[CBO^{+}23] André Ramos Carneiro, Jean Luca Bez, Carla Osthoff, Lucas Mello Schnorr, and Philippe O. A. Navaux. Uncovering I/O demands on HPC platforms: Peeking under the hood of Santos Dumont. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001144.

Chunduri:2023:DTW

[CC23] Raghavendra Kumar Chunduri and Aswani Kumar Cherukuri. Distributed three-way formal concept analysis for large formal contexts. Journal of Parallel and Distributed Computing, 171(??):141–156, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522002015.

Cano-Cano:2021:QPH

[CCAACS21] Javier Cano-Cano, Francisco J. Andújar, Francisco J. Alfaro-Cortés, and José L. Sánchez. QoS provision in hierarchical and non-hierarchical switch architectures. Journal of Parallel and Distributed Computing, 148(??):138–150, February 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303981.

Chen:2023:PPD

[CCC23] Chen-Chun Chen, Yu-Min Chou, and Jerry Chou. PHY: a performance-driven hybrid communication compression method for distributed training. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000898.

Catalan:2021:LTD

[CCSI21] Sandra Catalán, Rocío Carratalá-Sáez, and Sergio Iserte. Leveraging teaching on demand: Approaching HPC to undergrads. Journal of Parallel and Distributed Computing, 156(??):148– 162, October 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521001271.

Cheriere:2020:HFC

[CDA20] Nathanaël Cheriere, Matthieu Dorier, and Gabriel Antoniu. How fast can one resize a distributed file system? Journal of Parallel and Distributed Computing, 140(??):80–98, June 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305008.

Canonne:2024:LSC

[CDTS24] Lorenzo Canonne, Bilel Derbel, Miwako Tsuji, and Mitsuhisa Sato. Large-scale and cooperative graybox parallel optimization on the supercomputer Fugaku. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000856.

Chen:2023:NCG

[CDY23] Jianguo Chen, Qingying Deng, and Xulei Yang. Noncooperative game algorithms for computation offloading in mobile edge computing environments. Journal of Parallel and Distributed Computing, 172(??):18–31, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200212X.

Carrasco:2024:EGF

[CFNH24] Roberto Carrasco, Héctor Ferrada, Cristóbal A. Navarro, and Nancy Hitschfeld. An evaluation of GPU filters for accelerating the 2D convex hull. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001636.

Castano:2022:EID

[CFRGPM22] Germán Castaño, Youssef Faqir-Rhazoui, Carlos García, and Manuel Prieto-Matías. Evaluation of Intel’s DPC++ compatibility tool in heterogeneous computing. Journal of Parallel and Distributed Computing, 165(??):120–129, July 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000727.

Choo:2021:BES

[CGC21] Kim-Kwang Raymond Choo, Keke Gai, and Luca Chiaraviglio. Blockchain-enabled secure communications in smart cities. Journal of Parallel and Distributed Computing, 152(??):125– 127, June 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000423.

Chahal:2020:HGD

[CGDS20] Karanbir Singh Chahal, Manraj Singh Grover, Kuntal Dey, and Rajiv Ratn Shah. A hitchhiker’s guide on distributed training of deep neural networks. Journal of Parallel and Distributed Computing, 137(??):65–76, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308712.

Civit:2023:EAO

[CGG^{+}23] Pierre Civit, Seth Gilbert, Vincent Gramoli, Rachid Guerraoui, and Jovan Komatovic. As easy as ABC: Optimal (A)ccountable (B)yzantine (C)onsensus is easy! Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001132.

Choi:2020:LBD

[CGK20] Jinhwan Choi, Yu Gu, and Jinoh Kim. Learning-based dynamic cache management in a cloud. Journal of Parallel and Distributed Computing, 145(??):98–110, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303208.

Chen:2022:NNN

[CGL^{+}22] Xinhai Chen, Chunye Gong, Jie Liu, Yufei Pang, Liang Deng, Lihua Chi, and Kenli Li. A novel neural network approach for airfoil mesh quality evaluation. Journal of Parallel and Distributed Computing, 164(??):123–132, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000594.

Czappa:2023:SSP

[CGW23] Fabian Czappa, Alexander Geiß, and Felix Wolf. Simulating structural plasticity of the brain more scalable than expected. Journal of Parallel and Distributed Computing, 171(??):24–27, January 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522001927.

Chuang:2023:RTA

[CH23] Yung-Ting Chuang and Yuan-Tsang Hung. A real-time and ACO-based offloading algorithm in edge computing. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000667.

inodChandra:2024:MSD

[CH24] S. S. Vīnod Chandra and S. Anand Hareendran. Modified smell detection algorithm for optimal paths engineering in hybrid SDN. Journal of Parallel and Distributed Computing, 187(??): ??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523002046.

Chen:2023:SCM

[Che23] Chi-Yeh Chen. Scheduling coflows for minimizing the total weighted completion time in heterogeneous parallel networks. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001223.

Cheng:2024:ESA

[CHJ^{+}24] Yongli Cheng, Chuanjie Huang, Hong Jiang, Xianghao Xu, and Fang Wang. An efficient SSSP algorithm on timeevolving graphs with prediction of computation results. Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523002009.

Catalan:2023:PPD

[CIH^{+}23] Sandra Catalán, Francisco D. Igual, José R. Herrero, Rafael Rodríguez-Sánchez, and Enrique S. Quintana-Ortí. Programming parallel dense matrix factorizations and inversion for new-generation NUMA architectures. Journal of Parallel and Distributed Computing, 175(??):51–65, May 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000047.

Chen:2022:EIE

[CJZ^{+}22] Chen Chen, Jiange Jiang, Yang Zhou, Ning Lv, Xiaoxu Liang, and Shaohua Wan. An edge intelligence empowered flooding process prediction using Internet of Things in smart city. Journal of Parallel and Distributed Computing, 165(??):66–78, July 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000612.

Cerin:2022:DSC

[CKS22] Christophe Cérin, Keiji Kimura, and Mamadou Sow. Data stream clustering for low-cost machines. Journal of Parallel and Distributed Computing, 166(??):57–70, August 2022.

CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000831.

Cornebize:2022:SBO

[CL22a] Tom Cornebize and Arnaud Legrand. Simulation-based optimization and sensibility analysis of MPI applications: Variability matters. Journal of Parallel and Distributed Computing, 166(??):111–125, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522000806.

Coutinho:2022:GES

[CL22b] Rodolfo W. L. Coutinho and Frank Y. Li. Guest editorial: Smart mobility management and 5G/beyond 5G (B5G) wireless access. Journal of Parallel and Distributed Computing, 166(??):147–148, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522000958.

Chang:2023:MDS

[CLLM23] Qiong Chang, Xiang Li, Yun Li, and Jun Miyazaki. Multidirectional Sobel operator kernel on GPUs. Journal of Parallel and Distributed Computing, 177(??):160–170, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000400.

Chen:2022:CSA

[CLMH22] Siyi Chen, Jin Liu, Fengchao Ma, and Huixian Huang. Customer-satisfaction-aware and deadline-constrained profit maximization problem in cloud computing. Journal of Parallel and Distributed Computing, 163(??):198–213, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000399.

Cao:2020:BBA

[CLT^{+}20] Huikang Cao, Ruixuan Li, Wenlong Tian, Zhiyong Xu, and Weijun Xiao. Blockchain-based accountability for multi-party oblivious RAM. Journal of Parallel and Distributed Computing, 137(??):224–237, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303387.

Chang:2023:GGB

[CLW^{+}23a] Tao Chang, Li Li, MeiHan Wu, Wei Yu, Xiaodong Wang, and ChengZhong Xu. GraphCS: Graph-based client selection for heterogeneity in federated learning. Journal of Parallel and Distributed Computing, 177(??):131–143, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000394.

Chang:2023:PPA

[CLW^{+}23b] Tao Chang, Li Li, MeiHan Wu, Wei Yu, Xiaodong Wang, and ChengZhong Xu. PAGroup: Privacy-aware grouping framework for high-performance federated learning. Journal of Parallel and Distributed Computing, 175(??):37–50, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002635.

Cheng:2020:WSW

[CLZ20] Daning Cheng, Shigang Li, and Yunquan Zhang. WP-SGD: Weighted parallel SGD for distributed unbalancedworkload training system. Journal of Parallel and Distributed Computing, 145(??):202–216, November 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152030318X.

Cheng:2022:LBG

[CLZ^{+}22] Wen Cheng, Mi Luo, Lingfang Zeng, Yang Wang, and André Brinkmann. Lifespan-based garbage collection to improve SSD’s reliability and performance. Journal of Parallel and Distributed Computing, 164(??):28–39, June 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200048X.

Ciesko:2020:HAT

[CMFV^{+}20] Jan Ciesko, Pedro J. Martínez-Ferrer, Raúl Peñacoba Veigas, Xavier Teruel, and Vicenç Beltran. HDOT — an approach towards productive programming of hybrid applications. Journal of Parallel and Distributed Computing, 137(??):104–118, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305318.

Chen:2024:PCO

[CML^{+}24] Fei Chen, Fengming Meng, Zhipeng Li, Li Li, and Tao Xiang. Public cloud object storage auditing: Design, implementation, and analysis. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731524000340.

Chisholm:2020:IGN

[CMR20] Robert Chisholm, Steve Maddock, and Paul Richmond. Improved GPU near neighbours performance for multi-agent simulations. Journal of Parallel and Distributed Computing, 137(??): 53–64, March 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519301340.

Chidume:2022:IUC

[CN22] Chidiebere Sunday Chidume and Christantus O. Nnamani. Intelligent user-collaborative edge device APC-based MEC 5G IoT for computational offloading and resource allocation. Journal of Parallel and Distributed Computing, 169(??):286–300, November 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522001721.

Constantinescu:2020:PED

[CNFMA20] Denisa-Andreea Constantinescu, Angeles Navarro, Juan-Antonio Fernández-Madrigal, and Rafael Asenjo. Performance evaluation of decision making under uncertainty for low power heterogeneous platforms. Journal of Parallel and Distributed Computing, 137(??):119–133, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307585.

Castaneda:2024:RWF

[CP24] Armando Castañeda and Miguel Piña. Read/write fencefree work-stealing with multiplicity. Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001867.

Cao:2020:PAR

[CPZ^{+}20] Ting Cao, Xiaopu Peng, Chaowei Zhang, Taha Khalid Al Tekreeti, Jianzhou Mao, Xiao Qin, and Jianzhong Huang. A popularity-aware reconstruction technique in erasure-coded storage systems. Journal of Parallel and Distributed Computing, 146(??):122–138, December 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303439.

Chinnakkonda:2022:ASE

[CRS22] Diyanesh Chinnakkonda, Karthick Rajamani, and M. B. Srinivas. Architecture slack exploitation for phase classification and performance estimation in server-class processors. Journal of Parallel and Distributed Computing, 169(??):157–170, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001526.

Chan:2021:DAC

[CSS21] T.-H. Hubert Chan, Mauro Sozio, and Bintao Sun. Distributed approximate k-core decomposition and min-max edge orientation: Breaking the diameter barrier. Journal of Parallel and Distributed Computing, 147(??):87–99, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303506.

Chen:2023:AMC

[CSS^{+}23] Jou-An Chen, Hsin-Hsuan Sung, Xipeng Shen, Nathan Tallent, Kevin Barker, and Ang Li. Accelerating matrix-centric graph processing on GPUs through bit-level optimizations. Journal of Parallel and Distributed Computing, 177(??):53–67, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000357.

Cao:2024:NHA

[CSY^{+}24] Zijian Cao, Qiao Sun, Wenhao Yang, Changcheng Song, Zhe Wang, and Huiyuan Li. A novel HPL-AI approach for FP16-only accelerator and its instantiation on Kunpeng + Ascend AI-specific platform. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000480.

Chockalingam:2020:SPA

[CTA20] Sriram P. Chockalingam, Sharma V. Thankachan, and Srinivas Aluru. Sequential and parallel algorithms for all-pair kmismatch maximal common substrings. Journal of Parallel and Distributed Computing, 144(??):68–79, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302975.

Chen:2022:WAS

[CTFW22] Yu Chen, Wei Tong, Dan Feng, and Zike Wang. Workloadaware storage policies for cloud object storage. Journal of Parallel and Distributed Computing, 163(??):232–247, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000326.

Chenli:2022:PNB

[CTGJ22] Changhao Chenli, Wenyi Tang, Frank Gomulka, and Taeho Jung. ProvNet: Networked bi-directional blockchain for data sharing with verifiable provenance. Journal of Parallel and Distributed Computing, 166(??):32–44, August 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000818.

Casanova:2021:TPD

[CTKdS21] Henri Casanova, Ryan Tanaka, William Koch, and Rafael Ferreira da Silva. Teaching parallel and distributed computing concepts in simulation with WRENCH. Journal of Parallel and Distributed Computing, 156(??):53–63, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001143.

Chen:2020:MLP

[CW20] Huan Chen and Yijie Wang. MiniChain: a lightweight protocol to combat the UTXO growth in public blockchain. Journal of Parallel and Distributed Computing, 143(??):67–76, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302732.

Chen:2021:JCR

[CW21a] Yang Chen and Jie Wu. Joint coflow routing and scheduling in leaf-spine data centers. Journal of Parallel and Distributed Computing, 148(??):83–95, February 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303683.

Chu:2021:GOP

[CW21b] Dongliang Chu and Chase Q. Wu. Generalizing the over operator for parallelization and order-independency. Journal of Parallel and Distributed Computing, 151(??):52–60, May 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000228.

Chen:2022:AAL

[CWHC22] Chen-Chun Chen, Kai-Siang Wang, Yu-Tung Hsiao, and Jerry Chou. ALBERT: an automatic learning based execution and resource management system for optimizing Hadoop workload in clouds. Journal of Parallel and Distributed Computing, 168(??):45–56, October 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001289.

Cheng:2021:EPP

[CYWL21] Guoli Cheng, Shi Ying, Bingming Wang, and Yuhang Li. Efficient performance prediction for Apache Spark. Journal of Parallel and Distributed Computing, 149(??):40–51, March 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303993.

Cajahuaringa:2024:IMC

[CZR^{+}24] Samuel Cajahuaringa, Leandro N. Zanotto, Sandro Rigo, Hervé Yviquel, Munir S. Skaf, and Guido Araujo. Ionmolecule collision cross-section calculations using trajectory parallelization in distributed systems. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000662.

Datta:2024:ETE

[DAG24] Somoshree Datta, Sourav Kanti Addya, and Soumya K. Ghosh. ESMA: Towards elevating system happiness in a decentralized serverless edge computing framework. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001326.

deAssuncao:2021:PSI

[dARR21] Marcos Dias de Assuncao, Eduardo Rocha Rodrigues, and Bruno Raffin. Preface — special issue advances on high performance computing for artificial intelligence. Journal of Parallel and Distributed Computing, 156(??):131, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001313.

Dhamal:2022:SID

[DBAC^{+}22] Swapnil Dhamal, Walid Ben-Ameur, Tijani Chahed, Eitan Altman, Albert Sunny, and Sudheer Poojary. Strategic investments in distributed computing: a stochastic game perspective. Journal of Parallel and Distributed Computing, 169(??):317–333, November 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522001770.

Dai:2023:DMP

[DCM^{+}23] Guowei Dai, Yannan Chen, Yaping Mao, Dachuan Xu, Xiaoyan Zhang, and Zan-Bo Zhang. A distributed message passing algorithm for computing perfect demand matching. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000692.

Dai:2024:EPP

[DDC^{+}24] Zhe Dai, Liang Deng, YongGang Che, Ming Li, Jian Zhang, and Yueqing Wang. Evaluating performance portability of five shared-memory programming models using a high-order unstructured CFD solver. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523002010.

Devaraj:2020:HFI

[DED^{+}20] A. Francis Saviour Devaraj, Mohamed Elhoseny, S. Dhanasekaran, E. Laxmi Lydia, and K. Shankar. Hybridization of firefly and improved multi-objective particle swarm optimization algorithm for energy efficient load balancing in cloud computing environments. Journal of Parallel and Distributed Computing, 142(??):36–45, August 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520300459.

Dufrechou:2021:MLO

[DEFQO21] Ernesto Dufrechou, Pablo Ezzatti, Manuel Freire, and Enrique S. Quintana-Ortí. Machine learning for optimal selection of sparse triangular system solvers on GPUs. Journal of Parallel and Distributed Computing, 158(??):47–55, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001593.

Deb:2021:IBD

[DF21] Debzani Deb and Muztaba Fuad. Integrating big data and cloud computing topics into the computing curricula: a modular approach. Journal of Parallel and Distributed Computing, 157(??):303–315, November 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001581.

DAngelo:2022:APD

[DF22] Gabriele D’Angelo and Stefano Ferretti. Adaptive parallel and distributed simulation of complex networks. Journal of Parallel and Distributed Computing, 163(??):30–44, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000284.

Ding:2023:JJM

[DFL^{+}23] Yi Ding, Weiwei Fang, Mengran Liu, Meng Wang, Yusong Cheng, and Naixue Xiong. JMDC: a joint model and data compression system for deep neural networks collaborative computing in edge-cloud networks. Journal of Parallel and Distributed Computing, 173(??):83–93, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002416.

DAngelo:2020:MDM

[DFP20] Gianni D’Angelo, Massimo Ficco, and Francesco Palmieri. Malware detection in mobile environments based on Autoencoders and API-images. Journal of Parallel and Distributed Computing, 137(??):26–33, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519302436.

Daggitt:2022:DAI

[DG22] Matthew L. Daggitt and Timothy G. Griffin. Dynamic asynchronous iterations. Journal of Parallel and Distributed Computing, 164(??):168–177, June 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000685.

Dai:2024:SFG

[DGA^{+}24] Lingfei Dai, Luqi Gong, Zhulin An, Yongjun Xu, and Boyu Diao. Sketch-fusion: a gradient compression method with multi-layer fusion for communication-efficient distributed training. Journal of Parallel and Distributed Computing, 185(??): ??, March 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001818.

Delporte-Gallet:2021:WIF

[DGFR21] Carole Delporte-Gallet, Hugues Fauconnier, and Michel Raynal. On the weakest information on failures to solve mutual exclusion and consensus in asynchronous crash-prone read/write systems. Journal of Parallel and Distributed Computing, 153(??): 110–118, July 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000745.

Denninnart:2020:ETP

[DGMS20] Chavit Denninnart, James Gentry, Ali Mokhtari, and Mohsen Amini Salehi. Efficient task pruning mechanism to improve robustness of heterogeneous computing systems. Journal of Parallel and Distributed Computing, 142(??):46–61, August 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519307750.

Dewan:2021:TTV

[DGWD21] Prasun Dewan, Samuel George, Andrew Wortas, and Justin Do. Techniques and tools for visually introducing freshmen to object-based thread abstractions. Journal of Parallel and Distributed Computing, 157(??):179–200, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001258.

Diep:2023:GAS

[DHF23] Thanh-Dang Diep, Phuong Hoai Ha, and Karl Fürlinger. A general approach for supporting nonblocking data structures on distributed-memory systems. Journal of Parallel and Distributed Computing, 173(??):48–60, March 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002337.

Dogani:2024:PAS

[DK24] Javad Dogani and Farshad Khunjush. Proactive auto-scaling technique for web applications in container-based edge computing using federated learning model. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000017.

Dua:2021:PLH

[DKS21] Yaman Dua, Vinod Kumar, and Ravi Shankar Singh. Parallel lossless HSI compression based on RLS filter. Journal of Parallel and Distributed Computing, 150(??):60–68, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304202.

Ding:2021:SLT

[DLL^{+}21] Yan Ding, Kenli Li, Chubo Liu, Zhuo Tang, and Keqin Li. Short- and long-term cost and performance optimization for mobile user equipments. Journal of Parallel and Distributed Computing, 150(??):69–84, April 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304226.

Dong:2023:FTU

[DLWF23] Hui Dong, Mengjie Lv, Huaqun Wang, and Weibei Fan. Fault-tolerant unicast using conditional local safe model in the data center network BCube. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001028.

DiFrancescoMaesa:2020:BAS

[DM20] Damiano Di Francesco Maesa and Paolo Mori. Blockchain 3.0 applications survey. Journal of Parallel and Distributed Computing, 138(??):99–114, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519308664.

Dabbaghjamanesh:2020:RTM

[DMKFJ20] Morteza Dabbaghjamanesh, Amirhossein Moeini, Abdollah Kavousi-Fard, and Alireza Jolfaei. Real-time monitoring and operation of microgrid using distributed cloud-fog architecture. Journal of Parallel and Distributed Computing, 146(??):15–24, December 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303130.

Dunning:2021:MPP

[DMM^{+}21] Daniel J. Dunning, Nathaniel R. Morgan, Jacob L. Moore, Eappen Nelluvelil, Tanya V. Tafolla, and Robert W. Robey. MATAR: a performance portability and productivity implementation of data-oriented design with Kokkos. Journal of Parallel and Distributed Computing, 157(??):86–104, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000770.

DePorre:2020:CDP

[DMSB20] Kevin De Porre, Florian Myter, Christophe Scholliers, and Elisa Gonzalez Boix. CScript: a distributed programming language for building mixed-consistency applications. Journal of Parallel and Distributed Computing, 144(??):109–123, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302896.

Datta:2022:QC

[DO22] Anwitaman Datta and Frédérique Oggier. Quorums over codes. Journal of Parallel and Distributed Computing, 161(??):1–19, March 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521002094.

Diaz:2024:VSE

[DPEL24] Antonio F. Díaz, Beatriz Prieto, Juan José Escobar, and Thomas Lampert. Vampire: a smart energy meter for synchronous monitoring in a distributed computer system. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001648.

Dutta:2021:DLI

[DPSD21] Nitul Dutta, Shobhit K. Patel, Vadim Samusenkov, and Vigneswaran D. Deep learning inspired routing in ICN using Monte Carlo Tree Search algorithm. Journal of Parallel and Distributed Computing, 150(??):104–111, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304305.

Ding:2021:FSR

[DQH^{+}21] Hao Ding, Junyan Qian, Bisheng Huang, Lingzhong Zhao, and Zhongyi Zhai. Flexible scheme for reconfiguring 2D meshconnected VLSI subarrays under row and column rerouting. Journal of Parallel and Distributed Computing, 151(??):1–12, May 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000034.

Ding:2021:HPV

[DQZZ21] Hao Ding, Junyan Qian, Lingzhong Zhao, and Zhongyi Zhai. A high-performance VLSI array reconfiguration scheme based on network flow under row and column rerouting. Journal of Parallel and Distributed Computing, 158(??):176–185, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001775.

doRosario:2021:ESM

[dRBB21] Vanderson Martins do Rosario, Mauricio Breternitz, and Edson Borin. Efficiency and scalability of multi-lane capsule networks (MLCN). Journal of Parallel and Distributed Computing, 155(??):63–73, September 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000939.

daSilva:2023:ROF

[dSSE23] Leylane Graziele Ferreira da Silva, Djamel F. H. Sadok, and Patricia Takako Endo. Resource optimizing federated learning for use with IoT: a systematic review. Journal of Parallel and Distributed Computing, 175(??):92–108, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000060.

Diamond:2021:PMB

[DSZ^{+}21] Gerrett Diamond, Cameron W. Smith, Chonglin Zhang, Eisung Yoon, and Mark S. Shephard. PUMIPic: a mesh-based approach to unstructured mesh Particle-In-Cell on GPUs. Journal of Parallel and Distributed Computing, 157(??):1–12, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001337.

David:2021:DFC

[DT21] Jisa David and Ciza Thomas. Discriminating flash crowds from DDoS attacks using efficient thresholding algorithm. Journal of Parallel and Distributed Computing, 152(??):79–87, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100040X.

Dehury:2024:HHR

[DVS24] Chinmaya Kumar Dehury, Bharadwaj Veeravalli, and Satish Narayana Srirama. HeRAFC: Heuristic resource allocation and optimization in MultiFog-Cloud environment. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001302.

Duan:2023:ADM

[DW23] Yubin Duan and Jie Wu. Accelerating distributed machine learning with model compression and graph partition. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000680.

Dorier:2023:TES

[DWR^{+}23] Matthieu Dorier, Zhe Wang, Srinivasan Ramesh, Utkarsh Ayachit, Shane Snyder, Rob Ross, and Manish Parashar. Towards elastic in situ analysis for high-performance computing simulations. Journal of Parallel and Distributed Computing, 177(??): 106–116, July 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000369.

Dai:2021:TBS

[DWW^{+}21] Weiqi Dai, Qinyuan Wang, Zeli Wang, Xiaobin Lin, Deqing Zou, and Hai Jin. Trustzone-based secure lightweight wallet for hyperledger fabric. Journal of Parallel and Distributed Computing, 149(??):66–75, March 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304020.

Dai:2022:TSE

[DWWX22] Hao Dai, Jiashu Wu, Yang Wang, and Chengzhong Xu. Towards scalable and efficient Deep-RL in edge computing: a game-based partition approach. Journal of Parallel and Distributed Computing, 168(??):108–119, October 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001411.

Ding:2024:PIM

[DZ24] Yi Ding and Linhe Zhu. Parameter identification method of a reaction-diffusion network information propagation system based on optimization theory. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000522.

Dou:2023:TCE

[DZZ^{+}23] Hui Dou, Lei Zhang, Yiwen Zhang, Pengfei Chen, and Zibin Zheng. TurBO: a cost-efficient configuration-based auto-tuning approach for cluster-based big data frameworks. Journal of Parallel and Distributed Computing, 177(??):89–105, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000382.

ElJoubari:2022:SMM

[EBV22] Oumaima El Joubari, Jalel Ben Othman, and Veronique Veque. A stochastic mobility model for traffic forecasting in urban environments. Journal of Parallel and Distributed Computing, 165(??):142–155, July 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000636.

Ernst:2023:APE

[EHH^{+}23] Dominik Ernst, Markus Holzer, Georg Hager, Matthias Knorr, and Gerhard Wellein. Analytical performance estimation during code generation on modern GPUs. Journal of Parallel and Distributed Computing, 173(??):152–167, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002313.

Emara:2020:SBM

[EL20] Moustafa Emara and Bo-Cheng Lai. Selective bypassing and mapping for heterogeneous applications on GPGPUs. Journal of Parallel and Distributed Computing, 142(??):106–118, August 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518304714.

Entezari-Maleki:2020:EMP

[EMCE20] Reza Entezari-Maleki, Younghyun Cho, and Bernhard Egger. Evaluation of memory performance in NUMA architectures using stochastic reward nets. Journal of Parallel and Distributed Computing, 144(??):172–188, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303075.

El-Moursy:2020:PPB

[EMSEMM20] Ali A. El-Moursy, Fadi N. Sibai, Magdy A. El-Moursy, and Ahmed S. S. Mohamed. PMSMC: Priority-based multirequestor scheduler for embedded system memory controller. Journal of Parallel and Distributed Computing, 139(??):135– 147, May 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151930139X.

Eizaguirre:2024:SKB

[ESA24] Germán T. Eizaguirre and Marc Sánchez-Artigas. A seer knows best: Auto-tuned object storage shuffling for serverless analytics. Journal of Parallel and Distributed Computing, 183(??): ??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001338.

Fan:2021:PPP

[FBL^{+}21] Yongkai Fan, Jianrong Bai, Xia Lei, Weiguo Lin, Qian Hu, Guodong Wu, Jiaming Guo, and Gang Tan. PPMCK: Privacypreserving multi-party computing for K-means clustering. Journal of Parallel and Distributed Computing, 154(??):54–63, August 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000629.

Ferraz:2024:DEW

[FDT^{+}24] Samuel Ferraz, Vinicius Dias, Carlos H. C. Teixeira, Srinivasan Parthasarathy, George Teodoro, and Wagner Meira. DuMato: an efficient warp-centric subgraph enumeration system for GPU. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731524000674.

Feuilloley:2022:ESP

[FF22] Laurent Feuilloley and Pierre Fraigniaud. Error-sensitive prooflabeling schemes. Journal of Parallel and Distributed Computing, 166(??):149–165, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522000934.

Fernandez-Fabeiro:2021:DPH

[FFGEL21] Jorge Fernández-Fabeiro, Arturo Gonzalez-Escribano, and Diego R. Llanos. Distributed programming of a hyperspectral image registration algorithm for heterogeneous GPU clusters. Journal of Parallel and Distributed Computing, 151(??): 86–93, May 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000356.

Friha:2022:FFL

[FFS^{+}22] Othmane Friha, Mohamed Amine Ferrag, Lei Shu, Leandros Maglaras, Kim-Kwang Raymond Choo, and Mehdi Nafaa. FELIDS: Federated learning-based intrusion detection system for agricultural Internet of Things. Journal of Parallel and Distributed Computing, 165(??):17–31, July 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000570.

Fan:2020:PPC

[FHG^{+}20] Weibei Fan, Jing He, Mengjiao Guo, Peng Li, Zhijie Han, and Ruchuan Wang. Privacy preserving classification on local differential privacy in data centers. Journal of Parallel and Distributed Computing, 135(??):70–82, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303296.

Filipovic:2022:UHP

[FHN^{+}22] Jiří Filipovič, Jana Hozzová, Amin Nezarat, Jaroslav Ol’ha, and Filip Petrovič. Using hardware performance counters to speed up autotuning convergence on GPUs. Journal of Parallel and Distributed Computing, 160(??):16–35, February 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001945.

Flores-Lamas:2020:DAM

[FLFZTS20] Alejandro Flores-Lamas, José Alberto Fernández-Zepeda, and Joel Antonio Trejo-Sánchez. A distributed algorithm for a maximal 2-packing set in Halin graphs. Journal of Parallel and Distributed Computing, 142(??):62–76, August 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304459.

Fan:2020:SEO

[FLZ^{+}20] Kai Fan, Tingting Liu, Kuan Zhang, Hui Li, and Yintang Yang. A secure and efficient outsourced computation on data sharing scheme for privacy computing. Journal of Parallel and Distributed Computing, 135(??):169–176, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303351.

Freitas:2021:PSD

[FPdLS^{+}21] Vinicius Freitas, Laércio L. Pilla, Alexandre de L. Santana, Márcio Castro, and Johanne Cohen. PackStealLB: a scalable distributed load balancer based on work stealing and workload discretization. Journal of Parallel and Distributed Computing, 150(??):34–45, April 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304214.

Finol:2024:EIE

[FPGLSA24] Gerard Finol, Gerard París, Pedro García-López, and Marc Sánchez-Artigas. Exploiting inherent elasticity of serverless in algorithms with unbalanced and irregular workloads. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000558.

Feng:2023:TPE

[FQL^{+}23] Jiahui Feng, Jingze Qi, Yuanning Liu, Liyan Dong, and Zhen Liu. A task processing efficiency improvement scheme based on cloud-edge architecture in computationally intensive scenarios. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001120.

Feliu:2023:SIT

[FRAK23] Josué Feliu, Alberto Ros, Manuel E. Acacio, and Stefanos Kaxiras. Speculative inter-thread store-to-load forwarding in SMT architectures. Journal of Parallel and Distributed Computing, 173(??):94–106, March 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002349.

Freytag:2021:CEF

[FSL^{+}21] Gabriel Freytag, Matheus S. Serpa, João V. F. Lima, Paolo Rech, and Philippe O. A. Navaux. Collaborative execution of fluid flow simulation using non-uniform decomposition on heterogeneous architectures. Journal of Parallel and Distributed Computing, 152(??):11–20, June 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000277.

Farooq:2022:MLI

[FTA^{+}22] Umer Farooq, Noshina Tariq, Muhammad Asim, Thar Baker, and Ahmed Al-Shamma’a. Machine learning and the Internet of Things security: Solutions and open challenges. Journal of Parallel and Distributed Computing, 162(??):89–104, April 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000235.

Frei:2023:EDM

[FW23] Fabian Frei and Koichi Wada. Efficient deterministic MapReduce algorithms for parallelizable problems. Journal of Parallel and Distributed Computing, 177(??):28–38, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152300031X.

Fan:2020:DPD

[FWZ^{+}20] Yuqi Fan, Chen Wang, Bei Zhang, Shuyang Gu, Weili Wu, and Dingzhu Du. Data placement in distributed data centers for improved SLA and network cost. Journal of Parallel and Distributed Computing, 146(??):189–200, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152030335X.

Golshani:2021:PAS

[GA21] Ehsan Golshani and Mehrdad Ashtiani. Proactive autoscaling for cloud environments using temporal convolutional neural networks. Journal of Parallel and Distributed Computing, 154(??):119–141, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000836.

Gou:2022:MSP

[GBC^{+}22] Changjiang Gou, Anne Benoit, Mingsong Chen, Loris Marchal, and Tongquan Wei. Mapping series-parallel streaming applications on hierarchical platforms with reliability and energy constraints. Journal of Parallel and Distributed Computing, 163(??):45–61, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522000211.

Garralda-Barrio:2024:NFG

[GBEFBC24] Mariano Garralda-Barrio, Carlos Eiras-Franco, and Verónica Bolón-Canedo. A novel framework for generic Spark workload characterization and similar pattern recognition using machine learning. Journal of Parallel and Distributed Computing, 189 (??):??, July 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731524000455.

Gotfryd:2023:DDA

[GC23] Karol Gotfryd and Jacek Cichoń. On distributed data aggregation and the precision of approximate histograms. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000928.

Goscinski:2023:SID

[GDF^{+}23] Andrzej Goscinski, Flavia C. Delicato, Giancarlo Fortino, Anna Kobusińska, and Gautam Srivastava. Special issue on Distributed Intelligence at the Edge for the Future Internet of Things. Journal of Parallel and Distributed Computing, 171(??):157–162, January 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200209X.

Garcia-Font:2020:SAD

[GF20] Victor Garcia-Font. SocialBlock: an architecture for decentralized user-centric data management applications for communications in smart cities. Journal of Parallel and Distributed Computing, 145(??):13–23, November 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303117.

Gutierrez-Garcia:2023:ACL

[GGTSFD23] J. Octavio Gutierrez-Garcia, Joel Antonio Trejo-Sánchez, and Daniel Fajardo-Delgado. Agent coalitions for load balancing in cloud data centers. Journal of Parallel and Distributed Computing, 172(??):1–17, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002143.

Ghosal:2020:DDC

[GHD20] Amrita Ghosal, Subir Halder, and Sajal K. Das. Distributed on-demand clustering algorithm for lifetime optimization in wireless sensor networks. Journal of Parallel and Distributed Computing, 141(??):129–142, July 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305793.

Gebhard:2023:ITA

[GHKKL23] Oliver Gebhard, Max Hahn-Klimroth, Dominik Kaaser, and Philipp Loick. Information-theoretic and algorithmic aspects of parallel and distributed reconstruction from pooled data. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000886.

Georgiou:2022:ITE

[GHNS22] Chryssis Georgiou, Theophanis Hadjistasi, Nicolas Nicolaou, and Alexander A. Schwarzmann. Implementing three exchange read operations for distributed atomic storage. Journal of Parallel and Distributed Computing, 163(??):97–113, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000314.

Gschwandtner:2021:CCT

[GHT^{+}21] Philipp Gschwandtner, Alexander Hirsch, Peter Thoman, Peter Zangerl, Herbert Jordan, and Thomas Fahringer. The cluster coffer: Teaching HPC on the road. Journal of Parallel and Distributed Computing, 155(??):50–62, September 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000964.

Gao:2024:RTC

[GJL^{+}24] Jianhua Gao, Weixing Ji, Jie Liu, Yizhuo Wang, and Feng Shi. Revisiting thread configuration of SpMV kernels on GPU: a machine learning based approach. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001697.

Gergel:2021:PSM

[GK21] Victor Gergel and Evgeniy Kozinov. Parallel solving of multiple information-coordinated global optimization problems. Journal of Parallel and Distributed Computing, 154(??):153–162, August 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000927.

Gupta:2024:ELL

[GK24] Arya Tanmay Gupta and Sandeep S. Kulkarni. Eventually lattice-linear algorithms. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523001727.

Garg:2020:AEA

[GKB^{+}20] Sahil Garg, Kuljeet Kaur, Shalini Batra, Gagangeet Singh Aujla, Graham Morgan, Neeraj Kumar, Albert Y. Zomaya, and Rajiv Ranjan. En-ABC: an ensemble artificial bee colony based anomaly detection scheme for cloud environment. Journal of Parallel and Distributed Computing, 135(??):219–233, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304721.

Guerraoui:2021:PTF

[GKP21] Rachid Guerraoui, David Kozhaya, and Yvonne-Anne Pignolet. Probabilistic and temporal failure detectors for solving distributed problems. Journal of Parallel and Distributed Computing, 158(??):1–15, December 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001635.

Gkikopoulos:2024:RIA

[GKSS24] Panagiotis Gkikopoulos, Peter Kropf, Valerio Schiavoni, and Josef Spillner. Reliable IoT analytics at scale. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000042.

Gast:2021:AWS

[GKTW21] Nicolas Gast, Mohammed Khatiri, Denis Trystram, and Frédéric Wagner. Analysis of work stealing with latency. Journal of Parallel and Distributed Computing, 153(??):119– 129, July 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000630.

Gao:2022:FBB

[GLC^{+}22] Liang Gao, Li Li, Yingwen Chen, ChengZhong Xu, and Ming Xu. FGFL: a blockchain-based fair incentive governor for federated learning. Journal of Parallel and Distributed Computing, 163(??):283–299, May 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000259.

Ge:2020:BBD

[GLF20] Chunpeng Ge, Zhe Liu, and Liming Fang. A blockchain based decentralized data security mechanism for the Internet of Things. Journal of Parallel and Distributed Computing, 141 (??):1–9, July 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151930810X.

Guo:2021:PPA

[GLL21] Luanzheng Guo, Dong Li, and Ignacio Laguna. PARIS: Predicting application resilience using machine learning. Journal of Parallel and Distributed Computing, 152(??):111–124, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000368.

Guan:2021:AEP

[GLY^{+}21] Zhitao Guan, Xin Lu, Wenti Yang, Longfei Wu, Naiyu Wang, and Zijian Zhang. Achieving efficient and privacy-preserving energy trading based on blockchain and ABE in smart grid. Journal of Parallel and Distributed Computing, 147(??):34–45, January 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303609.

Gonzalez:2021:MGS

[GM21] Marc González and Enric Morancho. Multi-GPU systems and Unified Virtual Memory for scientific applications: the case of the NAS multi-zone parallel benchmarks. Journal of Parallel and Distributed Computing, 158(??):138–150, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001672.

Gul:2022:IBI

[GMA^{+}22] Faiza Gul, Imran Mir, Deemah Alarabiat, Hamzeh Mohammad Alabool, Laith Abualigah, and Suleman Mir. Implementation of bio-inspired hybrid algorithm with mutation operator for robotic path planning. Journal of Parallel and Distributed Computing, 169(??):171–184, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001496.

Guo:2024:RNE

[GMLW24] Chaoming Guo, Meijie Ma, Xiang-Jun Li, and Guijuan Wang. The Rabin numbers of enhanced hypercubes. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000698.

Gorain:2024:CDS

[GMMP24] Barun Gorain, Partha Sarathi Mandal, Kaushik Mondal, and Supantha Pandit. Collaborative dispersion by silent robots. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000169.

Goncharow:2021:CMS

[GMS^{+}21] Alec Goncharow, Matthew Mcquaigue, Erik Saule, Kalpathi Subramanian, Paula Goolkasian, and Jamie Payton. CS-Materials: a system for classifying and analyzing pedagogical materials to improve adoption of parallel and distributed computing topics in early CS courses. Journal of Parallel and Distributed Computing, 157(??):316–330, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100126X.

Gowanlock:2021:HKJ

[Gow21] Michael Gowanlock. Hybrid KNN-join: Parallel nearest neighbor searches exploiting CPU and GPU architectural features. Journal of Parallel and Distributed Computing, 149(??):119– 137, March 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520304056.

Gu:2023:SAH

[GPC23] Mei-Mei Gu, Kung-Jui Pai, and Jou-Ming Chang. Subversion analyses of hierarchical networks based on (edge) neighbor connectivity. Journal of Parallel and Distributed Computing, 171(??):54–65, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001988.

Gonzalez:2022:OCM

[GPH^{+}22] Jorge Gonzalez, Mauricio G. Palma, Maarten Hattink, Ruth Rubio-Noriega, Lois Orosa, Onur Mutlu, Keren Bergman, and Rodolfo Azevedo. Optically connected memory for disaggregated data centers. Journal of Parallel and Distributed Computing, 163(??):300–312, May 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200020X.

Gudeme:2021:CPP

[GPK21] Jaya Rao Gudeme, Syamkumar Pasupuleti, and Ramesh Kandukuri. Certificateless privacy preserving public auditing for dynamic shared data with group user revocation in cloud storage.

Journal of Parallel and Distributed Computing, 156(??):163– 175, October 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521001301.

Gu:2021:SEG

[GQW^{+}21] Rong Gu, Yang Qi, Tongyu Wu, Zhaokang Wang, Xiaolong Xu, Chunfeng Yuan, and Yihua Huang. SparkDQ: Efficient generic big data quality management on distributed data-parallel computation. Journal of Parallel and Distributed Computing, 156(??):132–147, October 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001246.

Guo:2020:SFT

[GQX20] Litao Guo, Chengfu Qin, and Liqiong Xu. Subgraph fault tolerance of distance optimally edge connected hypercubes and folded hypercubes. Journal of Parallel and Distributed Computing, 138(??):190–198, April 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302862.

Gao:2022:RWV

[GRZT22] Shuzhi Gao, Xuepeng Ren, Yimin Zhang, and Haihong Tang. Reliability of the weight vector generation method of the multiobjective evolutionary algorithm and application. Journal of Parallel and Distributed Computing, 169(??):130–156, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001514.

Geiser:2020:SMB

[GS20] Georg Geiser and Wolfgang Schröder. Structured multiblock grid partitioning using balanced cut trees. Journal of Parallel and Distributed Computing, 138(??):139–152, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518303502.

Giacaman:2021:VAV

[GSA21] Nasser Giacaman, Oliver Sinnen, and Joel Adams. Visual analogy videos for understanding fundamental parallel scheduling policies. Journal of Parallel and Distributed Computing, 153(??):64–74, July 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521000733.

Guo:2022:ECF

[GSM22] Huimei Guo, Eminjan Sabir, and Aygul Mamut. The gextra connectivity of folded crossed cubes. Journal of Parallel and Distributed Computing, 166(??):139–146, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000922.

Gonggiatgul:2023:PBB

[GSMÖ23] Taspon Gonggiatgul, Ghassan Shobaki, and Pınar Muyan-Özçelik. A parallel branch-and-bound algorithm with historybased domination and its application to the sequential ordering problem. Journal of Parallel and Distributed Computing, 172(??):131–143, February 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522002167.

Gupta:2021:DFT

[GSV21] Pushpanjali Gupta, Prasan Kumar Sahoo, and Bharadwaj Veeravalli. Dynamic fault tolerant scheduling with response time minimization for multiple failures in cloud. Journal of Parallel and Distributed Computing, 158(??):80–93, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001659.

Girolami:2022:MBD

[GVC^{+}22] Michele Girolami, Piergiorgio Vitello, Andrea Capponi, Claudio Fiandrino, Luca Foschini, and Paolo Bellavista. A mobilitybased deployment strategy for edge data centers. Journal of Parallel and Distributed Computing, 164(??):133–141, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000600.

Greguric:2024:IFD

[GVI24] Martin Gregurić, Filip Vrbanić, and Edouard Ivanjko. Impact of federated deep learning on vehicle-based speed control in mixed traffic flows. Journal of Parallel and Distributed Computing, 186 (??):??, April 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373152300182X.

Ganguly:2024:DRV

[GXJ^{+}24] Ritam Ganguly, Yingjie Xue, Aaron Jonckheere, Parker Ljung, Benjamin Schornstein, Borzoo Bonakdarpour, and Maurice Herlihy. Distributed runtime verification of metric temporal properties. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523001715.

Gao:2021:ADS

[GXYH21] Jiaquan Gao, Yifei Xia, Renjie Yin, and Guixia He. Adaptive diagonal sparse matrix-vector multiplication on GPU. Journal of Parallel and Distributed Computing, 157(??):287–302, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001532.

Harikrishna:2021:RMP

[HA21] Pillutla Harikrishna and A. Amuthan. Rival-Model Penalized Self-Organizing Map enforced DDoS attack prevention mechanism for software defined network-based cloud computing environment. Journal of Parallel and Distributed Computing, 154(??):142–152, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000587.

Hashmi:2020:FXZ

[HCC^{+}20] Jahanzeb Maqbool Hashmi, Ching-Hsiang Chu, Sourav Chakraborty, Mohammadreza Bayatpour, Hari Subramoni, and Dhabaleswar K. Panda. FALCON-X: Zero-copy MPI derived datatype processing on modern CPU and GPU architectures. Journal of Parallel and Distributed Computing, 144(??):1–13, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302872.

Huang:2021:TSS

[HCY^{+}21] Chenlin Huang, Wei Chen, Lu Yuan, Yan Ding, Songlei Jian, Yusong Tan, Hua Chen, and Dan Chen. Toward security as a service: a trusted cloud service architecture with policy customization. Journal of Parallel and Distributed Computing, 149(??):76–88, March 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304032.

Hegedus:2021:DLW

[HDJ21] István Hegedüs, Gábor Danner, and Márk Jelasity. Decentralized learning works: an empirical comparison of gossip learning and federated learning. Journal of Parallel and Distributed Computing, 148(??):109–124, February 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303890.

Hossain:2020:KDM

[HFA20] M. Anwar Hossain, Rahatara Ferdousi, and Mohammed F. Alhamid. Knowledge-driven machine learning based framework for early-stage disease risk prediction in edge environment. Journal of Parallel and Distributed Computing, 146(??):25–34, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303324.

He:2022:BBA

[HFP^{+}22] Songlin He, Eric Ficke, Mir Mehedi Ahsan Pritom, Huashan Chen, Qiang Tang, Qian Chen, Marcus Pendleton, Laurent Njilla, and Shouhuai Xu. Blockchain-based automated and robust cyber security management. Journal of Parallel and Distributed Computing, 163(??):62–82, May 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000089.

Haris:2023:STT

[HGC^{+}23] Jude Haris, Perry Gibson, José Cano, Nicolas Bohm Agostini, and David Kaeli. SECDA-TFLite: a toolkit for efficient development of FPGA-based DNN accelerators for edge inference. Journal of Parallel and Distributed Computing, 173(??):140– 151, March 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522002301.

Hosseinioun:2020:NEA

[HKTG20] Pejman Hosseinioun, Maryam Kheirabadi, Seyed Reza Kamel Tabbakh, and Reza Ghaemi. A new energy-aware tasks scheduling approach in fog computing using hybrid meta-heuristic algorithm. Journal of Parallel and Distributed Computing, 143(??):88–96, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373152030023X.

Hsieh:2020:UPA

[HLBZ20] Sun-Yuan Hsieh, Cheng-Sheng Liu, Rajkumar Buyya, and Albert Y. Zomaya. Utilization-prediction-aware virtual machine consolidation approach for energy-efficient cloud data centers. Journal of Parallel and Distributed Computing, 139(??):99– 109, May 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151930190X.

Hiessl:2022:CBF

[HLK^{+}22] Thomas Hiessl, Safoura Rezapour Lakani, Jana Kemnitz, Daniel Schall, and Stefan Schulte. Cohort-based federated learning services for industrial collaboration on the edge. Journal of Parallel and Distributed Computing, 167(??):64–76, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000995.

Hu:2021:CFD

[HLL^{+}21] Junyan Hu, Kenli Li, Chubo Liu, Jianguo Chen, and Keqin Li. Coalition formation for deadline-constrained resource procurement in cloud computing. Journal of Parallel and Distributed Computing, 149(??):1–12, March 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303877.

Hou:2022:NFV

[HLL^{+}22] Jian Hou, Fangai Liu, Hui Lu, Zhiyuan Tan, Xuqiang Zhuang, and Zhihong Tian. A novel flow-vector generation approach for malicious traffic detection. Journal of Parallel and Distributed Computing, 169(??):72–86, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001332.

Hosseininoorbin:2023:EET

[HLS^{+}23] Seyedehfaezeh Hosseininoorbin, Siamak Layeghy, Mohanad Sarhan, Raja Jurdak, and Marius Portmann. Exploring edge TPU for network intrusion detection in IoT. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000758.

Huang:2022:RLC

[HLX^{+}22] Binbin Huang, Xiao Liu, Yuanyuan Xiang, Dongjin Yu, Shuiguang Deng, and Shangguang Wang. Reinforcement learning for cost-effective IoT service caching at the edge. Journal of Parallel and Distributed Computing, 168(??):120–136, October 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001423.

Huang:2022:AMC

[HLZ22] Tian Huang, Tao Luo, and Joey Tianyi Zhou. APT: the master-copy-free training method for quantised neural network on edge devices. Journal of Parallel and Distributed Computing, 166(??):95–103, August 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200082X.

Huo:2020:DEP

[HMCG20] Zenan Huo, Gang Mei, Giampaolo Casolla, and Fabio Giampaolo. Designing an efficient parallel spectral clustering algorithm on multi-core processors in Julia. Journal of Parallel and Distributed Computing, 138(??):211–221, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519308743.

Halver:2020:KIE

[HMS20] Rene Halver, Jan H. Meinke, and Godehard Sutmann. Kokkos implementation of an Ewald Coulomb solver and analysis of performance portability. Journal of Parallel and Distributed Computing, 138(??):48–54, April 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519305040.

Hassanzadeh-Nazarabadi:2021:IFD

[HNKÖ21] Yahya Hassanzadeh-Nazarabadi, Alptekin Küpçü, and Öznur Özkasap. Interlaced: Fully decentralized churn stabilization for skip graph-based DHTs. Journal of Parallel and Distributed Computing, 149(??):13–28, March 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303919.

He:2020:PPE

[HNN^{+}20] Yuanyuan He, Jianbing Ni, Ben Niu, Fenghua Li, and Xuemin (Sherman) Shen. Privbus: a privacy-enhanced crowdsourced bus service via fog computing. Journal of Parallel and Distributed Computing, 135(??):156–168, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303284.

Huang:2022:ICO

[HQL^{+}22] Dan Huang, Zhenlu Qin, Qing Liu, Norbert Podhorszki, and Scott Klasky. Identifying challenges and opportunities of in-memory computing on large HPC systems. Journal of Parallel and Distributed Computing, 164(??):106–122, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000387.

Hezavehi:2023:IAB

[HR23] Sasha Mahdavi Hezavehi and Rouhollah Rahmani. Interactive anomaly-based DDoS attack detection method in cloud computing environments using a third party auditor. Journal of Parallel and Distributed Computing, 178(??):82–99, August 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000588.

He:2022:ODG

[HSHT22] Huaiwen He, Hong Shen, Qing Hao, and Hui Tian. Online delay-guaranteed workload scheduling to minimize power cost in cloud data centers using renewable energy. Journal of Parallel and Distributed Computing, 159(??):51–64, January 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001805.

Huang:2021:BBE

[HSX^{+}21] Haiping Huang, Xiang Sun, Fu Xiao, Peng Zhu, and Wenming Wang. Blockchain-based eHealth system for auditable EHRs manipulation in cloud environments. Journal of Parallel and Distributed Computing, 148(??):46–57, February 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303853.

Hao:2023:EAT

[HWM^{+}23] Yongsheng Hao, Qi Wang, Tinghuai Ma, Jinglin Du, and Jie Cao. Energy allocation and task scheduling in edge devices based on forecast solar energy with meteorological information. Journal of Parallel and Distributed Computing, 177(??):171– 181, July 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000412.

Hussain:2024:CAQ

[HWR^{+}24] Mehboob Hussain, Lian-Fu Wei, Amir Rehman, Muqadar Ali, Syed Muhammad Waqas, and Fakhar Abbas. Cost-aware quantum-inspired genetic algorithm for workflow scheduling in hybrid clouds. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731524000844.

Huang:2020:BRE

[HZL^{+}20] Bobo Huang, Rui Zhang, Zhihui Lu, Yiming Zhang, Jie Wu, Lu Zhan, and Patrick C. K. Hung. BPS: a reliable and efficient pub/sub communication model with blockchainenhanced paradigm in multi-tenant edge cloud. Journal of Parallel and Distributed Computing, 143(??):167–178, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302847.

Hu:2021:SSA

[HZY^{+}21] Tao Hu, Zhen Zhang, Peng Yi, Dong Liang, Ziyong Li, Quan Ren, Yuxiang Hu, and Julong Lan. SEAPP: a secure application management framework based on REST API access control in SDN-enabled cloud environment. Journal of Parallel and Distributed Computing, 147(??):108–123, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303671.

Isupov:2020:DIM

[IKK20] Konstantin Isupov, Vladimir Knyazkov, and Alexander Kuvaev. Design and implementation of multiple-precision BLAS Level 1 functions for graphics processing units. Journal of Parallel and Distributed Computing, 140(??):25–36, June 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303302.

Ignatov:2023:TPH

[IMP^{+}23] A. Ignatov, I. Maslova, M. Posypkin, W. Yang, and J. Wu. A two-phase heuristic algorithm for power-aware offline scheduling in IaaS clouds. Journal of Parallel and Distributed Computing, 178(??):1–10, August 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523000424.

Ibeid:2020:FFM

[IOG20] Huda Ibeid, Luke Olson, and William Gropp. FFT, FMM, and multigrid on the road to exascale: Performance challenges and opportunities. Journal of Parallel and Distributed Computing, 136(??):63–74, February 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305513.

Ibrahim:2020:EES

[IRA20] Godar J. Ibrahim, Tarik A. Rashid, and Mobayode O. Akinsolu. An energy efficient service composition mechanism using a hybrid meta-heuristic algorithm in a mobile cloud environment. Journal of Parallel and Distributed Computing, 143(??):77–87, September 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520302744.

Islam:2023:OPA

[IRLN23] Mohammad Mainul Islam, Fahimeh Ramezani, Hai Yan Lu, and Mohsen Naderpour. Optimal placement of applications in the fog environment: a systematic literature review. Journal of Parallel and Distributed Computing, 174(??):46–69, April 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002465.

Ivoghlian:2022:AAA

[IWS22] Ameer Ivoghlian, Kevin I-Kai Wang, and Zoran Salcic. Application-aware adaptive parameter control for LoRaWAN. Journal of Parallel and Distributed Computing, 166(??):166– 177, August 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522001010.

Jararweh:2020:EES

[Jar20] Yaser Jararweh. Enabling efficient and secure energy cloud using edge computing and 5G. Journal of Parallel and Distributed Computing, 145(??):42–49, November 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152030321X.

Jamieson:2020:HLP

[JB20] Maurice Jamieson and Nick Brown. High level programming abstractions for leveraging hierarchical memories with micro-core architectures. Journal of Parallel and Distributed Computing, 138(??):128–138, April 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305616.

Jeannot:2022:PMT

[Jea22] Emmanuel Jeannot. Process mapping on any topology with TopoMatch. Journal of Parallel and Distributed Computing, 170(??):39–52, December 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001794.

Jemmali:2021:BSP

[JH21] Mahdi Jemmali and Lotfi Hidri. Bounding schemes for the parallel machine scheduling problem with DeJong’s learning effect. Journal of Parallel and Distributed Computing, 156(??):101– 118, October 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521001088.

Javed:2021:PNC

[JHML21] Aqib Javed, Jim Harkin, Liam McDaid, and Junxiu Liu. Predicting networks-on-chip traffic congestion with spiking neural networks. Journal of Parallel and Distributed Computing, 154(??):82–93, August 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521000666.

Jiang:2020:AHD

[JHZ20] Tianming Jiang, Ping Huang, and Ke Zhou. Achieving high data reliability at low scrubbing cost via failure-aware scrubbing. Journal of Parallel and Distributed Computing, 144(??):220–229, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520302860.

Jayabalan:2022:SBM

[JJ22] Jayapriya Jayabalan and N. Jeyanthi. Scalable blockchain model using off-chain IPFS storage for healthcare data security and privacy. Journal of Parallel and Distributed Computing, 164(??):152–167, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522000648.

Javaid:2021:ASH

[JJJ21] Nadeem Javaid, Naeem Jan, and Muhammad Umar Javed. An adaptive synthesis to handle imbalanced big data with deep siamese network for electricity theft detection in smart grids. Journal of Parallel and Distributed Computing, 153(??):44– 52, July 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000551.

Jordan:2023:EAF

[JKK^{+}23] Michael Guilherme Jordan, Guilherme Korol, Tiago Knorst, Mateus Beck Rutzig, and Antonio Carlos Schneider Beck. Energy-aware fully-adaptive resource provisioning in collaborative CPU–FPGA cloud environments. Journal of Parallel and Distributed Computing, 176(??):55–69, June 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000308.

Junger:2022:GPG

[JKM^{+}22] Daniel Jünger, Robin Kobus, André Müller, Christian Hundt, Kai Xu, Weiguo Liu, and Bertil Schmidt. General-purpose GPU hashing data structures and their application in accelerated genomics. Journal of Parallel and Distributed Computing, 163(??): 256–268, May 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000132.

Jia:2023:EGB

[JL23] Zhuohao Jia and Simon Liao. An efficient GPU-based method to compute high-order Zernike moments. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000990.

Johnson:2022:FCP

[JP22] Martin Johnson and Daniel Playne. A fast and concise parallel implementation of the 8 × 8 2D forward and inverse DCTs using Halide. Journal of Parallel and Distributed Computing, 163(??):20–29, May 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000223.

Jafari:2021:FSM

[JSA21] Nazanin Jafari, Oguz Selvitopi, and Cevdet Aykanat. Fast shared-memory streaming multilevel graph partitioning. Journal of Parallel and Distributed Computing, 147(??):140–151, January 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303658.

Jebali:2022:SDO

[JSJC22] Adel Jebali, Salma Sassi, Abderrazak Jemai, and Richard Chbeir. Secure data outsourcing in presence of the inference problem: a graph-based approach. Journal of Parallel and Distributed Computing, 160(??):1–15, February 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001842.

Jian:2022:AAL

[JTV^{+}22] Yubing Jian, Ching-Lun Tai, Shyam Krishnan Venkateswaran, Mohit Agarwal, Yuchen Liu, Douglas M. Blough, and Raghupathy Sivakumar. Algorithms for addressing line-of-sight issues in mmWave WiFi networks using access point mobility. Journal of Parallel and Distributed Computing, 160(??):65–78, February 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002069.

Jahja:2022:PRD

[JYH22] Irvan Jahja, Haifeng Yu, and Ruomu Hou. On the power of randomization in distributed algorithms in dynamic networks with adaptive adversaries. Journal of Parallel and Distributed Computing, 159(??):35–50, January 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001829.

Jin:2020:TAD

[JZS^{+}20] Tong Jin, Fan Zhang, Qian Sun, Melissa Romanus, Hoang Bui, and Manish Parashar. Towards autonomic data management for staging-based coupled scientific workflows. Journal of Parallel and Distributed Computing, 146(??):35–51, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303312.

Jiang:2020:EAP

[JZWX20] Yan Jiang, Youwen Zhu, Jian Wang, and Yong Xiang. Efficient authentication protocol with anonymity and key protection for mobile Internet users. Journal of Parallel and Distributed Computing, 137(??):179–191, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303107.

Kudva:2021:SBB

[KBS^{+}21] Sowmya Kudva, Shahriar Badsha, Shamik Sengupta, Hung La, Ibrahim Khalil, and Mohammed Atiquzzaman. A scalable blockchain based trust management in VANET routing protocol. Journal of Parallel and Distributed Computing, 152(??):144–156, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521000459.

Kandekar:2020:DAC

[KBW20] S. A. Kandekar, Y. M. Borse, and B. N. Waphare. Decomposition of augmented cubes into regular connected pancyclic subgraphs. Journal of Parallel and Distributed Computing, 141(??): 74–81, July 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518306579.

Khalily-Dermany:2021:DAC

[KD21] Moammad Khalily-Dermany. A decentralized algorithm to combine topology control with network coding. Journal of Parallel and Distributed Computing, 149(??):174–185, March 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304172.

Kelefouras:2022:WSM

[KD22] Vasilios Kelefouras and Karim Djemame. Workflow simulation and multi-threading aware task scheduling for heterogeneous computing. Journal of Parallel and Distributed Computing, 168(??):17–32, October 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001265.

Karahoda:2020:MMP

[KEK^{+}20] Sertaç Karahoda, Osman Tufan Erenay, Kamer Kaya, Uraz Cengiz Türker, and Hüsnü Yenigün. Multicore and manycore parallelization of cheap synchronizing sequence heuristics. Journal of Parallel and Distributed Computing, 140(??):13–24, June 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303776.

Kahveci:2020:JES

[KG20] Basri Kahveci and Bugra Gedik. Joker: Elastic stream processing with organic adaptation. Journal of Parallel and Distributed Computing, 137(??):205–223, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519308184.

Kumar:2021:SIE

[KGP^{+}21] Mohit Kumar, Saurabh Gupta, Tirthak Patel, Michael Wilder, Weisong Shi, Song Fu, Christian Engelmann, and Devesh Tiwari. Study of interconnect errors, network congestion, and applications characteristics for throttle prediction on a large scale HPC system. Journal of Parallel and Distributed Computing, 153(??):29–43, July 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100054X.

Keramatian:2021:MCM

[KGPT21] Amir Keramatian, Vincenzo Gulisano, Marina Papatriantafilou, and Philippas Tsigas. MAD-C: Multi-stage approximate distributed cluster-combining for obstacle detection and localization. Journal of Parallel and Distributed Computing, 147(??):248–267, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520303610.

Keramatian:2023:PCF

[KGPT23] Amir Keramatian, Vincenzo Gulisano, Marina Papatriantafilou, and Philippas Tsigas. PARMA-CC: a family of parallel multiphase approximate cluster combining algorithms. Journal of Parallel and Distributed Computing, 177(??):68–88, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000163.

Kumari:2020:BAA

[KGTK20] Aparna Kumari, Rajesh Gupta, Sudeep Tanwar, and Neeraj Kumar. Blockchain and AI amalgamation for energy cloud management: Challenges, solutions, and future directions. Journal of Parallel and Distributed Computing, 143(??):148–166, September 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373152030277X.

Khemili:2022:EAF

[KHO22] Wided Khemili, Jalel Eddine Hajlaoui, and Mohamed Nazih Omri. Energy aware fuzzy approach for placement and consolidation in cloud data centers. Journal of Parallel and Distributed Computing, 161(??):130–142, March 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002227.

Khalid:2022:CET

[KJA^{+}22] Muhammad Usman Khalid, Nadeem Javaid, Ahmad Almogren, Abrar Ahmed, Sardar Muhammad Gulfam, and Ayman Radwan. Cooperative energy transactions in micro and utility grids integrating energy storage systems. Journal of Parallel and Distributed Computing, 161(??):48–62, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002136.

Kumar:2021:RAU

[KK21] R. Anantha Kumar and K. Kartheeban. Resource allocation using Dynamic Pricing Auction Mechanism for supporting emergency demands in Cloud Computing. Journal of Parallel and Distributed Computing, 158(??):213–226, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001623.

Kadri:2022:MOB

[KK22] Nassima Kadri and Mouloud Koudil. Multi-objective biogeographybased optimization and reinforcement learning hybridization for network-on chip reliability improvement. Journal of Parallel and Distributed Computing, 161(??):20–36, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002124.

Kumar:2023:BOD

[KKG^{+}23] Prabhat Kumar, Randhir Kumar, Govind P. Gupta, Rakesh Tripathi, Alireza Jolfaei, and A. K. M. Najmul Islam. A blockchain-orchestrated deep learning approach for secure data transmission in IoT-enabled healthcare system. Journal of Parallel and Distributed Computing, 172(??):69–83, February 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002106.

Kao:2023:PAC

[KKH^{+}23] Shih-Shun Kao, Ralf Klasing, Ling-Ju Hung, Chia-Wei Lee, and Sun-Yuan Hsieh. A parallel algorithm for constructing multiple independent spanning trees in bubble-sort networks. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001016.

Kumar:2022:DID

[KKT^{+}22] Randhir Kumar, Prabhat Kumar, Rakesh Tripathi, Govind P. Gupta, Sahil Garg, and Mohammad Mehedi Hassan. A distributed intrusion detection system to detect DDoS attacks in blockchain-enabled IoT network. Journal of Parallel and Distributed Computing, 164(??):55–68, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000351.

Kokocinski:2023:CHA

[KKW23] Maciej Kokociński, Tadeusz Kobus, and Paweł T. Wojciechowski. On the correctness of highly available systems in the presence of failures. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000709.

Keller:2022:SSS

[KL22] Jörg Keller and Sebastian Litzinger. Systematic search space design for energy-efficient static scheduling of moldable tasks. Journal of Parallel and Distributed Computing, 162(??):44– 58, April 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000107.

Koo:2021:ESS

[KLL^{+}21] Donghun Koo, Jaehwan Lee, Jialin Liu, Eun-Kyu Byun, Jae-Hyuck Kwak, Glenn K. Lockwood, Soonwook Hwang, Katie Antypas, Kesheng Wu, and Hyeonsang Eom. An empirical study of I/O separation for burst buffers in HPC systems. Journal of Parallel and Distributed Computing, 148(??):96–108, February 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303907.

Kshemkalyani:2022:DMR

[KMS22] Ajay D. Kshemkalyani, Anisur Rahaman Molla, and Gokarna Sharma. Dispersion of mobile robots using global communication. Journal of Parallel and Distributed Computing, 161(??):100–117, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521002203.

Kawtikwar:2024:HHL

[KN24] Samiran Kawtikwar and Rakesh Nagi. HyLAC: Hybrid linear assignment solver in CUDA. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000029.

Kashi:2023:IBS

[KNK^{+}23] Aditya Kashi, Pratik Nayak, Dhruva Kulkarni, Aaron Scheinberg, Paul Lin, and Hartwig Anzt. Integrating batched sparse iterative solvers for the collision operator in fusion plasma simulations on GPUs. Journal of Parallel and Distributed Computing, 178(??):69–81, August 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000540.

Kim:2022:GDS

[KPS^{+}22] Yongho Kim, Seongha Park, Sean Shahkarami, Rajesh Sankaran, Nicola Ferrier, and Pete Beckman. Goal-driven scheduling model in edge computing for smart city applications. Journal of Parallel and Distributed Computing, 167(??):97–108, September 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522001009.

Khan:2020:EEA

[KRK20] Muhammad Nawaz Khan, Haseeb Ur Rahman, and Muhamad Zahid Khan. An Energy Efficient Adaptive Scheduling Scheme (EASS) for mesh grid wireless sensor networks. Journal of Parallel and Distributed Computing, 146(??):139–157, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303476.

Klinkenberg:2020:CRL

[KSB^{+}20] Jannis Klinkenberg, Philipp Samfass, Michael Bader, Christian Terboven, and Matthias S. Müller. CHAMELEON: Reactive load balancing for hybrid MPI + OpenMP task-parallel applications. Journal of Parallel and Distributed Computing, 138(??): 55–64, April 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519305180.

Krause:2024:HAR

[KSLN24] Arthur M. Krause, Paulo C. Santos, Arthur F. Lorenzon, and Philippe O. A. Navaux. HBPB, applying reuse distance to improve cache efficiency proactively. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000832.

Khodayarseresht:2023:MOC

[KSS23] Ehsan Khodayarseresht and Alireza Shameli-Sendi. A multiobjective cloud energy optimizer algorithm for federated environments. Journal of Parallel and Distributed Computing, 174(??):81–99, April 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522002520.

Khaund:2023:RFR

[KST^{+}23] Abhigyan Khaund, Abhishek Mukesh Sharma, Abhishek Tiwari, Shashwat Garg, and Sriram Kailasam. RD-FCA: a resilient distributed framework for formal concept analysis. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000734.

Kong:2020:PGB

[KSV^{+}20a] Wei Kong, Jian Shen, Pandi Vijayakumar, Youngju Cho, and Victor Chang. A practical group blind signature scheme for privacy protection in smart grid. Journal of Parallel and Distributed Computing, 136(??):29–39, February 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519301285.

Kshemkalyani:2020:PCE

[KSV20b] Ajay D. Kshemkalyani, Min Shen, and Bhargav Voleti. Prime clock: Encoded vector clock to characterize causality in distributed systems. Journal of Parallel and Distributed Computing, 140(??):37–51, June 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519304939.

Kumar:2021:SDD

[KTM^{+}21] Randhir Kumar, Rakesh Tripathi, Ningrinla Marchang, Gautam Srivastava, Thippa Reddy Gadekallu, and Neal N. Xiong. A secured distributed detection system based on IPFS and blockchain for industrial image and video data security. Journal of Parallel and Distributed Computing, 152(??):128–143, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000435.

Kwadjo:2022:TCB

[KTMB22] Danielle Tchuinkou Kwadjo, Erman Nghonda Tchinda, Joel Mandebi Mbongue, and Christophe Bobda. Towards a componentbased acceleration of convolutional neural networks on FP-GAs. Journal of Parallel and Distributed Computing, 167 (??):123–135, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001095.

Kuru:2021:MGD

[Kur21] Kaya Kuru. Management of geo-distributed intelligence: Deep Insight as a Service (DINSaaS) on Forged Cloud Platforms (FCP). Journal of Parallel and Distributed Computing, 149(??):103–118, March 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520304160.

Khan:2024:MOG

[KuR24] Minhaj Ahmad Khan and Raihan ur Rasool. A multi-objective grey-wolf optimization based approach for scheduling on cloud platforms. Journal of Parallel and Distributed Computing, 187 (??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373152400011X.

Kalaiarasi:2023:HPH

[KVMR23] M. Kalaiarasi, V. R. Venkatasubramani, M. S. K. Manikandan, and S. Rajaram. High performance HITA based Binary Edward Curve Crypto processor for FPGA platforms. Journal of Parallel and Distributed Computing, 178(??):56–68, August 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000503.

Khanchandani:2020:TEI

[KW20] Pankaj Khanchandani and Roger Wattenhofer. Two elementary instructions make compare-and-swap. Journal of Parallel and Distributed Computing, 145(??):176–187, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303129.

Kambatla:2020:OSS

[KYGG20] Karthik Kambatla, Vamsee Yarlagadda, Íñigo Goiri, and Ananth Grama. Optimistic scheduling with service guarantees. Journal of Parallel and Distributed Computing, 135(??):246– 258, January 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518305926.

Khalajzadeh:2020:CED

[KYZ^{+}20] Hourieh Khalajzadeh, Dong Yuan, Bing Bing Zhou, John Grundy, and Yun Yang. Cost effective dynamic data placement for efficient access of social networks. Journal of Parallel and Distributed Computing, 141(??):82–98, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518301710.

Lai:2021:OND

[Lai21] Cheng-Nan Lai. Optimal node-disjoint paths in folded hypercubes. Journal of Parallel and Distributed Computing, 147(??):100–107, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373152030366X.

LoBello:2020:SAT

[LAPB20] Lucia Lo Bello, Mohammad Ashjaei, Gaetano Patti, and Moris Behnam. Schedulability analysis of time-sensitive networks with scheduled traffic and preemption support. Journal of Parallel and Distributed Computing, 144(??):153–171, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303087.

Liu:2020:PSP

[LBHW20] Qin Liu, Md Zakirul Alam Bhuiyan, Jiankun Hu, and Jie Wu. Preface: Security & privacy in social big data. Journal of Parallel and Distributed Computing, 141(??):59–60, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152030215X.

Liao:2020:IMD

[LCC20] Ling Xia Liao, Han-Chieh Chao, and Mu-Yen Chen. Intelligently modeling, detecting, and scheduling elephant flows in software defined energy cloud: a survey. Journal of Parallel and Distributed Computing, 146(??):64–78, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303373.

Li:2021:PMS

[LCH^{+}21] Zhongjin Li, Victor Chang, Haiyang Hu, Dongjin Yu, Jidong Ge, and Binbin Huang. Profit maximization for securityaware task offloading in edge-cloud environment. Journal of Parallel and Distributed Computing, 157(??):43–55, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001283.

Lemeire:2023:AAP

[LCK23] Jan Lemeire, Jan G. Cornelis, and Elias Konstantinidis. Analysis of the analytical performance models for GPUs and extracting the underlying Pipeline model. Journal of Parallel and Distributed Computing, 173(??):32–47, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002295.

Li:2021:DSP

[LCW^{+}21] Zhuozhao Li, Ryan Chard, Logan Ward, Kyle Chard, Tyler J. Skluzacek, Yadu Babuji, Anna Woodard, Steven Tuecke, Ben Blaiszik, Michael J. Franklin, and Ian Foster. DL-Hub: Simplifying publication, discovery, and use of machine learning models in science. Journal of Parallel and Distributed Computing, 147(??):64–76, January 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303464.

Li:2021:FGA

[LCZL21] Guangsong Li, Wei Chen, Bin Zhang, and Siqi Lu. A finegrained anonymous handover authentication protocol based on consortium blockchain for wireless networks. Journal of Parallel and Distributed Computing, 157(??):157–167, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001362.

Lien:2021:RBK

[LD21] Li-Chuan Lien and Unurjargal Dolgorsuren. Rule-based knowledge discovery of satellite imagery using evolutionary classification tree. Journal of Parallel and Distributed Computing, 147(??):132–139, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520303646.

Lv:2024:CAF

[LFC^{+}24] Mengjie Lv, Jianxi Fan, Baolei Cheng, Jia Yu, and Xiaojua Jia. Construction algorithms of fault-tolerant paths and disjoint paths in k-ary n-cube networks. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001314.

Li:2020:EAI

[LFJ^{+}20] Liandeng Li, Jiarui Fang, Jinlei Jiang, Lin Gan, Weijie Zheng, Haohuan Fu, and Guangwen Yang. Efficient AES implementation on Sunway TaihuLight supercomputer: a systematic approach. Journal of Parallel and Distributed Computing, 138(??):178–189, April 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519301108.

Lopez-Fernandez:2021:MGB

[LFRBGV^{+}21] Aurelio Lopez-Fernandez, Domingo Rodriguez-Baena, Francisco Gomez-Vela, Federico Divina, and Miguel Garcia-Torres. A multi-GPU biclustering algorithm for binary datasets. Journal of Parallel and Distributed Computing, 147(??):209–219, January 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303701.

Long:2023:DLB

[LGZZ23] Xinjian Long, Xiangyang Gong, Bo Zhang, and Huiyang Zhou. Deep learning based data prefetching in CPU-GPU unified virtual memory. Journal of Parallel and Distributed Computing, 174(??):19–31, April 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002490.

Liu:2021:BAB

[LHL21] Han Liu, Dezhi Han, and Dun Li. Behavior analysis and blockchain based trust management in VANETs. Journal of Parallel and Distributed Computing, 151(??):61–69, May 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000320.

Li:2022:HMG

[Li22a] Jian-Guo Li. Hybrid multi-grid parallelisation of WAVE-WATCH III model on spherical multiple-cell grids. Journal of Parallel and Distributed Computing, 167(??):187–198, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001113.

Li:2022:DIC

[Li22b] Keqin Li. Distributed and individualized computation offloading optimization in a fog computing environment. Journal of Parallel and Distributed Computing, 159(??):24–34, January 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001817.

Li:2023:SET

[Li23] Yi Li. A secure and efficient three-factor authentication protocol for IoT environments. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000849.

Li:2024:SIT

[Li24] Keqin Li. Scheduling independent tasks on multiple cloudassisted edge servers with energy constraint. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152300151X.

Liu:2022:DLB

[LJH^{+}22] Bowen Liu, Xutong Jiang, Xin He, Lianyong Qi, Xiaolong Xu, Xiaokang Wang, and Wanchun Dou. A deep learning-based edge caching optimization method for costdriven planning process over IIoT. Journal of Parallel and Distributed Computing, 168(??):80–89, October 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001435.

Liu:2022:ZRM

[LJW^{+}22] Wei Liu, Jiangming Jin, Hao Wu, Yifan Gong, Ziyue Jiang, and Jidong Zhai. Zoro: a robotic middleware combining high performance and high reliability. Journal of Parallel and Distributed Computing, 166(??):126–138, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000879.

Lee:2022:ISP

[LKAB^{+}22] Sunwoo Lee, Qiao Kang, Reda Al-Bahrani, Ankit Agrawal, Alok Choudhary, and Wei keng Liao. Improving scalability of parallel CNN training by adaptively adjusting parameter update frequency. Journal of Parallel and Distributed Computing, 159(??):10–23, January 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001830.

Liu:2020:HEP

[LLC20] Wei Liu, Wei Li, and Yong Beom Cho. High-efficiency parallelism solution for a Multiview High-Efficiency Video Coding decoder. Journal of Parallel and Distributed Computing, 141(??): 115–128, July 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151930214X.

Lin:2023:FCR

[LLD^{+}23] Wenmin Lin, Hui Leng, Ruihan Dou, Lianyong Qi, Zhigeng Pan, and Md. Arafatur Rahman. A federated collaborative recommendation model for privacy-preserving distributed recommender applications based on microservice framework. Journal of Parallel and Distributed Computing, 174(??):70–80, April 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002477.

Li:2020:SII

[LLF^{+}20] Fenghua Li, Zifu Li, Liang Fang, Yongjun Li, Yaobing Xu, and Yunchuan Guo. Securing instruction interaction for hierarchical management. Journal of Parallel and Distributed Computing, 137(??):91–103, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303144.

Li:2024:ABL

[LLG^{+}24] Yue Li, Han Liu, Jianbo Gao, Jiashuo Zhang, Zhi Guan, and Zhong Chen. Accelerating block lifecycle on blockchain via hardware transactional memory. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001491.

Li:2022:SDT

[LLGC22] Xiao-Yan Li, Wanling Lin, Wenzhong Guo, and Jou-Ming Chang. A secure data transmission scheme based on multiprotection routing in datacenter networks. Journal of Parallel and Distributed Computing, 167(??):222–231, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001253.

Laccetti:2020:PED

[LLM^{+}20] Giuliano Laccetti, Marco Lapegna, Valeria Mele, Diego Romano, and Lukasz Szustak. Performance enhancement of a dynamic K-means algorithm through a parallel adaptive strategy on multicore CPUs. Journal of Parallel and Distributed Computing, 145(??):34–41, November 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303178.

Li:2024:RAA

[LLM24] Si-Yu Li, Xiang-Jun Li, and Meijie Ma. Reliability assessment for k-ary n-cubes with faulty edges. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000509.

Lahderanta:2021:ECS

[LLR^{+}21] Tero Lähderanta, Teemu Leppänen, Leena Ruha, Lauri Lovén, Erkki Harjula, Mika Ylianttila, Jukka Riekki, and Mikko J. Sillanpää. Edge computing server placement with capacitated location allocation. Journal of Parallel and Distributed Computing, 153(??):130–149, July 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000605.

Liu:2020:ADV

[LLW^{+}20] Xiao Liu, Anfeng Liu, Tian Wang, Kaoru Ota, Mianxiong Dong, Yuxin Liu, and Zhiping Cai. Adaptive data and verified message disjoint security routing for gathering big data in energy harvesting networks. Journal of Parallel and Distributed Computing, 135(??):140–155, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830964X.

Li:2023:LNF

[LLX^{+}23] Xiaoming Li, Yi Luo, Neal Xiong, Wei Yu, Guangquan Xu, Changzheng Liu, and Xiaoping Yang. Local node feature modeling for edge computing based on network embedding in dynamic networks. Journal of Parallel and Distributed Computing, 171(??):98–110, January 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002027.

Li:2021:ICH

[LLXG21] Ming Li, Qiang Li, Guangzhe Xuan, and Dong Guo. Identifying compromised hosts under APT using DNS request sequences. Journal of Parallel and Distributed Computing, 152(??):67– 78, June 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000381.

Li:2024:FRL

[LLXX24] Qiliang Li, Min Lyu, Liangliang Xu, and Yinlong Xu. Fast recovery for large disk enclosures based on RAID2.0: Algorithms and evaluation. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731524000182.

Liao:2023:OCO

[LLYZ23] Linbo Liao, Yongxuan Lai, Fan Yang, and Wenhua Zeng. Online computation offloading with double reinforcement learning algorithm in mobile edge computing. Journal of Parallel and Distributed Computing, 171(??):28–39, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001976.

Li:2021:HEM

[LMG^{+}21] Biwei Li, Zhenqiang Mi, Yu Guo, Yang Yang, and Mohammad S. Obaidat. A high efficient multi-robot simultaneous localization and mapping system using partial computing offloading assisted cloud point registration strategy. Journal of Parallel and Distributed Computing, 149(??):89–102, March 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304019.

Lasserre:2021:EFL

[LNW21] Alice Lasserre, Raymond Namyst, and Pierre-André Wacrenier. EasyPAP: a framework for learning parallel programming. Journal of Parallel and Distributed Computing, 158(??):94–114, December 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521001647.

Lee:2024:FKG

[LoKS24] Chun-Hee Lee, Dong oh Kang, and Hwa Jeon Song. Fast knowledge graph completion using graphics processing units. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000492.

Liang:2020:PPR

[LQX^{+}20] Jinwen Liang, Zheng Qin, Sheng Xiao, Jixin Zhang, Hui Yin, and Keqin Li. Privacy-preserving range query over multisource electronic health records in public clouds. Journal of Parallel and Distributed Computing, 135(??):127–139, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151930053X.

Li:2020:TSG

[LRV20] Zhongmiao Li, Paolo Romano, and Peter Van Roy. Transparent speculation in geo-replicated transactional data stores. Journal of Parallel and Distributed Computing, 143(??):129–147, September 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520302768.

Lin:2022:EHH

[LSC22] Yu-Hsiang Lin, Wen-Chi Shih, and Yeim-Kuan Chang. Efficient hierarchical hash tree for OpenFlow packet classification with fast updates on GPUs. Journal of Parallel and Distributed Computing, 167(??):136–147, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200096X.

Li:2020:STF

[LSWY20] Guangshun Li, Xiaofei Sheng, Junhua Wu, and Haili Yu. Securing transmissions by friendly jamming scheme in wireless networks. Journal of Parallel and Distributed Computing, 144(??):260–267, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520302720.

Li:2020:ERM

[LSZL20] Chunlin Li, Mingyang Song, Min Zhang, and Youlong Luo. Effective replica management for improving reliability and availability in edge-cloud computing environment. Journal of Parallel and Distributed Computing, 143(??):107–128, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302719.

Liu:2020:ABW

[LTBY20] Lin Liu, Wei-Tek Tsai, Md Zakirul Alam Bhuiyan, and Dong Yang. Automatic blockchain whitepapers analysis via heterogeneous graph neural network. Journal of Parallel and Distributed Computing, 145(??):1–12, November 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302938.

Lin:2020:GTD

[LW20] Huanxin Lin and Cho-Li Wang. On-GPU thread-data remapping for nested branch divergence. Journal of Parallel and Distributed Computing, 139(??):75–86, May 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308967.

Lu:2020:TCE

[LWD^{+}20] Shuaibing Lu, Jie Wu, Yubin Duan, Ning Wang, and Juan Fang. Towards cost-efficient resource provisioning with multiple mobile users in fog computing. Journal of Parallel and Distributed Computing, 146(??):96–106, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303427.

Li:2022:MDO

[LWHF22] Huifang Li, Yizhu Wang, Jingwei Huang, and Yushun Fan. Mutation and dynamic objective-based farmland fertility algorithm for workflow scheduling in the cloud. Journal of Parallel and Distributed Computing, 164(??):69–82, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000478.

Li:2022:ZSB

[LWL^{+}22] Yuxian Li, Jian Weng, Ming Li, Wei Wu, Jiasi Weng, Jia-Nan Liu, and Shun Hu. ZeroCross: a sidechain-based privacypreserving cross-chain solution for Monero. Journal of Parallel and Distributed Computing, 169(??):301–316, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001733.

Li:2023:PPB

[LWW^{+}23] Yuxian Li, Jian Weng, Wei Wu, Ming Li, Yingjiu Li, Haoxin Tu, Yongdong Wu, and Robert H. Deng. PRI: PCHbased privacy-preserving with reusability and interoperability for enhancing blockchain scalability. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000916.

Lin:2022:UIC

[LXC^{+}22] Longxin Lin, Zhenxiong Xu, Chien-Ming Chen, Ke Wang, Md. Rafiul Hassan, Md. Golam Rabiul Alam, Mohammad Mehedi Hassan, and Giancarlo Fortino. Understanding the impact on convolutional neural networks with different model scales in AIoT domain. Journal of Parallel and Distributed Computing, 170(??):1–12, December 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001769.

Liu:2023:IEE

[LYC23] Jing Liu, Pei Yang, and Cen Chen. Intelligent energyefficient scheduling with ant colony techniques for heterogeneous edge computing. Journal of Parallel and Distributed Computing, 172(??):84–96, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002131.

Li:2024:EEO

[LYF^{+}24] Zengpeng Li, Huiqun Yu, Guisheng Fan, Jiayin Zhang, and Jin Xu. Energy-efficient offloading for DNN-based applications in edge-cloud computing: a hybrid chaotic evolutionary approach. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000145.

Lin:2022:GBM

[LYZ^{+}22a] Weiwei Lin, Kun Yao, Lan Zeng, Fagui Liu, Chun Shan, and Xiaobin Hong. A GAN-based method for time-dependent cloud workload generation. Journal of Parallel and Distributed Computing, 168(??):33–44, October 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200123X.

Liu:2022:MMB

[LYZ^{+}22b] Shun Liu, Qiang Yang, Shaobo Zhang, Tian Wang, and Neal N. Xiong. MIDP: an MDP-based intelligent big data processing scheme for vehicular edge computing. Journal of Parallel and Distributed Computing, 167(??):1–17, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000880.

Li:2021:WSC

[LZD21] Xiaoyun Li, Zibin Zheng, and Hong-Ning Dai. When services computing meets blockchain: Challenges and opportunities. Journal of Parallel and Distributed Computing, 150(??): 1–14, April 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520304196.

Li:2022:ELT

[LZGL22] Chunlin Li, Yong Zhang, Xiang Gao, and Youlong Luo. Energylatency tradeoffs for edge caching and dynamic service migration based on DQN in mobile edge computing. Journal of Parallel and Distributed Computing, 166(??):15–31, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000569.

Li:2022:IDP

[LZL22] Chunlin Li, Yong Zhang, and Youlong Luo. Intermediate data placement and cache replacement strategy under Spark platform. Journal of Parallel and Distributed Computing, 163(??): 114–135, May 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000260.

Li:2024:IAO

[LZS^{+}24] Hongliang Li, Hairui Zhao, Ting Sun, Xiang Li, Haixiao Xu, and Keqin Li. Interference-aware opportunistic job placement for shared distributed deep learning clusters. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001466.

Liu:2022:CDT

[LZWZ22] Jiafei Liu, Shuming Zhou, Dajin Wang, and Hong Zhang. Component diagnosability in terms of component connectivity of hypercube-based compound networks. Journal of Parallel and Distributed Computing, 162(??):17–26, April 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002252.

Luo:2023:RRA

[LZY23] Ming Luo, Jie Zhou, and Ping Yang. RATS: a regulatory anonymous transaction system based on blockchain. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001211.

Liu:2020:NRV

[LZZ^{+}20] Yaping Liu, Shuo Zhang, Haojin Zhu, Peng-Jun Wan, Lixin Gao, Yaoxue Zhang, and Zhihong Tian. A novel routing verification approach based on blockchain for inter-domain routing in smart metropolitan area networks. Journal of Parallel and Distributed Computing, 142(??):77–89, August 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519308317.

Li:2022:MPP

[LZZ^{+}22] Junli Li, Chaowei Zhang, Jifu Zhang, Xiao Qin, and Lihua Hu. MiCS-P:Parallel mutual-information computation of big categorical data on Spark. Journal of Parallel and Distributed Computing, 161(??):118–129, March 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002239.

Li:2023:NSE

[LZZJ23] Xiao-Yan Li, Kun Zhao, Hongbin Zhuang, and Xiaohua Jia. Novel schemes for embedding Hamiltonian paths and cycles in balanced hypercubes with exponential faulty edges. Journal of Parallel and Distributed Computing, 177(??):182–191, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000436.

Mallach:2021:NLM

[Mal21] Sven Mallach. A note on labeling methods to schedule unit execution time tasks in the presence of delayed precedence constraints. Journal of Parallel and Distributed Computing, 156(??):1–6, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521001076.

Malkis:2022:RPP

[Mal22] Alexander Malkis. Reachability in parallel programs is polynomial in the number of threads. Journal of Parallel and Distributed Computing, 162(??):1–16, April 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002215.

Mondal:2023:TLR

[MAL^{+}23] Sagnik Mondal, Sameh Abdulah, Hatem Ltaief, Ying Sun, Marc G. Genton, and David E. Keyes. Tile low-rank approximations of non-Gaussian space and space-time Tukey g-and-h random field likelihoods and predictions on large-scale systems. Journal of Parallel and Distributed Computing, 180(??): ??, October 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000850.

Marowka:2020:PDB

[Mar20] Ami Marowka. On the performance difference between theory and practice for parallel algorithms. Journal of Parallel and Distributed Computing, 138(??):199–210, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304964.

Moreno:2021:IMR

[MAR21] Pedro Moreno, Miguel Areias, and Ricardo Rocha. On the implementation of memory reclamation methods in a lock-free hash trie design. Journal of Parallel and Distributed Computing, 155(??):1–13, September 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000903.

Mansouri:2021:REC

[MB21] Yaser Mansouri and M. Ali Babar. A review of edge computing: Features and resource virtualization. Journal of Parallel and Distributed Computing, 150(??):155–183, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304317.

Mirsaraei:2022:STF

[MBB22] AmirHossein Ghafouri Mirsaraei, Ali Barati, and Hamid Barati. A secure three-factor authentication scheme for IoT environments. Journal of Parallel and Distributed Computing, 169(??):87–105, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001460.

M:2020:LBE

[MBM^{+}20] Swarna Priya R. M., Sweta Bhattacharya, Praveen Kumar Reddy Maddikunta, Siva Rama Krishnan Somayaji, Kuruva Lakshmanna, Rajesh Kaluri, Aseel Hussien, and Thippa Reddy Gadekallu. Load balancing of energy cloud using wind driven and firefly algorithms in Internet of Everything. Journal of Parallel and Distributed Computing, 142(??):16–26, August 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520300356.

Monteleone:2022:DMM

[MBN22] Alessandra Monteleone, Gaetano Burriesci, and Enrico Napoli. A distributed-memory MPI parallelization scheme for multidomain incompressible SPH. Journal of Parallel and Distributed Computing, 170(??):53–67, December 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001885.

Moran:2024:EES

[MBRR24] Marina Morán, Javier Balladini, Dolores Rexachs, and Enzo Rucci. Exploring energy saving opportunities in fault tolerant HPC systems. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523001673.

Martinez:2020:RVS

[MBS^{+}20] Daniel Martínez, Wesley Brewer, Andrew Strelzoff, Andrew Wilson, and Daniel Wade. Rotorcraft virtual sensors via deep regression. Journal of Parallel and Distributed Computing, 135(??):114–126, January 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300413.

Mukherjee:2020:TAD

[MCC20] Priyajit Mukherjee, Navonil Chatterjee, and Santanu Chattopadhyay. Thermal-aware detour routing in 3D NoCs. Journal of Parallel and Distributed Computing, 144(??):230–245, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302690.

Maneas:2021:AIC

[MCD^{+}21] Stathis Maneas, Nikos Chondros, Panos Diamantopoulos, Christos Patsonakis, and Mema Roussopoulos. On achieving interactive consistency in real-world distributed systems. Journal of Parallel and Distributed Computing, 147(??):220–235, January 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303713.

Mabed:2020:SLT

[MD20] Hakim Mabed and Eugen Dedu. Short and long term optimization for micro-object conveying with air-jet modular distributed system. Journal of Parallel and Distributed Computing, 144(??):98–108, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302951.

Marszalkowski:2020:TET

[MDS20] Jedrzej Marszałkowski, Maciej Drozdowski, and Gaurav Singh. Time-energy trade-offs in processing divisible loads on heterogeneous hierarchical memory systems. Journal of Parallel and Distributed Computing, 144(??):206–219, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152030294X.

Martinez-Ferrer:2023:IPC

[MFAB23] Pedro J. Martinez-Ferrer, Tufan Arslan, and Vicenç Beltran. Improving the performance of classical linear algebra iterative methods via hybrid parallelism. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000746.

Mirsky:2020:LCA

[MGE20] Yisroel Mirsky, Tomer Golomb, and Yuval Elovici. Lightweight collaborative anomaly detection for the IoT using blockchain. Journal of Parallel and Distributed Computing, 145(??):75–97, November 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303154.

Mao:2024:ATI

[MGW24] Chunyu Mao, Wojciech Golab, and Bernard Wong. Antipaxos: Taking interactive consistency to the next level. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000030.

Morales:2023:IMT

[MHdC^{+}23] Catalina Munoz Morales, Bruno Honorio, Joao P. L. de Carvalho, Alexandro Baldassin, and Guido Araujo. On the impact of mode transition on phased transactional memory performance. Journal of Parallel and Distributed Computing, 173(??):126–139, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522002428.

Meyer:2022:DFA

[MKP22] Marius Meyer, Tobias Kenter, and Christian Plessl. Indepth FPGA accelerator performance evaluation with single node benchmarks from the HPC challenge benchmark suite for Intel and Xilinx FPGAs using OpenCL. Journal of Parallel and Distributed Computing, 160(??):79–89, February 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002057.

Mecheraoui:2021:PNE

[MLB21] Khalil Mecheraoui, Irina A. Lomazova, and Nabil Belala. A Petri net extension for systems of concurrent communicating agents with durable actions. Journal of Parallel and Distributed Computing, 155(??):14–23, September 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000940.

Molina:2021:ETD

[MLGC^{+}21] Romina Molina, Fernando Loor, Veronica Gil-Costa, Franco Maria Nardini, Raffaele Perego, and Salvatore Trani. Efficient traversal of decision tree ensembles with FPGAs. Journal of Parallel and Distributed Computing, 155(??):38–49, September 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000915.

Mwasinga:2023:RRA

[MLR^{+}23] Lusungu Josh Mwasinga, Duc-Tai Le, Syed M. Raza, Rajesh Challa, Moonseong Kim, and Hyunseung Choo. RASM: Resource-aware service migration in edge computing based on deep reinforcement learning. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001156.

Meng:2020:TBE

[MLTT20] Weizhi Meng, Wenjuan Li, Steven Tug, and Jiao Tan. Towards blockchain-enabled single character frequency-based exclusive signature matching in IoT-assisted smart cities. Journal of Parallel and Distributed Computing, 144(??):268–277, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302926.

Macedo:2024:APG

[MM24] Emerson A. Macedo and Alba C. M. A. Melo. Adaptive patch grid strategy for parallel protein folding using atomic burials with NAMD. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731524000327.

Monifi:2022:TVD

[MMA22] Elham Monifi and Nezam Mahdavi-Amiri. Time-varying dual accelerated gradient ascent: a fast network optimization algorithm. Journal of Parallel and Distributed Computing, 165(??): 130–141, July 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000661.

Maglione-Mathey:2021:LIC

[MMESG^{+}21] German Maglione-Mathey, Jesus Escudero-Sahuquillo, Pedro Javier Garcia, Francisco J. Quiles, and Eitan Zahavi. Leveraging InfiniBand controller to configure deadlock-free routing engines for Dragonflies. Journal of Parallel and Distributed Computing, 147(??):16–33, January 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303397.

Mullen:2021:TLH

[MMM21] Julia Mullen, Lauren Milechin, and Dennis Milechin. Teaching and learning HPC through serious games. Journal of Parallel and Distributed Computing, 158(??):115–125, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100160X.

Mostefaoui:2022:SIC

[MMM22] Ahmed Mostefaoui, Geyong Min, and Peter Müller. Special issue: Connected vehicles meet big data technologies: Recent advances and future trends. Journal of Parallel and Distributed Computing, 167(??):221, September 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001277.

Mahajan:2024:GAT

[MN24] Mohit Mahajan and Rakesh Nagi. GPU-accelerated transportation simplex algorithm. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0743731523001600.

Moradi:2024:TTD

[MPAS24] Fereidoun Moradi, Bahman Pourvatan, Sara Abbaspour Asadollah, and Marjan Sirjani. Tiny Twins for detecting cyberattacks at runtime using concise Rebeca time transition system. Journal of Parallel and Distributed Computing, 184(??): ??, February 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001508.

Medeiros:2021:MPA

[MPR^{+}21] Thiarles S. Medeiros, Luan Pereira, Fábio D. Rossi, Marcelo C. Luizelli, Antonio Carlos S. Beck, and Arthur F. Lorenzon. Mitigating the processor aging through dynamic concurrency throttling. Journal of Parallel and Distributed Computing, 156(??):86–100, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521001118.

Mirmahaleh:2020:FMM

[MRB20] Seyedeh Yasaman Hosseini Mirmahaleh, Midia Reshadi, and Nader Bagherzadeh. Flow mapping on mesh-based deep learning accelerator. Journal of Parallel and Distributed Computing, 144(??):80–97, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302707.

Meftah:2021:EMC

[MRC21] Lakhdar Meftah, Romain Rouvoy, and Isabelle Chrisment. Empowering mobile crowdsourcing apps with user privacy control. Journal of Parallel and Distributed Computing, 147(??):1–15, January 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303403.

Mohan:2020:OPR

[MRPH20] Lakshmi J. Mohan, Ketan Rajawat, Udaya Parampalli, and Aaron Harwood. Optimal placement for repair-efficient erasure codes in geo-diverse storage centres. Journal of Parallel and Distributed Computing, 135(??):101–113, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302187.

Marowka:2020:ESI

[MS20] Ami Marowka and Przemysław Stpiczyński. Editorial on the special issue on advances in parallel programming: Languages, models and algorithms. Journal of Parallel and Distributed Computing, 140(??):63–64, June 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520301830.

Monteiro:2023:PCF

[MS23] Ana M. Monteiro and António A. F. Santos. Parallel computing in finance for estimating risk-neutral densities through option prices. Journal of Parallel and Distributed Computing, 173(??): 61–69, March 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373152200243X.

Mahmud:2020:PAA

[MSRB20] Redowan Mahmud, Satish Narayana Srirama, Kotagiri Ramamohanarao, and Rajkumar Buyya. Profit-aware application placement for integrated fog–cloud computing environments. Journal of Parallel and Distributed Computing, 135(??):177– 190, January 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519300346.

Mukosey:2024:GBR

[MST24] Anatoly Mukosey, Alexander Semenov, and Aleksandr Tretiakov. Graph based routing algorithm for torus topology and its evaluation for the angara interconnect. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001351.

Mencagli:2024:GPD

[MTG^{+}24] Gabriele Mencagli, Massimo Torquati, Dalvan Griebler, Alessandra Fais, and Marco Danelutto. General-purpose data stream processing on heterogeneous architectures with Wind-Flow. Journal of Parallel and Distributed Computing, 184(??): ??, February 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001521.

Mendes:2022:DGV

[MTR22] Francisco Mendes, Pedro Tomás, and Nuno Roma. Decoupling GPGPU voltage-frequency scaling for deep-learning applications. Journal of Parallel and Distributed Computing, 165(??): 32–51, July 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000624.

Mommessin:2023:AAR

[MYS^{+}23] Clément Mommessin, Renyu Yang, Natalia V. Shakhlevich, Xiaoyang Sun, Satish Kumar, Junqing Xiao, and Jie Xu. Affinity-aware resource provisioning for longrunning applications in shared clusters. Journal of Parallel and Distributed Computing, 177(??):1–16, July 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000321.

Ma:2021:ETM

[MZMM21] Wenhuan Ma, Mingzu Zhang, Jixiang Meng, and Tianlong Ma. Exponential type of many-to-many edge disjoint paths on ternary n-cubes. Journal of Parallel and Distributed Computing, 158(??):67–79, December 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001660.

Ma:2021:ESS

[MZZW21] Zhanyou Ma, Changzhen Zhang, Liyuan Zhang, and Shunzhi Wang. Energy saving strategy and Nash equilibrium of hybrid P2P networks. Journal of Parallel and Distributed Computing, 157(??):145–156, November 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001453.

Nakka:2023:CEC

[NA23] Kalyan Nakka and Habib M. Ammari. k-CSqu: Ensuring connected k-coverage using cusp squares of square tessellation. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001193.

Nunes:2023:CHS

[NCR23] Diogo Nunes, Daniel Castro, and Paolo Romano. CSMV: a highly scalable multi-versioned software transactional memory for GPUs. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523000576.

Nived:2023:MPI

[NE23] M. R. Nived and Vinayak Eswaran. A massively parallel implicit 3D unstructured grid solver for computing turbulent flows on latest distributed memory computational architectures. Journal of Parallel and Distributed Computing, 182(??):??, December 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152300120X.

Nawrocki:2021:ARP

[NGS21] Piotr Nawrocki, Mikolaj Grzywacz, and Bartlomiej Sniezynski. Adaptive resource planning for cloud-based services using machine learning. Journal of Parallel and Distributed Computing, 152(??):88–97, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521000393.

Niknia:2022:SBA

[NHR22] Farnaz Niknia, Vesal Hakami, and Kiamehr Rezaee. An SMDP-based approach to thermal-aware task scheduling in NoC-based MPSoC platforms. Journal of Parallel and Distributed Computing, 165(??):79–106, July 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000715.

Nuriyev:2024:SSU

[NMA^{+}24] Emin Nuriyev, Ravi Reddy Manumachu, Samar Aseeri, Mahendra K. Verma, and Alexey L. Lastovetsky. SUARA: a scalable universal allreduce communication algorithm for acceleration of parallel deep learning applications. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001375.

Nejat:2020:CMD

[NMPS20] Mehrzad Nejat, Madhavan Manivannan, Miquel Pericàs, and Per Stenström. Coordinated management of DVFS and cache partitioning under QoS constraints to save energy in multi-core systems. Journal of Parallel and Distributed Computing, 144(??):246–259, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302859.

Nascimento:2020:ERL

[NPAM20] Dimas Cassimiro Nascimento, Carlos Eduardo Santos Pires, Tiago Brasileiro Araujo, and Demetrio Gomes Mestre. Estimating record linkage costs in distributed environments. Journal of Parallel and Distributed Computing, 143(??):97–106, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302756.

Netti:2023:MPS

[NPO^{+}23] Alessio Netti, Yang Peng, Patrik Omland, Michael Paulitsch, Jorge Parra, Gustavo Espinosa, Udit Agarwal, Abraham Chan, and Karthik Pattabiraman. Mixed precision support in HPC applications: What about reliability? Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001168.

Nguyen:2023:ESI

[NPY^{+}23] Truong Thao Nguyen, Kien Trung Pham, Hiroshi Yamaguchi, Yutaka Urino, and Michihiro Koibuchi. Effective switchless inter-FPGA memory networks. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000837.

Naranjo:2020:ASC

[NRdA^{+}20] Diana M. Naranjo, Sebastián Risco, Carlos de Alfonso, Alfonso Pérez, Ignacio Blanquer, and Germán Moltó. Accelerated serverless computing based on GPU virtualization. Journal of Parallel and Distributed Computing, 139(??):32–42, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303533.

Nuriyev:2022:MBS

[NRGL22] Emin Nuriyev, Juan-Antonio Rico-Gallego, and Alexey Lastovetsky. Model-based selection of optimal MPI broadcast algorithms for multi-core clusters. Journal of Parallel and Distributed Computing, 165(??):1–16, July 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000697.

Nan:2020:VAP

[NT20] Lihao Nan and Dacheng Tao. Variational approach for privacy funnel optimization on continuous data. Journal of Parallel and Distributed Computing, 137(??):17–25, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300899.

Ngoc:2023:HUA

[NTT^{+}23] Tu Dinh Ngoc, Boris Teabe, Alain Tchana, Gilles Muller, and Daniel Hagimont. HyperTP: a unified approach for live hypervisor replacement in datacenters. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152300103X.

Nguyen:2021:SBE

[NVE^{+}21] Gia Nhu Nguyen, Nin Ho Le Viet, Mohamed Elhoseny, K. Shankar, B. B. Gupta, and Ahmed A. Abd El-Latif. Secure blockchain enabled cyber-physical systems in healthcare using deep belief network with ResNet model. Journal of Parallel and Distributed Computing, 153(??):150–160, July 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000642.

Nosouhi:2020:BSL

[NYZ^{+}20] Mohammad Reza Nosouhi, Shui Yu, Wanlei Zhou, Marthie Grobler, and Habiba Keshtiar. Blockchain for secure location verification. Journal of Parallel and Distributed Computing, 136(??):40–51, February 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373151930320X.

Ni:2023:EZK

[NZ23] Ning Ni and Yongxin Zhu. Enabling zero knowledge proof by accelerating zk-SNARK kernels on GPU. Journal of Parallel and Distributed Computing, 173(??):20–31, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002246.

Ouyang:2021:COS

[ODXX21] Shuo Ouyang, Dezun Dong, Yemao Xu, and Liquan Xiao. Communication optimization strategies for distributed deep neural network training: a survey. Journal of Parallel and Distributed Computing, 149(??):52–65, March 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304068.

Orhean:2024:SSC

[OGR^{+}24] Alexandru Iulian Orhean, Anna Giannakou, Lavanya Ramakrishnan, Kyle Chard, Boris Glavic, and Ioan Raicu. SCIPIS: Scalable and concurrent persistent indexing and search in high-end computing systems. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152400042X.

Ogundoyin:2021:PPP

[OK21] Sunday Oyinlola Ogundoyin and Ismaila Adeniyi Kamil. PAASH: a privacy-preserving authentication and fine-grained access control of outsourced data for secure smart health in smart cities. Journal of Parallel and Distributed Computing, 155(??):101–119, September 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373152100099X.

Odegbile:2023:PET

[OMCW23] Olufemi Odegbile, Chaoyi Ma, Shigang Chen, and Yuanda Wang. Policy enforcement in traditional non-SDN networks. Journal of Parallel and Distributed Computing, 177(??):39– 52, July 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000205.

Orr:2020:OTS

[OS20] Michael Orr and Oliver Sinnen. Optimal task scheduling benefits from a duplicate-free state-space. Journal of Parallel and Distributed Computing, 146(??):158–174, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303348.

Ouyang:2022:CAC

[OZ22] Xiangzhen Ouyang and Yian Zhu. Core-aware combining: Accelerating critical section execution on heterogeneous multi-core systems via combining synchronization. Journal of Parallel and Distributed Computing, 162(??):27–43, April 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000077.

Palchaudhuri:2022:FFC

[PAD22] Ayan Palchaudhuri, Digvijay Anand, and Anindya Sundar Dhar. FPGA fabric conscious architecture design and automation of speed-area efficient Margolus neighborhood based cellular automata with variegated scan path insertion. Journal of Parallel and Distributed Computing, 167(??):50–63, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000983.

Palkowski:2020:PTC

[PB20] Marek Palkowski and Wlodzimierz Bielecki. Parallel tiled cache and energy efficient codes for o(n_{4}) RNA folding algorithms. Journal of Parallel and Distributed Computing, 137(??):252– 258, March 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519304873.

Popov:2021:FBF

[PB21] Serguei Popov and William J. Buchanan. FPC-BI: Fast probabilistic consensus within Byzantine infrastructures. Journal of Parallel and Distributed Computing, 147(??):77–86, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303634.

Padhy:2021:MCA

[PC21] Satyajit Padhy and Jerry Chou. MIRAGE: a consolidation aware migration avoidance genetic job scheduling algorithm for virtualized data centers. Journal of Parallel and Distributed Computing, 154(??):106–118, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000575.

Pai:2020:PRS

[PCC20] Kung-Jui Pai, Ruay-Shiung Chang, and Jou-Ming Chang. A protection routing with secure mechanism in Möbius cubes. Journal of Parallel and Distributed Computing, 140(??):1–12, June 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731518308761.

Palchaudhuri:2021:SAO

[PD21] Ayan Palchaudhuri and Anindya Sundar Dhar. Speed-area optimized VLSI architecture of multi-bit cellular automaton cell based random number generator on FPGA with testable logic support. Journal of Parallel and Distributed Computing, 151(??):13–23, May 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521000058.

Peretz:2022:FPM

[PF22] Yossi Peretz and Yigal Fischler. A fast parallel max-flow algorithm. Journal of Parallel and Distributed Computing, 169(??):226–241, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001691.

Prikopa:2020:FTL

[PG20] Karl E. Prikopa and Wilfried N. Gansterer. Fault-tolerant least squares solvers for wireless sensor networks based on gossiping. Journal of Parallel and Distributed Computing, 136(??):52–62, February 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519306252.

Prasad:2022:KTT

[PGB^{+}22] Sushil Prasad, Sheikh Ghafoor, Martina Barnas, Felix Wolf, Erik Saule, Noemi Rodriguez, and Rizos Sakellariou. Keeping up with technology: Teaching parallel, distributed, and high-performance computing. Journal of Parallel and Distributed Computing, 160(??):36–38, February 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001957.

Padmaa:2022:OCG

[PJV^{+}22] M. Padmaa, T. Jayasankar, S. Venkatraman, Ashit Kumar Dutta, Deepak Gupta, Shahab Shamshirband, and Joel J. P. C. Rodrigues. Oppositional chaos game optimization based clustering with trust based data transmission protocol for intelligent IoT edge systems. Journal of Parallel and Distributed Computing, 164(??):142–151, June 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200065X.

Plimpton:2021:RAL

[PK21a] Steven J. Plimpton and Christopher Knight. Rendezvous algorithms for large-scale modeling and simulation. Journal of Parallel and Distributed Computing, 147(??):184–195, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303622.

Przybyla-Kasperek:2021:PAD

[PK21b] Małgorzata Przybyła-Kasperek. The power of agents in a dispersed system — the Shapley–Shubik power index. Journal of Parallel and Distributed Computing, 157(??):105–124, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001465.

Perotin:2024:MRS

[PKSR24] Lucas Perotin, Sandhya Kandaswamy, Hongyang Sun, and Padma Raghavan. Multi-resource scheduling of moldable workflows. Journal of Parallel and Distributed Computing, 184(??): ??, February 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001624.

Ponce:2021:DLE

[PLBG21] Lucas M. Ponce, Daniele Lezzi, Rosa M. Badia, and Dorgival Guedes. DDF Library: Enabling functional programming in a task-based model. Journal of Parallel and Distributed Computing, 151(??):112–124, May 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000307.

Platz:2020:PCU

[PMV20] Kenneth Platz, Neeraj Mittal, and S. Venkatesan. Practical concurrent unrolled linked lists using lazy synchronization. Journal of Parallel and Distributed Computing, 139(??):110–134, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519307610.

Pournaras:2020:PWP

[Pou20] Evangelos Pournaras. Proof of witness presence: Blockchain consensus for augmented democracy in smart cities. Journal of Parallel and Distributed Computing, 145(??):160–175, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303282.

Patel:2020:DCS

[PPN^{+}20] Yashwant Singh Patel, Aditi Page, Manvi Nagdev, Anurag Choubey, Rajiv Misra, and Sajal K. Das. On demand clock synchronization for live VM migration in distributed cloud data centers. Journal of Parallel and Distributed Computing, 138(??): 15–31, April 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519303582.

Panwar:2022:DRP

[PS22] Reena Panwar and M. Supriya. Dynamic resource provisioning for service-based cloud applications: a Bayesian learning approach. Journal of Parallel and Distributed Computing, 168(??):90–107, October 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001307.

Perez:2021:SAT

[PSBB21] Borja Pérez, E. Stafford, J. L. Bosque, and R. Beivide. Sigmoid: an auto-tuned load balancing algorithm for heterogeneous systems. Journal of Parallel and Distributed Computing, 157(??):30–42, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521001325.

Penna:2021:IKC

[PSU^{+}21] Pedro Henrique Penna, João Vicente Souto, João Fellipe Uller, Márcio Castro, Henrique Freitas, and Jean-François Méhaut. Inter-kernel communication facility of a distributed operating system for NoC-based lightweight manycores. Journal of Parallel and Distributed Computing, 154(??):1–15, August 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000794.

Phisutthangkoon:2021:SPR

[PW21] Nuntipat Phisutthangkoon and Jeeraporn Werapun. Shortestpath routing for optimal all-to-all personalized-exchange embedding on hierarchical hypercube networks. Journal of Parallel and Distributed Computing, 150(??):139–154, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000046.

Peng:2022:DCA

[PWL^{+}22] Yuhuai Peng, Chenlu Wang, Qiming Li, Lei Liu, and Keping Yu. Distributed collaboration and anti-interference optimization in edge computing for IoT. Journal of Parallel and Distributed Computing, 163(??):156–165, May 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200034X.

Pal:2020:HDD

[PXY^{+}20] Soumitra Pal, Tingyang Xu, Tianbao Yang, Sanguthevar Rajasekaran, and Jinbo Bi. Hybrid-DCA: a double asynchronous approach for stochastic dual coordinate ascent. Journal of Parallel and Distributed Computing, 143(??):47–66, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302483.

Peng:2022:UUL

[PYX^{+}22] Lujie Peng, Junyu Yang, Jianbiao Xiao, Mingxue Yang, Yujiang Wang, Haojie Qin, Xiaorong Li, and Jun Zhou. ULSED: an ultra-lightweight SED model for IoT devices. Journal of Parallel and Distributed Computing, 166(??):104–110, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000867.

Park:2022:AAM

[PYYO22] Juwon Park, Daegun Yoon, Sangho Yeo, and Sangyoon Oh. AMBLE: Adjusting mini-batch and local epoch for federated learning with heterogeneous devices. Journal of Parallel and Distributed Computing, 170(??):13–23, December 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001757.

Qasem:2021:MBI

[QBS21] Apan Qasem, David P. Bunde, and Philip Schielke. A modulebased introduction to heterogeneous computing in core courses. Journal of Parallel and Distributed Computing, 158(??):56–66, December 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373152100157X.

Qiao:2022:QPV

[QDD^{+}22] Wenxuan Qiao, Ping Dong, Xiaojiang Du, Yuyang Zhang, Hongke Zhang, and Mohsen Guizani. QoS provision for vehicle big data by parallel transmission based on heterogeneous network characteristics prediction. Journal of Parallel and Distributed Computing, 163(??):83–96, May 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200017X.

Quislant:2024:EMA

[QGP24] Ricardo Quislant, Eladio Gutierrez, and Oscar Plata. Exploring multiprocessor approaches to time series analysis. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000194.

Quaranta:2021:NMH

[QM21] Lionel Quaranta and Lalith Maddegedara. A novel MPI+MPI hybrid approach combining MPI-3 shared memory windows and C11/C++11 memory model. Journal of Parallel and Distributed Computing, 157(??):125–144, November 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100143X.

Quilis:2021:CFP

[QMB21] J. Damian Segrelles Quilis, Germán Moltó, and Ignacio Blanquer. A cloud framework for problem-based learning on grid computing. Journal of Parallel and Distributed Computing, 155(??):24–37, September 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000952.

Qian:2024:ETR

[QZW^{+}24] Junyan Qian, Chuanfang Zhang, Zheng Wu, Hao Ding, and Long Li. Efficient topology reconfiguration for NoCbased multiprocessors: a greedy-memetic algorithm. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000686.

Rustam:2023:MDU

[RAJ^{+}23] Furqan Rustam, Imran Ashraf, Anca Delia Jurcut, Ali Kashif Bashir, and Yousaf Bin Zikria. Malware detection using image representation of malware data and transfer learning. Journal of Parallel and Distributed Computing, 172(??):32–50, February 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002118.

Ramesh:2021:DES

[RBS21] Shriram Ramesh, Animesh Baranawal, and Yogesh Simmhan. G ranite: a distributed engine for scalable path queries over temporal property graphs. Journal of Parallel and Distributed Computing, 151(??):94–111, May 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000253.

Rosendo:2022:DIE

[RCVA22] Daniel Rosendo, Alexandru Costan, Patrick Valduriez, and Gabriel Antoniu. Distributed intelligence on the edge-to-cloud continuum: a systematic literature review. Journal of Parallel and Distributed Computing, 166(??):71–94, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000843.

Rivera:2021:THP

[RCX^{+}21] Cody Rivera, Jieyang Chen, Nan Xiong, Jing Zhang, Shuaiwen Leon Song, and Dingwen Tao. TSM2X: High-performance tall-and-skinny matrix-matrix multiplication on GPUs. Journal of Parallel and Distributed Computing, 151(??):70–85, May 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000344.

Ruchel:2024:SAB

[RdCR^{+}24] Lucas V. Ruchel, Edson Tavares de Camargo, Luiz Antonio Rodrigues, Rogério C. Turchetti, Luciana Arantes, and Elias Procópio Duarte. Scalable atomic broadcast: a leaderless hierarchical algorithm. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001594.

Rahman:2020:BBA

[RGB20] Mohsin Ur Rahman, Barbara Guidi, and Fabrizio Baiardi. Blockchain-based access control management for decentralized online social networks. Journal of Parallel and Distributed Computing, 144(??):41–54, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302902.

Rocher-Gonzalez:2021:TEC

[RGESG^{+}21] Jose Rocher-Gonzalez, Jesus Escudero-Sahuquillo, Pedro J. García, Francisco J. Quiles, and Gaspar Mora. Towards an efficient combination of adaptive routing and queuing schemes in fat-tree topologies. Journal of Parallel and Distributed Computing, 147(??):46–63, January 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303385.

Rjaibi:2020:ESD

[RH20] Walid Rjaibi and Mohammad Hammoudeh. Enhancing and simplifying data security and privacy for multitiered applications. Journal of Parallel and Distributed Computing, 139(??): 53–64, May 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373151930632X.

Reddy:2024:EEB

[RK24] B. Naresh Kumar Reddy and Aruru Sai Kumar. Evaluating the effectiveness of Bat optimization in an adaptive and energy-efficient network-on-chip routing framework. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000170.

Rose:2020:HAB

[RKAA20] Thomas Rose, Kashif Kifayat, Sohail Abbas, and Muhammad Asim. A hybrid anomaly-based intrusion detection system to improve time complexity in the Internet of Energy environment. Journal of Parallel and Distributed Computing, 145(??):124–139, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520303191.

Rang:2024:UHM

[RLW^{+}24] Wei Rang, Huanghuang Liang, Ye Wang, Xiaobo Zhou, and Dazhao Cheng. A unified hybrid memory system for scalable deep learning and big data applications. Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001909.

Rojas:2021:UFT

[RMJM21] Elvis Rojas, Esteban Meneses, Terry Jones, and Don Maxwell. Understanding failures through the lifetime of a top-level supercomputer. Journal of Parallel and Distributed Computing, 154(??):27–41, August 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521000782.

Rojas:2024:CSE

[RPM24] Elvis Rojas, Diego Pérez, and Esteban Meneses. A characterization of soft-error sensitivity in data-parallel and model-parallel distributed deep learning. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000431.

Roy:2020:ISA

[RSGA20] Indranil Roy, Ankit Srivastava, Matt Grimm, and Srinivas Aluru. Interval stabbing on the Automata Processor. Journal of Parallel and Distributed Computing, 135(??):234–245, January 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518300510.

Reza:2023:DAM

[RSSP23] Tahsin Reza, Trevor Steil, Geoffrey Sanders, and Roger Pearce. Distributed approximate minimal Steiner trees with millions of seed vertices on billion-edge graphs. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000874.

Reano:2021:E

[RSV21] Carlos Reaño, Federico Silla, and Blesson Varghese. Editorial. Journal of Parallel and Distributed Computing, 147(??):268– 269, January 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303865.

Rubai:2022:HHB

[Rub22] Saleh Muhammad Rubai. Hybrid heuristic-based key generation protocol for intelligent privacy preservation in cloud sector. Journal of Parallel and Distributed Computing, 163(??):166– 180, May 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000119.

Ren:2021:ESM

[RWF^{+}21] Rui Ren, Chunghsuan Wu, Zhouwang Fu, Tao Song, Yanqiang Liu, Zhengwei Qi, and Haibing Guan. Efficient shuffle management for DAG computing frameworks based on the FRQ model. Journal of Parallel and Distributed Computing, 149(??):163– 173, March 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520304159.

Saraereh:2021:RAB

[SAATK21] Omar A. Saraereh, Ashraf Ali, Luae Al-Tarawneh, and Imran Khan. A robust approach for barrier-reinforcing in wireless sensor networks. Journal of Parallel and Distributed Computing, 149(??):186–192, March 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304238.

Sens:2024:SFC

[SADM24] Pierre Sens, Luciana Arantes, Anubis Graciela De Moraes Rossetto, and Olivier Marin. Stab-FD: a cooperative and adaptive failure detector for wide area networks. Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001739.

Stylianopoulos:2020:MPM

[SALP20] Charalampos Stylianopoulos, Magnus Almgren, Olaf Landsiedel, and Marina Papatriantafilou. Multiple pattern matching for network security applications: Acceleration through vectorization. Journal of Parallel and Distributed Computing, 137(??): 34–52, March 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519301984.

Stephan:2020:AII

[SATJ^{+}20] Thompson Stephan, Fadi Al-Turjman, K. Suresh Joseph, Balamurugan Balusamy, and Sweta Srivastava. Artificial intelligence inspired energy and spectrum aware cluster based routing protocol for cognitive radio sensor networks. Journal of Parallel and Distributed Computing, 142(??):90–105, August 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519306847.

Salem:2020:ELB

[SAY20] Fatima K. Abu Salem, Mira Al Arab, and Laurence T. Yang. Extending the limits for big data RSA cracking: Towards cache-oblivious TU decomposition. Journal of Parallel and Distributed Computing, 138(??):65–77, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300425.

Spiliotis:2020:PII

[SBB20] Iraklis M. Spiliotis, Michael P. Bekakos, and Yiannis S. Boutalis. Parallel implementation of the Image Block Representation using OpenMP. Journal of Parallel and Distributed Computing, 137(??):134–147, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519307622.

Souza:2021:IMP

[SBB21] Jeckson Dellagostin Souza, Pedro Henrique Exenberger Becker, and Antonio Carlos Schneider Beck. Improving multitask performance and energy consumption with partial-ISA multicores. Journal of Parallel and Distributed Computing, 153(??): 1–14, July 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000289.

Saurabh:2020:ESC

[SBBP20] Nishant Saurabh, Shajulin Benedict, Jorge G. Barbosa, and Radu Prodan. Expelliarmus: Semantic-centric virtual machine image management in IaaS clouds. Journal of Parallel and Distributed Computing, 146(??):107–121, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303415.

Schubiger:2020:GAA

[SBL20] Michel Schubiger, Goran Banjac, and John Lygeros. GPU acceleration of ADMM for large-scale quadratic programming. Journal of Parallel and Distributed Computing, 144(??):55–67, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303063.

Stillmaker:2020:SEE

[SBSB20] Aaron Stillmaker, Brent Bohnenstiehl, Lucas Stillmaker, and Bevan Baas. Scalable energy-efficient parallel sorting on a fine-grained many-core processor array. Journal of Parallel and Distributed Computing, 138(??):32–47, April 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151830532X.

Souto:2023:ICM

[SC23] João Vicente Souto and Márcio Castro. Improving concurrency and memory usage in distributed operating systems for lightweight manycores via cooperative time-sharing lightweight tasks. Journal of Parallel and Distributed Computing, 174(??): 2–18, April 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522002519.

Schryen:2024:SEC

[Sch24] Guido Schryen. Speedup and efficiency of computational parallelization: a unifying approach and asymptotic analysis. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523002058.

Sadatdiynov:2023:IHM

[SCZ^{+}23] Kuanishbay Sadatdiynov, Laizhong Cui, Lei Zhang, Joshua Zhexue Huang, Neal N. Xiong, and Chengwen Luo. An intelligent hybrid method: Multi-objective optimization for MEC-enabled devices of IoE. Journal of Parallel and Distributed Computing, 171(??):1–13, January 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001940.

Shen:2024:SDA

[SCZ24] Shuling Shen, Xinlin Chen, and Linhe Zhu. Spatiotemporal dynamics analysis and parameter optimization of a network epidemic-like propagation model based on neural network method. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731524000704.

Sudo:2020:SST

[SDLM20] Yuichi Sudo, Ajoy K. Datta, Lawrence L. Larmore, and Toshimitsu Masuzawa. Self-stabilizing token distribution on trees with constant space. Journal of Parallel and Distributed Computing, 146(??):201–211, December 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303361.

Soheili:2020:DDQ

[SEM20] Majid Soheili and Amir Masoud Eftekhari-Moghadam. DQPFS: Distributed quadratic programming based feature selection for big data. Journal of Parallel and Distributed Computing, 138(??):1–14, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519302758.

Shahid:2021:IAE

[SFML21] Arsalan Shahid, Muhammad Fahad, Ravi Reddy Manumachu, and Alexey Lastovetsky. Improving the accuracy of energy predictive models for multicore CPUs by combining utilization and performance events model variables. Journal of Parallel and Distributed Computing, 151(??):38–51, May 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000137.

Souza:2021:OMR

[SFT^{+}21] Rafael Souza, André Fernandes, Thiago S. F. X. Teixeira, George Teodoro, and Renato Ferreira. Online multimedia retrieval on CPU–GPU platforms with adaptive work partition. Journal of Parallel and Distributed Computing, 148(??):31–45, February 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520303841.

Suvizi:2023:PCA

[SFZ23] Ali Suvizi, Azim Farghadan, and Morteza Saheb Zamani. A parallel computing architecture based on cellular automata for hydraulic analysis of water distribution networks. Journal of Parallel and Distributed Computing, 178(??):11–28, August 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000527.

Saeed:2022:CLB

[SHI22] Fahad Saeed, Muhammad Haseeb, and S. S. Iyengar. Communication lower-bounds for distributed-memory computations for mass spectrometry based omics data. Journal of Parallel and Distributed Computing, 161(??):37–47, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002082.

Saikia:2021:IDA

[SK21] Parikshit Saikia and Sushanta Karmakar. Improved distributed approximation for Steiner tree in the CONGEST model. Journal of Parallel and Distributed Computing, 158(??):196–212, December 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521001702.

Simecek:2023:PAA

[ŠK23] Ivan Šimeček and Claudio Kozický. A parallel algorithm for approximating the silhouette using a ball tree. Journal of Parallel and Distributed Computing, 173(??):115–125, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002283.

Sakouhi:2021:HLG

[SKB21] Chayma Sakouhi, Abir Khaldi, and Henda Ben Ghezala. Hammer lightweight graph partitioner based on graph data volumes. Journal of Parallel and Distributed Computing, 158(??):16–28, December 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521001544.

Sabelfeld:2021:PIC

[SKK21] Karl K. Sabelfeld, Sergey Kireev, and Anastasiya Kireeva. Parallel implementation of cellular automata model of electronhole transport in a semiconductor. Journal of Parallel and Distributed Computing, 158(??):186–195, December 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001787.

Sood:2021:EEI

[SKS21] Sandeep K. Sood, Amandeep Kaur, and Vaishali Sood. Energy efficient IoT-Fog based architectural paradigm for prevention of Dengue fever infection. Journal of Parallel and Distributed Computing, 150(??):46–59, April 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304184.

Sosa:2022:AHS

[SLFC22] Juan Sebastian Sosa, Paul Leger, Hiroaki Fukuda, and Nicolás Cardozo. Ad hoc systems management and specification with distributed Petri nets. Journal of Parallel and Distributed Computing, 169(??):117–129, November 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001502.

Seisa:2024:EAE

[SLSN24] Achilleas Santi Seisa, Björn Lindqvist, Sumeet Gajanan Satpute, and George Nikolakopoulos. An edge architecture for enabling autonomous aerial navigation with embedded collision avoidance through remote nonlinear model predictive control. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000133.

Su:2024:CNC

[SLZP24] Liangkuan Su, Mingwei Lin, Jianpeng Zhang, and Yubiao Pan. CCFTL: a novel continuity compressed page-level flash address mapping method for SSDs. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000819.

Sabale:2022:PPM

[SM22] Ketan Sabale and S. Mini. Path planning mechanism for mobile anchor-assisted localization in wireless sensor networks. Journal of Parallel and Distributed Computing, 165(??):52–65, July 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000703.

Siavashi:2023:GMO

[SM23] Ahmad Siavashi and Mahmoud Momtazpour. gVMP: a multi-objective joint VM and vGPU placement heuristic for API remoting-based GPU virtualization and disaggregation in cloud data centers. Journal of Parallel and Distributed Computing, 172(??):97–113, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002234.

Singh:2021:CDS

[SMHK21] Parminder Singh, Mehedi Masud, M. Shamim Hossain, and Avinash Kaur. Cross-domain secure data sharing using blockchain for industrial IoT. Journal of Parallel and Distributed Computing, 156(??):176–184, October 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100112X.

Saxena:2020:SEE

[SMMG20] Sagar Saxena, Deekshith Shenoy Manur, Naseef Mansoor, and Amlan Ganguly. Scalable and energy efficient wireless inter chip interconnection fabrics using THz-band antennas. Journal of Parallel and Distributed Computing, 139(??):148–160, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518308177.

Suetterlein:2022:EAR

[SMMG22] Joshua Suetterlein, Joseph Manzano, Andres Marquez, and Guang R. Gao. Extending an asynchronous runtime system for high throughput applications: a case study. Journal of Parallel and Distributed Computing, 163(??):214–231, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000338.

Sorokin:2022:CPG

[SMT22] Aleksei Sorokin, Sergey Malkovsky, and Georgiy Tsoy. Comparing the performance of general matrix multiplication routine on heterogeneous computing systems. Journal of Parallel and Distributed Computing, 160(??):39–48, February 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001933.

Setlur:2020:EFT

[SNSK20] Amrith Rajagopal Setlur, S. Jaya Nirmala, Har Simrat Singh, and Sudhanshu Khoriya. An efficient fault tolerant workflow scheduling approach using replication heuristics and checkpointing in the cloud. Journal of Parallel and Distributed Computing, 136(??):14–28, February 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518306580.

Suita:2020:ECP

[SNT^{+}20] Shunsuke Suita, Takahiro Nishimura, Hiroki Tokura, Koji Nakano, Yasuaki Ito, Akihiko Kasagi, and Tsuguchika Tabaru. Efficient convolution pooling on the GPU. Journal of Parallel and Distributed Computing, 138(??):222–229, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519305052.

Sonbol:2020:EDS

[SÖAOA20] Karim Sonbol, Öznur Özkasap, Ibrahim Al-Oqily, and Moayad Aloqaily. EdgeKV: Decentralized, scalable, and consistent storage for the edge. Journal of Parallel and Distributed Computing, 144(??):28–40, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302884.

Sokolinsky:2021:BPC

[Sok21] Leonid B. Sokolinsky. BSF: a parallel computation model for scalability estimation of iterative numerical algorithms on cluster computing systems. Journal of Parallel and Distributed Computing, 149(??):193–206, March 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304251.

Sun:2022:ACP

[SOL22] Chenglong Sun, Yiming Ouyang, and Huaguo Liang. Architecting a congestion pre-avoidance and load-balanced wireless network-on-chip. Journal of Parallel and Distributed Computing, 161(??):143–154, March 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002240.

Syafiq:2024:PFE

[SOS^{+}24] Nik Amir Syafiq, Mohamed Othman, Norazak Senu, Fudziah Ismail, and Nor Asilah Wati Abdul Hamid. A parallel fractional explicit group modified AOR iterative method for solving fractional Poisson equation with multi-core architecture. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001685.

Silla:2020:IPP

[SPBR20] Federico Silla, Javier Prades, Elvira Baydal, and Carlos Reaño. Improving the performance of physics applications in atom-based clusters with rCUDA. Journal of Parallel and Distributed Computing, 137(??):160–178, March 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304034.

Sousa:2023:TSO

[SPK^{+}23] Rafael Sousa, Marcio Pereira, Yongin Kwon, Taeho Kim, Namsoon Jung, Chang Soo Kim, Michael Frank, and Guido Araujo. Tensor slicing and optimization for multicore NPUs. Journal of Parallel and Distributed Computing, 175(??):66–79, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002532.

Scravaglieri:2023:OPE

[SPP^{+}23] Lana Scravaglieri, Mihail Popov, Laércio Lima Pilla, Amina Guermouche, Olivier Aumage, and Emmanuelle Saillard. Optimizing performance and energy across problem sizes through a search space exploration and machine learning. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000904.

Sapre:2021:EMB

[SS21] Saunhita Sapre and Mini S. Emulous mechanism based multi-objective moth-flame optimization algorithm. Journal of Parallel and Distributed Computing, 150(??):15–33, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304263.

Sarroca:2024:CMA

[SSA24a] Pablo Gimeno Sarroca and Marc Sánchez-Artigas. Corrigendum to “MLLess: Achieving Cost Efficiency in Serverless Machine Learning Training” [Journal of Parallel and Distributed Computing 183 (2024) 104764]. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000352. See [SSA24b].

Sarroca:2024:MAC

[SSA24b] Pablo Gimeno Sarroca and Marc Sánchez-Artigas. MLLess: Achieving cost efficiency in serverless machine learning training. Journal of Parallel and Distributed Computing, 183(??): ??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373152300134X. See corri-

gendum [SSA24a].

Shen:2020:GPC

[SSD^{+}20] Qi Shen, Craig Sharp, Richard Davison, Gary Ushaw, Rajiv Ranjan, Albert Y. Zomaya, and Graham Morgan. A general purpose contention manager for software transactions on the GPU. Journal of Parallel and Distributed Computing, 139(??): 1–17, May 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519301376.

Saharan:2020:QEV

[SSG^{+}20] Shweta Saharan, Gaurav Somani, Gaurav Gupta, Robin Verma, Manoj Singh Gaur, and Rajkumar Buyya. QuickDedup: Efficient VM deduplication in cloud computing environments. Journal of Parallel and Distributed Computing, 139(??):18–31, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303442.

Singh:2021:FCT

[SSG21] Jagdeep Singh, Parminder Singh, and Sukhpal Singh Gill. Fog computing: a taxonomy, systematic review, current trends and research challenges. Journal of Parallel and Distributed Computing, 157(??):56–85, November 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001349.

Sanchez:2024:DQD

[SSG24] Lizeth Patricia Aguirre Sanchez, Yao Shen, and Minyi Guo. DQS: a QoS-driven routing optimization approach in SDN using deep reinforcement learning. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000157.

Shahbazinia:2023:PPC

[SSH23] Amirhossein Shahbazinia, Saber Salehkaleybar, and Matin Hashemi. ParaLiNGAM: Parallel causal structure learning for linear non-Gaussian acyclic models. Journal of Parallel and Distributed Computing, 176(??):114–127, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000138.

Shafiq:2020:AFI

[STG^{+}20] Mohsin Shafiq, Imtiaz A. Taj, Mubeen Ghafoor, Syed Ali Tariq, Assad Abbas, and Albert Y. Zomaya. Accelerating fingerprint identification using FPGA for large-scale applications. Journal of Parallel and Distributed Computing, 141(??):35–48, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518307226.

Stpiczynski:2020:ALB

[Stp20] Przemysław Stpiczyński. Algorithmic and language-based optimization of Marsa-LFIB4 pseudorandom number generator using OpenMP, OpenACC and CUDA. Journal of Parallel and Distributed Computing, 137(??):238–245, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519304885.

Soother:2022:TSR

[SUD^{+}22] Dileep Kumar Soother, Sanaullah Mehran Ujjan, Kapal Dev, Sunder Ali Khowaja, Naveed Anwar Bhatti, and Tanweer Hussain. Towards soft real-time fault diagnosis for edge devices in industrial IoT using deep domain adaptation training strategy. Journal of Parallel and Distributed Computing, 160(??):90–99, February 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521002033.

Szaban:2022:ALA

[SW22] Miroslaw Szaban and Anna Wawrzynczak. Application of the Layered Algorithm in search of an airborne contaminant source. Journal of Parallel and Distributed Computing, 159(??):1–9, January 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521001799.

Shi:2022:AFL

[SWF^{+}22] Heyuan Shi, Guyu Wang, Ying Fu, Chao Hu, Houbing Song, Jian Dong, Kun Tang, and Kai Liang. Abaci-finder: Linux kernel crash classification through stack trace similarity learning. Journal of Parallel and Distributed Computing, 168(??):70– 79, October 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522001320.

Smolka:2023:EAB

[SWM23] Sven Smolka, Leon Wißenberg, and Zoltán Ádám Mann. EdgeDecAp: an auction-based decentralized algorithm for optimizing application placement in edge computing. Journal of Parallel and Distributed Computing, 175(??):22–36, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000023.

Song:2021:JOC

[SWY^{+}21] Youmei Song, Tianyu Wo, Renyu Yang, Qi Shen, and Jie Xu. Joint optimization of cache placement and request routing in unreliable networks. Journal of Parallel and Distributed Computing, 157(??):168–178, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001350.

Song:2023:WBN

[SZQ^{+}23] Jie Song, Pengyi Zhang, Qiang Qu, Yongjie Bai, Yu Gu, and Ge Yu. Why blockchain needs graph: a survey on studies, scenarios, and solutions. Journal of Parallel and Distributed Computing, 180(??):??, October 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001004.

Singh:2022:DMT

[SZW^{+}22] Navjot Singh, Zecheng Zhang, Xiaoxiao Wu, Naijing Zhang, Siyuan Zhang, and Edgar Solomonik. Distributed-memory tensor completion for generalized loss functions in Python using new sparse tensor kernels. Journal of Parallel and Distributed Computing, 169(??):269–285, November 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001708.

Tabbakh:2024:ESC

[TA24] Abdulaziz Tabbakh and Murali Annavaram. An efficient sequential consistency implementation with dynamic race detection for GPUs. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373152300206X.

Taranco:2023:LLP

[TAG23] Raúl Taranco, José-Maria Arnau, and Antonio González. LO-CATOR: Low-power ORB accelerator for autonomous cars. Journal of Parallel and Distributed Computing, 174(??):32– 45, April 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522002507.

Torres:2023:SEO

[TAGEL23] Yuri Torres, Francisco J. Andújar, Arturo Gonzalez-Escribano, and Diego R. Llanos. Supporting efficient overlapping of host-device operations for heterogeneous programming with CtrlEvents. Journal of Parallel and Distributed Computing, 179(??):??, September 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523000722.

Tariq:2022:OPC

[TB22] Hafza Nida Tariq and Muhammad Bilal Bashir. An optimized protocol for cost effective communication in a multi-agent environment. Journal of Parallel and Distributed Computing, 169(??):24–41, November 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001459.

Tong:2021:DND

[TDCM21] Zhao Tong, Xiaomei Deng, Hongjian Chen, and Jing Mei. DDMTS: a novel dynamic load balancing scheduling scheme under SLA constraints in cloud computing. Journal of Parallel and Distributed Computing, 149(??):138–148, March 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304147.

Tang:2022:CBE

[TDL^{+}22] Songtao Tang, Xin Du, Zhihui Lu, Keke Gai, Jie Wu, Patrick C. K. Hung, and Kim-Kwang Raymond Choo. Coordinatebased efficient indexing mechanism for intelligent IoT systems in heterogeneous edge computing. Journal of Parallel and Distributed Computing, 166(??):45–56, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000892.

Titos-Gil:2020:PCP

[TGFPRA20] Rubén Titos-Gil, Ricardo Fernández-Pascual, Alberto Ros, and Manuel E. Acacio. PfTouch: Concurrent page-fault handling for Intel restricted transactional memory. Journal of Parallel and Distributed Computing, 145(??):111–123, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303166.

Thomas:2022:USG

[THPM22] Samuel Thomas, Roxana Hayne, Jonad Pulaj, and Hammurabi Mendes. Using skip graphs for increased NUMA locality. Journal of Parallel and Distributed Computing, 167(??):31–49, September 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000910.

Taneja:2023:RBN

[TK23] Harsh Taneja and Supreet Kaur. Reputation based novel trust management framework with enhanced availability for cloud. Journal of Parallel and Distributed Computing, 178(??):43–55, August 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000539.

Trotter:2020:CSI

[TLC20] James D. Trotter, Johannes Langguth, and Xing Cai. Cache simulation for irregular memory traffic on multicore CPUs: Case study on performance models for sparse matrix-vector multiplication. Journal of Parallel and Distributed Computing, 144(??):189–205, October 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302999.

Tang:2023:JSA

[TLD^{+}23] Xiaoyong Tang, Yi Liu, Tan Deng, Zexin Zeng, Haowei Huang, Qiyu Wei, Xiaorong Li, and Li Yang. A job scheduling algorithm based on parallel workload prediction on computational grid. Journal of Parallel and Distributed Computing, 171(??):88–97, January 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522001952.

Tang:2023:PMD

[TlW23] Yifeng Tang and Cho li Wang. Performance modeling on DaVinci AI core. Journal of Parallel and Distributed Computing, 175(??):134–149, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373152300014X.

Tabani:2021:PAO

[TMB^{+}21] Hamid Tabani, Fabio Mazzocchetti, Pedro Benedicte, Jaume Abella, and Francisco J. Cazorla. Performance analysis and optimization opportunities for NVIDIA automotive GPUs. Journal of Parallel and Distributed Computing, 152(??):21–32, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000290.

Toliopoulos:2022:SRT

[TNM^{+}22] Theodoros Toliopoulos, Nikodimos Nikolaidis, Anna-Valentini Michailidou, Andreas Seitaridis, Theodoros Nestoridis, Chrysa Oikonomou, Anastasios Temperekidis, Fotios Gioulekas, Anastasios Gounaris, Nick Bassiliades, Panagiotis Katsaros, Apostolos Georgiadis, and Fotis K. Liotopoulos. Sboing4Real: a real-time crowdsensing-based traffic management system. Journal of Parallel and Distributed Computing, 162(??):59–75, April 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000193.

ng:2022:SCG

[TSTY22] Nguy˜ ên Kim Th´ ăng, Abhinav Srivastav, Denis Trystram, and Paul Youssef. A stochastic conditional gradient algorithm for decentralized online convex optimization. Journal of Parallel and Distributed Computing, 169(??):334–351, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001745.

Tian:2024:MES

[TTD24] Jin Tian, JunFeng Tian, and RuiZhong Du. MSLShard: an efficient sharding-based trust management framework for blockchain-empowered IoT access control. Journal of Parallel and Distributed Computing, 185(??):??, March 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152300165X.

Tiwari:2022:PPC

[TV22] Manasi Tiwari and Sathish Vadhiyar. Pipelined preconditioned conjugate gradient methods for real and complex linear systems for distributed memory architectures. Journal of Parallel and Distributed Computing, 163(??):147–155, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000156.

Tian:2023:AOD

[TY23] Junfeng Tian and Qian Yang. An arbitrable outsourcing data audit scheme supporting credit reward and punishment and multi-user sharing. Journal of Parallel and Distributed Computing, 178(??):100–111, August 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000564.

Tong:2022:NTO

[TYM^{+}22] Zhao Tong, Feng Ye, Jing Mei, Bilan Liu, and Keqin Li. A novel task offloading algorithm based on an integrated trust mechanism in mobile edge computing. Journal of Parallel and Distributed Computing, 169(??):185–198, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200171X.

Tang:2024:HSB

[TYOC24] Wenjie Tang, Yiping Yao, Lizhen Ou, and Kai Chen. Hierarchical sort-based parallel algorithm for dynamic interest matching. Journal of Parallel and Distributed Computing, 188(??):??, June 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000315.

Tang:2024:TDT

[TZC^{+}24] Chunhua Tang, Shuangyao Zhao, Binbin Chen, Xiaonong Lu, and Qiang Zhang. A two-dimensional time-aware cloud service recommendation approach with enhanced similarity and trust. Journal of Parallel and Distributed Computing, 190(??): ??, August 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731524000534.

Tang:2021:RNA

[TZDC21] Shuyang Tang, Jilai Zheng, Yao Deng, and Qinxiang Cao. Resisting newborn attacks via shared proof-of-space. Journal of Parallel and Distributed Computing, 150(??):85–95, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304275.

Tang:2020:SDA

[TZZ^{+}20a] Qi Tang, Li-Hua Zhu, Li Zhou, Jun Xiong, and Ji-Bo Wei. Scheduling directed acyclic graphs with optimal duplication strategy on homogeneous multiprocessor systems. Journal of Parallel and Distributed Computing, 138(??):115–127, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305653.

Tang:2020:DMA

[TZZ^{+}20b] Zhuo Tang, Ailing Zeng, Xuedong Zhang, Li Yang, and Kenli Li. Dynamic memory-aware scheduling in Spark computing environment. Journal of Parallel and Distributed Computing, 141(??):10–22, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373151930824X.

UlHaque:2024:SSI

[UHAH^{+}24] Rakib Ul Haque, A. S. M. Touhidul Hasan, Mohammed Ali Mohammed Al-Hababi, Yuqing Zhang, and Dianxiang Xu. SSI − FL: Self-sovereign identity based privacypreserving federated learning. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000716.

Uhlmann:2020:MLG

[Uhl20] Jeffrey Uhlmann. On the monotonic Lagrangian grid as antecedent to the neighborhood grid. Journal of Parallel and Distributed Computing, 142(??):13–15, August 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302471.

UlHassan:2020:DPB

[URC20] Muneeb Ul Hassan, Mubashir Husain Rehmani, and Jinjun Chen. Differential privacy in blockchain technology: a futuristic approach. Journal of Parallel and Distributed Computing, 145(??):50–74, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520303105.

Upadhyay:2023:FGD

[URM23] Bhargavi R. Upadhyay, Alberto Ros, and Supriya M. Finegrain data classification to filter token coherence traffic. Journal of Parallel and Distributed Computing, 171(??):40–53, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001939.

Upadhyay:2021:ECP

[URS21] Bhargavi R. Upadhyay, Alberto Ros, and Jalpa Shah. Efficient classification of private memory blocks. Journal of Parallel and Distributed Computing, 157(??):256–268, November 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001519.

Vousden:2022:ASA

[VBB22] Mark Vousden, Graeme M. Bragg, and Andrew D. Brown. Asynchronous simulated annealing on the placement problem: a beneficial race condition. Journal of Parallel and Distributed Computing, 169(??):242–251, November 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001666.

Vishwakarma:2021:SIS

[VD21] Lokendra Vishwakarma and Debasis Das. SCAB–IoTA: Secure communication and authentication for IoT applications using blockchain. Journal of Parallel and Distributed Computing, 154(??):94–105, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000800.

Vitali:2024:GOA

[VFB^{+}24] Emanuele Vitali, Federico Ficarelli, Mauro Bisson, Davide Gadioli, Gianmarco Accordi, Massimiliano Fatica, Andrea R. Beccari, and Gianluca Palermo. GPU-optimized approaches to molecular docking-based virtual screening in drug discovery: a comparative analysis. Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001892.

Vano-Garcia:2020:KMK

[VGMG20] Fernando Vano-Garcia and Hector Marco-Gisbert. KASLR-MT: Kernel address space layout randomization for multitenant cloud systems. Journal of Parallel and Distributed Computing, 137(??):77–90, March 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731519304095.

Valero:2021:LET

[VGTSG^{+}21] Alejandro Valero, Rubén Gran-Tejero, Darío Suárez-Gracia, Emanuel A. Georgescu, Joaquín Ezpeleta, Pedro Álvarez, Adolfo Muñoz, Luis M. Ramos, and Pablo Ibáñez. A learning experience toward the understanding of abstraction-level interactions in parallel applications. Journal of Parallel and Distributed Computing, 156(??):38–52, October 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001131.

Venkataramana:2020:PMF

[VJR20] Lokeswari Venkataramana, Shomona Gracia Jacob, and Rajavel Ramadoss. A parallel multilevel feature selection algorithm for improved cancer classification. Journal of Parallel and Distributed Computing, 138(??):78–98, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303363.

Valero-Lara:2020:SFA

[VLCM^{+}20] Pedro Valero-Lara, Sandra Catalán, Xavier Martorell, Tetsuzo Usui, and Jesús Labarta. sLASs: a fully automatic auto-tuned linear algebra library based on OpenMP extensions implemented in OmpSs (LASs library). Journal of Parallel and Distributed Computing, 138(??):153–171, April 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519303417.

Vanahalli:2020:DLB

[VP20] Manjunath K. Vanahalli and Nagamma Patil. Distributed load balancing frequent colossal closed itemset mining algorithm for high dimensional dataset. Journal of Parallel and Distributed Computing, 144(??):136–152, October 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520302963.

Venetis:2022:ALI

[VP22] Ioannis E. Venetis and Astero Provata. Analysis of the leaky integrate-and-fire neuron model for GPU implementation. Journal of Parallel and Distributed Computing, 163(??):1–19, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000272.

Venugopal:2022:TLW

[VTT^{+}22] Vinu Ellampallil Venugopal, Martin Theobald, Damien Tassetti, Samira Chaychi, and Amal Tawakuli. Targeting a light-weight and multi-channel approach for distributed stream processing. Journal of Parallel and Distributed Computing, 167(??):77–96, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001022.

Wong:2022:FQA

[WC22] Renata Wong and Weng-Long Chang. Fast quantum algorithm for protein structure prediction in hydrophobic-hydrophilic model. Journal of Parallel and Distributed Computing, 164 (??):178–190, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522000673.

Wu:2024:PPO

[WCG^{+}24] Guowen Wu, Xihang Chen, Zhengjun Gao, Hong Zhang, Shui Yu, and Shigen Shen. Privacy-preserving offloading scheme in multi-access mobile edge computing based on MADRL. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001454.

Wang:2021:PPB

[WCLD21] Ruoxi Wang, Chao Chen, Jonghyun Lee, and Eric Darve. PBBFMM3D: a parallel black-box algorithm for kernel matrix–vector multiplication. Journal of Parallel and Distributed Computing, 154(??):64–73, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000824.

Wang:2024:MMS

[WCMS24] Sheng Wang, Shiping Chen, Fei Meng, and Yumei Shi. MSHGN: Multi-scenario adaptive hierarchical spatial graph convolution network for GPU utilization prediction in heterogeneous GPU clusters. Journal of Parallel and Distributed Computing, 184(??):??, February 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731523001661.

Wang:2020:TCE

[WCR^{+}20] Yang Wang, Shan Cao, Hongshuai Ren, Jianjun Li, Kejiang Ye, Chengzhong Xu, and Xi Chen. Towards cost-effective service migration in mobile edge: a Q-learning approach. Journal of Parallel and Distributed Computing, 146(??):175–188, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303488.

Wu:2022:FBP

[WCT^{+}22] Zhiqiang Wu, Zhubin Cai, Xiaoyong Tang, Yuming Xu, and Tan Deng. A forward and backward private oblivious RAM for storage outsourcing on edge-cloud computing. Journal of Parallel and Distributed Computing, 166(??):1–14, August 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000855.

Wu:2022:HDH

[WCTW22] Guoqing Wu, Liqiang Cao, Hongyun Tian, and Wei Wang. HY-DBSCAN: a hybrid parallel DBSCAN clustering algorithm scalable on distributed-memory computers. Journal of Parallel and Distributed Computing, 168(??):57–69, October 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200140X.

Wang:2022:UNC

[WDL22] Yuyang Wang, Dezun Dong, and Fei Lei. Understanding node connection modes in Multi-Rail Fat-tree. Journal of Parallel and Distributed Computing, 167(??):199–210, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000971.

Wu:2024:RCB

[WDY^{+}24] Xiangyu Wu, Xuehui Du, Qiantao Yang, Na Wang, and Wenjuan Wang. Redactable consortium blockchain based on verifiable distributed chameleon hash functions. Journal of Parallel and Distributed Computing, 183(??):??, January 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001478.

Wang:2022:FTH

[WFS^{+}22] Yihong Wang, Jianxi Fan, Xueli Sun, Baolei Cheng, and Yan Wang. Fault-tolerability of the hypercube and variants with faulty subcubes. Journal of Parallel and Distributed Computing, 167(??):148–156, September 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001101.

Wei:2021:PID

[WHC21] Chao Wei, Rong-Xia Hao, and Jou-Ming Chang. Packing internally disjoint Steiner trees to compute the κ_{3}- connectivity in augmented cubes. Journal of Parallel and Distributed Computing, 154(??):42–53, August 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000812.

Wu:2023:EAM

[WHL^{+}23] Zhiwei Wu, Li Han, Jing Liu, Yves Robert, and Frédéric Vivien. Energy-aware mapping and scheduling strategies for real-time workflows under reliability constraints. Journal of Parallel and Distributed Computing, 176(??):1–16, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000199.

Wang:2021:RBB

[WHY^{+}21] Huaqun Wang, Debiao He, Jia Yu, Neal N. Xiong, and Bin Wu. RDIC: a blockchain-based remote data integrity checking scheme for IoT in 5G networks. Journal of Parallel and Distributed Computing, 152(??):1–10, June 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000332.

Wang:2023:FEP

[WLL^{+}23a] Jing Wang, Chao Li, Yibo Liu, Taolei Wang, Junyi Mei, Lu Zhang, Pengyu Wang, and Minyi Guo. Fargraph+: Excavating the parallelism of graph processing workload on RDMA-based far memory system. Journal of Parallel and Distributed Computing, 177(??):144–159, July 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000345.

Wang:2023:ADR

[WLL^{+}23b] Weidong Wang, Dian Li, Wangda Luo, Yujian Kang, and Liqiang Wang. Anthropomorphic diagnosis of runtime hidden behaviors in OpenMP multi-threaded applications. Journal of Parallel and Distributed Computing, 177(??):17–27, July 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000333.

Wang:2020:GBE

[WLZ20] Ziyu Wang, Nanqing Luo, and Pan Zhou. GuardHealth: Blockchain empowered secure data management and graph convolutional network enabled anomaly detection in smart healthcare. Journal of Parallel and Distributed Computing, 142(??): 1–12, August 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519308470.

Wang:2020:EBD

[WMJ^{+}20] Tian Wang, Yaxin Mei, Weijia Jia, Xi Zheng, Guojun Wang, and Mande Xie. Edge-based differential privacy computing for sensor-cloud systems. Journal of Parallel and Distributed Computing, 136(??):75–85, February 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373151930293X.

Wang:2021:SAT

[WWH^{+}21] Bo Wang, Changhai Wang, Wanwei Huang, Ying Song, and Xiaoyun Qin. Security-aware task scheduling with deadline constraints on heterogeneous hybrid clouds. Journal of Parallel and Distributed Computing, 153(??):15–28, July 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000563.

Wang:2021:VDL

[WWL^{+}21] Zhaokang Wang, Shen Wang, Junhong Li, Chunfeng Yuan, Rong Gu, and Yihua Huang. VSIM: Distributed local structural vertex similarity calculation on big graphs. Journal of Parallel and Distributed Computing, 158(??):29–46, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001556.

Wang:2022:ASM

[WXAL22] Zhe Wang, Chen Xu, Kunal Agrawal, and Jing Li. Adaptive scheduling of multiprogrammed dynamic-multithreading applications. Journal of Parallel and Distributed Computing, 162(??): 76–88, April 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000144.

Wu:2023:MAD

[WXZ^{+}23] Guowen Wu, Zhiqi Xu, Hong Zhang, Shigen Shen, and Shui Yu. Multi-agent DRL for joint completion delay and energy consumption with queuing theory in MEC-based IIoT. Journal of Parallel and Distributed Computing, 176(??):80–94, June 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000291.

Wang:2021:ISR

[WYA^{+}21] MengJie Wang, Xiaomin Yang, Marco Anisetti, Rongzhu Zhang, Marcelo Keese Albertini, and Kai Liu. Image superresolution via enhanced multi-scale residual network. Journal of Parallel and Distributed Computing, 152(??):57–66, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100037X.

Wang:2021:SOD

[WYH^{+}21] Chun-Yi Wang, Chi-Yu You, Fu-Hau Hsu, Chia-Hao Lee, Che-Hao Liu, and YungYu Zhuang. SMS Observer: a dynamic mechanism to analyze the behavior of SMS-based malware. Journal of Parallel and Distributed Computing, 156(??):25–37, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100109X.

Wang:2023:ISI

[WYH^{+}23] Haotian Wang, Wangdong Yang, Rong Hu, Renqiu Ouyang, Kenli Li, and Keqin Li. IAP-SpTV: an input-aware adaptive pipeline SpTV via GCN on CPU-GPU. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001119.

Wu:2024:HRC

[WYH^{+}24] Jiaxi Wu, Wenquan Yang, Xinming Han, Yunzhe Qiu, Andrei Gudkov, and Jie Song. Hotspot resolution in cloud computing: a Γ-robust knapsack approach for virtual machine migration. Journal of Parallel and Distributed Computing, 186(??): ??, April 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523001879.

Wang:2020:BVR

[WYW^{+}20] En Wang, Yongjian Yang, Jie Wu, Kaihao Lou, Wenbin Liu, and Yuanbo Xu. Budgeted video replacement policy in mobile crowdsensing. Journal of Parallel and Distributed Computing, 136(??):1–13, February 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731518305951.

Wang:2024:CEE

[WYZ^{+}24] Shuang Wang, Zian Yuan, Xiaodong Zhang, Jiawen Wu, and Yamin Wang. Cloud-edge-end workflow scheduling with multiple privacy levels. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731524000467.

Wei:2020:PRA

[WZC^{+}20] Mengze Wei, Wenyi Zhao, Quan Chen, Hao Dai, Jingwen Leng, Chao Li, Wenli Zheng, and Minyi Guo. Predicting and reining in application-level slowdown on spatial multitasking GPUs. Journal of Parallel and Distributed Computing, 141(??):99–114, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519307361.

Wang:2021:EES

[WZO^{+}21] Yi Wang, Hao Zhang, Kwang-Il Oh, Jae-Jin Lee, and Seok-Bum Ko. Energy efficient spiking neural network processing using approximate arithmetic units and variable precision weights. Journal of Parallel and Distributed Computing, 158(??):164–175, December 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521001696.

Xu:2020:MRE

[XB20] Minxian Xu and Rajkumar Buyya. Managing renewable energy and carbon footprint in multi-cloud computing environments. Journal of Parallel and Distributed Computing, 135(??):191– 202, January 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519303132.

Xu:2022:SPS

[XBX^{+}22] Guangquan Xu, Hongpeng Bai, Jun Xing, Tao Luo, Neal N. Xiong, Xiaochun Cheng, Shaoying Liu, and Xi Zheng. SG-PBFT: a secure and highly efficient distributed blockchain PBFT consensus algorithm for intelligent Internet of Vehicles. Journal of Parallel and Distributed Computing, 164(??): 1–11, June 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000363.

Xu:2022:FSR

[XDM^{+}22] Kai Xu, Xiaohui Duan, André Müller, Robin Kobus, Bertil Schmidt, and Weiguo Liu. FMapper: Scalable read mapper based on succinct hash index on SunWay TaihuLight. Journal of Parallel and Distributed Computing, 161(??):72–82, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521002112.

Xue:2021:IFG

[XJR21] Weicheng Xue, Charles W. Jackson, and Christoper J. Roy. An improved framework of GPU computing for CFD applications on structured grids using OpenACC. Journal of Parallel and Distributed Computing, 156(??):64–85, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001155.

Xiao:2020:MEM

[XLCL20] Guoqing Xiao, Jingning Li, Yuedan Chen, and Kenli Li. MalFCS: an effective malware classification framework with automated feature extraction based on deep convolutional neural networks. Journal of Parallel and Distributed Computing, 141(??):49–58, July 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S074373151930869X.

Xu:2020:AAR

[XLL^{+}20] Yajing Xu, Junnan Li, Zhihui Lu, Jie Wu, Patrick C. K. Hung, and Abdulhameed Alelaiwi. ARVMEC: Adaptive recommendation of virtual machines for IoT in edge-cloud environment. Journal of Parallel and Distributed Computing, 141(??):23– 34, July 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520302045.

Xu:2021:BBR

[XLL^{+}21] Zisang Xu, Wei Liang, Kuan-Ching Li, Jianbo Xu, and Hai Jin. A blockchain-based roadside unit-assisted authentication and key agreement protocol for Internet of Vehicles. Journal of Parallel and Distributed Computing, 149(??):29–39, March 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304044.

Xu:2022:SSE

[XMJG22] Chengshuo Xu, Abbas Mazloumi, Xiaolin Jiang, and Rajiv Gupta. SimGQ+: Simultaneously evaluating iterative pointto-all and point-to-point graph queries. Journal of Parallel and Distributed Computing, 164(??):12–27, June 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000120.

Xu:2022:AOS

[XPW^{+}22] Jinlai Xu, Balaji Palanisamy, Qingyang Wang, Heiko Ludwig, and Sandeep Gopisetty. Amnis: Optimized stream processing for edge computing. Journal of Parallel and Distributed Computing, 160(??):49–64, February 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001921.

Xu:2021:TDS

[XRBT21] Jun Xu, Rouhollah Rahmatizadeh, Ladislau Bölöni, and Damla Turgut. A taxi dispatch system based on prediction of demand and destination. Journal of Parallel and Distributed Computing, 157(??):269–279, November 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001489.

Xia:2021:DRA

[XTGJ21] Zhuoqun Xia, Jingjing Tan, Ke Gu, and WeiJia Jia. Detection resource allocation scheme for two-layer cooperative IDSs in smart grids. Journal of Parallel and Distributed Computing, 147(??):236–247, January 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303725.

Xiao:2022:MRC

[XWCJ22] Zheng Xiao, Mengyuan Wang, Anthony Theodore Chronopoulos, and Jiuchuan Jiang. A method for reducing cloud service request peaks based on game theory. Journal of Parallel and Distributed Computing, 165(??):107–119, July 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000582.

Xu:2023:LLO

[XWJ^{+}23] Xianghao Xu, Fang Wang, Hong Jiang, Yongli Cheng, Yu Hua, Dan Feng, and Yongxuan Zhang. LOSC: a locality-optimized subgraph construction scheme for out-of-core graph processing. Journal of Parallel and Distributed Computing, 172(??):51–68, February 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522002155.

Xu:2020:MPV

[XWL^{+}20] Li Xu, Yueyao Wang, Thomas Lux, Tyler Chang, Jon Bernard, Bo Li, Yili Hong, Kirk Cameron, and Layne Watson. Modeling I/O performance variability in high-performance computing systems using mixture distributions. Journal of Parallel and Distributed Computing, 139(??):87–98, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519302746.

Xie:2024:FVG

[XXZ^{+}24] Wei Xie, Runqun Xiong, Jinghui Zhang, Jiahui Jin, and Junzhou Luo. Federated variational generative learning for heterogeneous data in distributed environments. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000807.

Xie:2022:NBB

[XZH^{+}22] Mande Xie, Qiting Zhao, Haibo Hong, Chen Chen, and Jun Yu. A novel blockchain-based and proxy-oriented public audit scheme for low performance terminal devices. Journal of Parallel and Distributed Computing, 169(??):58–71, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001319.

Xu:2024:EET

[XZPL24] Hongzhi Xu, Binlian Zhang, Chen Pan, and Keqin Li. Energyefficient triple modular redundancy scheduling on heterogeneous multi-core real-time systems. Journal of Parallel and Distributed Computing, 191(??):??, September 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000790.

Xiong:2023:ESO

[XZY^{+}23] Zhi Xiong, Min Zhao, Ziyue Yuan, Jianlong Xu, and Lingru Cai. Energy-saving optimization of application server clusters based on mixed integer linear programming. Journal of Parallel and Distributed Computing, 171(??):111–129, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200199X.

Yazdanpanah:2023:LPW

[Yaz23a] Fahimeh Yazdanpanah. A low-power WNoC transceiver with a novel energy consumption management scheme for dependable IoT systems. Journal of Parallel and Distributed Computing, 172(??):144–158, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200226X.

Yazdanpanah:2023:TLN

[Yaz23b] Fahimeh Yazdanpanah. A two-level network-on-chip architecture with multicast support. Journal of Parallel and Distributed Computing, 172(??):114–130, February 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002258.

Yu:2022:CSD

[YDX^{+}22] Enda Yu, Dezun Dong, Yemao Xu, Shuo Ouyang, and Xiangke Liao. CP-SGD: Distributed stochastic gradient descent with compression and periodic compensation. Journal of Parallel and Distributed Computing, 169(??):42–57, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001290.

Yantir:2022:HSC

[YES22] Hasan Erdem Yantır, Ahmed M. Eltawil, and Khaled N. Salama. A hardware/software co-design methodology for inmemory processors. Journal of Parallel and Distributed Computing, 161(??):63–71, March 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521002070.

Young:2020:DBO

[YHKR20] M. Todd Young, Jacob D. Hinkle, Ramakrishnan Kannan, and Arvind Ramanathan. Distributed Bayesian optimization of deep reinforcement learning algorithms. Journal of Parallel and Distributed Computing, 139(??):43–52, May 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519300231.

Yang:2023:EJR

[YLDY23] Wenxiang Yang, Xiangke Liao, Dezun Dong, and Jie Yu. Exploring job running path to predict runtime on multiple production supercomputers. Journal of Parallel and Distributed Computing, 175(??):109–120, May 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000011.

Yin:2021:SCS

[YLL21] Luxiu Yin, Pengfei Li, and Juan Luo. Smart contract service migration mechanism based on container in edge computing. Journal of Parallel and Distributed Computing, 152(??):157– 166, June 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000447.

Yue:2020:BBV

[YLZ^{+}20] Dongdong Yue, Ruixuan Li, Yan Zhang, Wenlong Tian, and Yongfeng Huang. Blockchain-based verification framework for data integrity in edge-cloud storage. Journal of Parallel and Distributed Computing, 146(??):1–14, December 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303142.

Yilmazer-Metin:2021:GWA

[YM21] Ayse Yilmazer-Metin. Graph-waving architecture: Efficient execution of graph applications on GPUs. Journal of Parallel and Distributed Computing, 148(??):69–82, February 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303889.

Yasudo:2022:GAS

[YNI^{+}22] Ryota Yasudo, Koji Nakano, Yasuaki Ito, Ryota Katsuki, Yusuke Tabata, Takashi Yazane, and Kenichiro Hamano. GPUaccelerated scalable solver with bit permutated cyclic-min algorithm for quadratic unconstrained binary optimization. Journal of Parallel and Distributed Computing, 167(??):109–122, September 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731522000946.

Yazdinejad:2020:COS

[YPD^{+}20] Abbas Yazdinejad, Reza M. Parizi, Ali Dehghantanha, Gautam Srivastava, Senthilkumar Mohan, and Abedallah M. Rababah. Cost optimization of secure routing with untrusted devices in software defined networking. Journal of Parallel and Distributed Computing, 143(??):36–46, September 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519306288.

Yin:2020:FGA

[YQZ^{+}20] Hui Yin, Zheng Qin, Jixin Zhang, Hua Deng, Fangmin Li, and Keqin Li. A fine-grained authorized keyword secure search scheme with efficient search permission update in cloud computing. Journal of Parallel and Distributed Computing, 135(??):56– 69, January 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519303004.

Yarinezhad:2021:OCB

[YS21] Ramin Yarinezhad and Masoud Sabaei. An optimal clusterbased routing algorithm for lifetime maximization of Internet of Things. Journal of Parallel and Distributed Computing, 156(??):7–24, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731521001106.

Younis:2021:DPP

[YSMB21] Awad A. Younis, Rajshekhar Sunderraman, Mike Metzler, and Anu G. Bourgeois. Developing parallel programming and soft skills: a project based learning approach. Journal of Parallel and Distributed Computing, 158(??):151–163, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001611.

Yang:2023:DIA

[YSZL23] Lei Yang, Xiaoyuan Shen, Changyi Zhong, and Yuwei Liao. On-demand inference acceleration for directed acyclic graph neural networks over edge-cloud collaboration. Journal of Parallel and Distributed Computing, 171(??):79–87, January 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001964.

Yu:2022:TTM

[YTLF22] Jinyu Yu, Wei Tong, Pengze Lv, and Dan Feng. TERMS: Task management policies to achieve high performance for mixed workloads using surplus resources. Journal of Parallel and Distributed Computing, 170(??):74–85, December 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001897.

Yang:2021:IRA

[YWF21] Fan Yang, Huaqiong Wang, and Jianjing Fu. Improvement of recommendation algorithm based on collaborative deep learning and its parallelization on Spark. Journal of Parallel and Distributed Computing, 148(??):58–68, February 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303750. See corrigendum [YWF23].

Yang:2023:CIR

[YWF23] Fan Yang, Huaqiong Wang, and Jianjing Fu. Corrigendum to “Improvement of recommendation algorithm based on Collaborative Deep Learning and its Parallelization on Spark” [J. Parallel Distrib. Comput. 148 (2021) 58–68]. Journal of Parallel and Distributed Computing, 174(??):1, April 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002453. See [YWF21].

Yeh:2022:RDR

[YY22] Tsozen Yeh and Shengchieh Yu. Realizing dynamic resource orchestration on cloud systems in the cloud-to-edge continuum. Journal of Parallel and Distributed Computing, 160(??):100– 109, February 2022. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521002045.

Yan:2022:CDL

[YZC22] Ke Yan, Xiaokang Zhou, and Jinjun Chen. Collaborative deep learning framework on IoT data with bidirectional NLSTM neural networks for energy consumption forecasting. Journal of Parallel and Distributed Computing, 163(??):248–255, May 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000247.

Yang:2023:FTA

[YZM23] Yayu Yang, Mingzu Zhang, and Jixiang Meng. Fault tolerance analysis for Hamming graphs with large-scale faulty links based on k-component edge-connectivity. Journal of Parallel and Distributed Computing, 173(??):107–114, March 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522002441.

Zhang:2020:PAF

[ZAB20] Yongzhe Zhang, Ariful Azad, and Aydin Buluç. Parallel algorithms for finding connected components using linear algebra. Journal of Parallel and Distributed Computing, 144(??):14–27, October 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731520302689.

Zulian:2024:DCW

[ZBF^{+}24] P. Zulian, S. Ben Bader, G. Fourestey, R. Krause, and D. Rossinelli. Data-centric workloads with MPI Sort. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523002034.

Zhou:2021:CDS

[ZCD^{+}21] Jiang Zhou, Yong Chen, Dong Dai, Yu Zhuang, and Weiping Wang. I/O characteristic discovery for storage system optimizations. Journal of Parallel and Distributed Computing, 148(??):1–13, February 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303452.

Zhao:2021:RRN

[ZCY^{+}21] Shuang Zhao, Shuhui Chen, Hui Yang, Fei Wang, and Ziling Wei. RF-RISA: a novel flexible random forest accelerator based on FPGA. Journal of Parallel and Distributed Computing, 157(??):220–232, November 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001477.

Zode:2022:OEC

[ZD22] Pravin Zode and Raghavendra Deshmukh. Optimization of elliptic curve scalar multiplication using constraint based scheduling. Journal of Parallel and Distributed Computing, 167 (??):232–239, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731522001216.

Zrigui:2022:IPB

[ZdCLT22] Salah Zrigui, Raphael Y. de Camargo, Arnaud Legrand, and Denis Trystram. Improving the performance of batch schedulers using online job runtime classification. Journal of Parallel and Distributed Computing, 164(??):83–95, June 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000090.

Zeni:2024:SKO

[ZDD^{+}24] Alberto Zeni, Emanuele Del Sozzo, Eleonora D’Arnese, Davide Conficconi, and Marco D. Santambrogio. Starlight: a kernel optimizer for GPU processing. Journal of Parallel and Distributed Computing, 187(??):??, May 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523002022.

Zhang:2024:SML

[ZDL^{+}24] Xiaoyun Zhang, Dezun Dong, Cunlu Li, Shaocong Wang, and Liquan Xiao. A survey of machine learning for Network-on-Chips. Journal of Parallel and Distributed Computing, 186(??): ??, April 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S074373152300148X.

Zhao:2021:EUB

[ZDZ^{+}21] Jia Zhao, Yan Ding, Yunan Zhai, Yuqiang Jiang, Yujuan Zhai, and Ming Hu. Explore unlabeled big data learning to online failure prediction in safety-aware cloud environment. Journal of Parallel and Distributed Computing, 153(??):53–63, July 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000460.

Zhao:2023:VPA

[ZFL^{+}23] Hui Zhao, Nanzhi Feng, Jianhua Li, Guobin Zhang, Jing Wang, Quan Wang, and Bo Wan. VM performance-aware virtual machine migration method based on ant colony optimization in cloud environment. Journal of Parallel and Distributed Computing, 176(??):17–27, June 2023. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000187.

Zhang:2024:LDH

[ZGTM24] Haitao Zhang, Tongyu Guo, Wei Tian, and Huadong Ma. Learning-driven hybrid scaling for multi-type services in cloud. Journal of Parallel and Distributed Computing, 189(??):??, July 2024. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000443.

Zhang:2023:EFD

[Zha23] Jianwei Zhang. Extreme flow decomposition for multi-source multicast with intra-session network coding. Journal of Parallel and Distributed Computing, 175(??):80–91, May 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523000035.

Zhang:2021:CPT

[ZJW^{+}21] Yongxuan Zhang, Hong Jiang, Fang Wang, Yu Hua, Dan Feng, Yongli Cheng, Yuchong Hu, and Renzhi Xiao. CIC-PIM: Trading spare computing power for memory space in graph processing. Journal of Parallel and Distributed Computing, 147(??):152–165, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520303695.

Zhang:2022:MDA

[ZJW22] Zhaorui Zhang, Zhuoran Ji, and Choli Wang. Momentumdriven adaptive synchronization model for distributed DNN training on HPC clusters. Journal of Parallel and Distributed Computing, 159(??):65–84, January 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152100191X.

Zhang:2021:HPG

[ZKL21] Tao Zhang, Wang Kan, and Xiao-Yang Liu. High performance GPU primitives for graph-tensor learning operations. Journal of Parallel and Distributed Computing, 148(??):125–137, February 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304007.

Zhuang:2024:PDP

[ZLCL24] Hongbin Zhuang, Xiao-Yan Li, Jou-Ming Chang, and Ximeng Liu. Paired 2-disjoint path covers of k-ary n-cubes under the partitioned edge fault model. Journal of Parallel and Distributed Computing, 190(??):??, August 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731524000510.

Zhang:2022:HAA

[ZLD22] Haibin Zhang, Chao Liu, and Sisi Duan. How to achieve adaptive security for asynchronous BFT? Journal of Parallel and Distributed Computing, 169(??):252–268, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001678.

Zhou:2022:IEC

[ZLLC22] Zhou Zhou, Yangfan Li, Fangmin Li, and Hongbing Cheng. An intelligence energy consumption model based on BP neural network in mobile edge computing. Journal of Parallel and Distributed Computing, 167(??):211–220, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001149.

Zhang:2023:PDL

[ZLS23] Xiangping Zhang, Jianxun Liu, and Min Shi. A parallel deep learning-based code clone detection model. Journal of Parallel and Distributed Computing, 181(??):??, November 2023. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152300117X.

Zhang:2021:TID

[ZNX^{+}21] Cheng Zhang, Zhifei Ni, Yang Xu, Entao Luo, Linweiya Chen, and Yaoxue Zhang. A trustworthy industrial data management scheme based on redactable blockchain. Journal of Parallel and Distributed Computing, 152(??):167–176, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000472.

Zacarias:2021:ICH

[ZPN^{+}21] Felippe Vieira Zacarias, Vinicius Petrucci, Rajiv Nishtala, Paul Carpenter, and Daniel Mossé. Intelligent colocation of HPC workloads. Journal of Parallel and Distributed Computing, 151(??):125–137, May 2021. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000319.

Zheng:2021:DBC

[ZQL^{+}21] Jiaqi Zheng, Liulan Qin, Kexin Liu, Bingchuan Tian, Chen Tian, Bo Li, and Guihai Chen. Django: Bilateral coflow scheduling with predictive concurrent connections. Journal of Parallel and Distributed Computing, 152(??):45–56, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000125.

Zuk:2022:RRL

[ZR22] Pawel Zuk and Krzysztof Rzadca. Reducing response latency of composite functions-as-a-service through scheduling. Journal of Parallel and Distributed Computing, 167(??):18–30, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522000909.

Zhai:2023:HHC

[ZSL^{+}23] Wenbin Zhai, Shanshan Sun, Liang Liu, Youwei Ding, and Wanying Lu. HOTD: a holistic cross-layer time-delay attack detection framework for unmanned aerial vehicle networks. Journal of Parallel and Distributed Computing, 177(??):117– 130, July 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000370.

Zhou:2020:EOP

[ZT20] Hongyang Zhou and Gábor Tóth. Efficient OpenMP parallelization to a complex MPI parallel magnetohydrodynamics code. Journal of Parallel and Distributed Computing, 139(??): 65–74, May 2020. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731519304903.

Zhai:2021:HPF

[ZTKL^{+}21] Yanlong Zhai, Jude Tchaye-Kondi, Kwei-Jay Lin, Liehuang Zhu, Wenjun Tao, Xiaojiang Du, and Mohsen Guizani. Hadoop Perfect File: a fast and memory-efficient metadata access archive file to face small files problem in HDFS. Journal of Parallel and Distributed Computing, 156(??):119–130, October 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001234.

Zhang:2021:PEM

[ZWCL21] Yulai Zhang, Jiachen Wang, Gang Cen, and Kueiming Lo. Parallel ensemble methods for causal direction inference. Journal of Parallel and Distributed Computing, 150(??):96–103, April 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520304287.

Zhang:2020:BBB

[ZWS^{+}20] Junwei Zhang, Zhuzhu Wang, Lei Shang, Di Lu, and Jianfeng Ma. BTNC: a blockchain based trusted network connection protocol in IoT. Journal of Parallel and Distributed Computing, 143(??):1–16, September 2020. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731519308457.

Zhang:2022:PTR

[ZWSL22] Mi Zhang, Qiuping Wang, Zhirong Shen, and Patrick P. C. Lee. POCache: Toward robust and configurable straggler tolerance with parity-only caching. Journal of Parallel and Distributed Computing, 167(??):157–172, September 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001137.

Zhang:2021:RAA

[ZXY21] Qifan Zhang, Liqiong Xu, and Weihua Yang. Reliability analysis of the augmented cubes in terms of the extra edgeconnectivity and the component edge-connectivity. Journal of Parallel and Distributed Computing, 147(??):124–131, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152030349X.

Zhang:2021:PAO

[ZYL^{+}21] Yufeng Zhang, Wangdong Yang, Kenli Li, Dahai Tang, and Keqin Li. Performance analysis and optimization for SpMV based on aligned storage formats on an ARM processor. Journal of Parallel and Distributed Computing, 158(??):126–137, December 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521001684.

Zhang:2024:AQM

[ZYW24] Changzhen Zhang, Jun Yang, and Ning Wang. An active queue management for wireless sensor networks with priority scheduling strategy. Journal of Parallel and Distributed Computing, 187 (??):??, May 2024. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731524000121.

Zhao:2022:ECA

[ZZ22] Daming Zhao and Jiantao Zhou. An energy and carbonaware algorithm for renewable energy usage maximization in distributed cloud data centers. Journal of Parallel and Distributed Computing, 165(??):156–166, July 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S074373152200079X.

Zhang:2024:CMC

[ZZ24] Hong Zhang and Shuming Zhou. Characterization of matroidal connectivity of regular networks. Journal of Parallel and Distributed Computing, 186(??):??, April 2024. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731523001880.

Zhu:2022:RAN

[ZZLM22] Rui Zhu, Xue-Qian Zeng, Xiang-Jun Li, and Meijie Ma. Reliability of arrangement networks in terms of the hrestricted edge connectivity. Journal of Parallel and Distributed Computing, 170(??):68–73, December 2022. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731522001873.

Zeng:2021:AES

[ZZS^{+}21a] Hanqing Zeng, Hongkuan Zhou, Ajitesh Srivastava, Rajgopal Kannan, and Viktor Prasanna. Accurate, efficient and scalable training of graph neural networks. Journal of Parallel and Distributed Computing, 147(??):166–183, January 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731520303579.

Zhang:2021:NCR

[ZZS^{+}21b] Bo Zhang, Zeng Zeng, Xiupeng Shi, Jianxi Yang, Bharadwaj Veeravalli, and Keqin Li. A novel cooperative resource provisioning strategy for multi-cloud load balancing. Journal of Parallel and Distributed Computing, 152(??):98–107, June 2021. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0743731521000241.

Zhong:2020:EPD

[ZZZ^{+}20] Kai Zhong, Xu Zhou, Liqian Zhou, Zhibang Yang, Chubo Liu, and Na Xiao. An efficient parallel direction-based clustering algorithm. Journal of Parallel and Distributed Computing, 145(??):24–33, November 2020. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0743731520303099.

Zhang:2023:RRA

[ZZZ^{+}23] Yang Zhang, Kaige Zhu, Xuan Zhao, Quancheng Zhao, Zhenjiang Zhang, and Ali Kashif Bashir. Research on resource allocation technology in highly trusted environment of edge computing. Journal of Parallel and Distributed Computing, 178(??):29– 42, August 2023. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731523000552.

Zhang:2021:NHR

[ZZZG21] Xudong Zhang, Liang Zhao, Wei Zhong, and Feng Gu. A novel hybrid resampling algorithm for parallel/distributed particle filters. Journal of Parallel and Distributed Computing, 151(??): 24–37, May 2021. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0743731521000265.